A Bibliography of Publications of Per-Olov Löwdin

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

04 June 2020 Version 2.61

Abstract This bibliography records publications of the late Per-Olov Löwdin.

Dedication

This bibliography is dedicated in fond memory of, and to, Per-Olov Löwdin (28 October 1916 – 6 October 2000): advisor, alpinist, friend, host, leader, linear algebraist, mentor, musician, role model, scientist, and teacher of a great many of us in his extended scientific family.

Title word cross-reference

$14.50 [Kut67b]. $17.00 [Kön68]. 1s [HL59, HL65]. 2 [RG76, SW92, TL81]. $22 [Kle66b]. $225.00 [Boy05]. $27 [Löw69c]. $450.00 [Boy04]. $67.50 [Ahl83]. ^{−} [MY04]. ^{1}Σ_{g}^{+} [HL76]. _{3} [MY04, VB75]. α [Jon84, JBW87, Jon92, LSZ^{+}00, MK08, MK09, MÇ10a, MÇ10b, MÇ12]. b_{Kk}(LM|l) [Suz85]. c_{12} [Löw48d]. c_{44} [Löw48d]. d [LAS99]. E [JBW87]. h_{n,2n−i}(LM|l) [Suz85]. h_{n,2n−i}(LM|l)/(2n − i)! [Suz90]. i [Suz85]. J = (1/2)(E_{singlet} − E_{triplet}) [Har76]. n [Suz85]. n + ℓ, n [AK02]. Ω [WS82]. r_{12} [CL62]. s [LAS99]. SU_{p,q}(2) [WS92]. SU_{pq}(2) [SW92]. T [WS82]. U [Lar04]. v [Lev76].

-Aminopurine [RG76]. -function [Jon84, Jon92]. -Hydrogen [HL65, HL59]. -Orbital [LAS99]. -Parameter [SW92]. -radial [MK08, MK09, MÇ10a, MÇ10b]. -representability [Lev76].

1 [Boy05, FH58]. 1-4020-1286-1 [Boy04]. 1-4020-2583-1 [Boy05]. 13 [VBHG95]. 1520 [Roo00]. 1911 [Meh75]. 1955 [FHL55, LFH55, Löw87a]. 1958 [Löw58f]. 1963 [Löw63d, Löw63g]. 1978 [LCG82a]. 1979 [Löw80a]. 1988 [LGBC89, LGJLCB89]. 1989 [Löw90i]. 1991 [Löw91c].

2000 [Lin03]. 25 [Suz84a]. 26 [Suz87].

3rd [Löw80a, Löw80f].

500th [LGC78b].

60th [BDL^{+}94]. 6th [Löw89a, Löw89d].

86j [Suz87].

90-years [LGJLCB89]. 90th [LGBC89]. 97 [Löw87a].

A. [ES87, Löw87d]. åldrande [Löw77d]. Abb [Kle66b, Kut67b]. Academic [Boy04, Boy05, Kle66b, Kön68, Kut67b]. Accuracy [Par09]. accurate [MÇ11a]. acid [BS71a]. Acoustics [Löw91i]. Acta [FG76]. actinide [CSHM04]. Activities [Löw76c, Jan77]. actual [Car88]. Adapted [Kat04, Pan82a, Pan82b, Pan83, Pan84]. Addendum [Löw81b]. Addition [Bou05, Pan82a]. Address [Löw80f]. Adjoint [FL83, Löw83c, Löw96h, Löwxxd, Sco73]. Advances [Bos83, Gor84, Löw64c, Löw65b, Löw67a, Löw68c, Löw70a, Löw72b, Löw73a, Löw74a, Löw75c, Löw77c, Löw78a, Löw80b, Löw81c, Löw81d, Löw82a, Löw82b, Löw83a, Löw85b, Löw86a, LSZ88, LSZ89, LSZT90, LSZ91, LSZ92a, LSZ92b, LSZ94, LSZ95, LSZ96, LSZ^{+}97b, LSZ^{+}97a, LSZ^{+}98a, LSZ^{+}98c, LSZ^{+}98d, LSZ^{+}98b, LSZB99, LSZ^{+}99b, LSZ^{+}99a, LSZ^{+}00, LSZB00, SB02, SB17, Kön68, Kut67b, Kut67a, Par66]. Advancing [LÖS02]. Affinities [Lin04]. agenda [SG14]. Aging [Löw64e, Löw65h, Löw77d, Löw07b]. aktiviteter [Löw76c]. Akustiken [Löw91i]. Albert [ES87]. Alberte [LÖ79, LS79]. Algebra [BC04, Löw64f, Löw66e, Löw67d, Löw77a, Löw77b, Löw78e, Löw82d, Löw92k, Löw92l, Löw98a, Löwxxb, Löwxxa, SW92, Jon92, Löw81b]. algebras [WS92]. algorithm [Sch77]. Alkali [Löw51e, LCH72, Löw48e]. Alkenes [TTK04]. Alkyl [CPR04]. Allen [Sce06]. Alpha [Jon88, JW89, Ant85, De 76, Duf71, JW78, Jon91, MC09, MÇ11a, MC11b, ST05, Suz84a, Suz84b, Suz85, Suz87, Suz90, Suz92]. Alpha-Function [JW89, Jon91, ST05, Suz84a, Suz84b]. Alpha-Functions [Jon88]. alpha-radial [MC09]. Alternant [LCLL73, Löw54b, PdHL62a, PdHL62b]. alternatives [SM13]. American [Löw76b, Löw77f]. American-Swedish [Löw76b]. Aminopurine [RG76]. Analysis [BJ76, GÖL70, Löw92l, BDMC06, CSHM04, Per12, dA12]. Analytic [Jon91, LA56, TSH^{+}19, JBW87, MRH10]. Analytical [Jon88, Tul04]. analyzed [LAS99]. Angular [Har04, Löw64d, Cro77, MK08, MK09]. annihilation [Pan82b]. Anniversary [LGC78b, LGJLCB89]. Announcement [Löw86b, Löw88b, Löw94b]. Antone [Suz87]. any [FMPM^{+}14]. Aperiodic [Löw62g, Löw65h]. Appendix [SS04]. Application [Cha90, JH15, JW89, Kil69, Löw53b, Löw92q, Löw93g, Löw94f, Löw95j, Löw96f, Löw97c, Löw98b, Löw99b, SL58b, Gra77, PdHL62b]. Applications [DFKS05, LBH04, Löw51e, Löw66e, Löw67d, LCH72, Löw92m, Löw48e, May10, HL83, LFM89a, SL59]. Approach [Laz04, Löw85g, Mor78, SW92, ASW20, May02a, WS92]. Approaches [Lin04, Löw59b, Löw65j, Löw07a, Mei03]. Approximants [ČVV87]. Approximate [Löw53a, Löw85c, Mak81, Mez97]. Approximation [Dog77, LL68, LPLH63, WJ12, Löw55c, Löw88l]. Approximations [DDH75, Löw68l, Löw82e, Löw85e]. April [LSZÖ89, Löw89f]. Arbitrary [Kat04]. Argon [LA56]. Argon-Like [LA56]. Aromatic [VBHG95]. arranged [Löw63c, Löw63d]. Art [Löw86g]. Artificial [Löw86l]. Arvsanlagen [Löw65c]. Aspects [Heh76, Löw58g, Löw60c, Löw62k, Löw64j, Löw64i, Löw65h, Löw68a, Löw68b, Löw68h, Löw68i, Löw69f, Löw70d, Löw75a, Löw75g, Löw76b, Löw77g, Löw80j, Löw84, Löw85g, Löw86j, Löw87b, Löw87c, Löw92r, Löw97d, Löw98c, Löwxxc, Löw07b, Löw07c, Car88, Löw77f, Löw88j, Meh75]. Associated [Löw85d, Löw88e, LFM89a, Löw93h, TL81, TL82]. Associations [Löw68d, Löw81a]. Asymptotic [SPH14, LS51]. Atom [ABD04, BK76, Löw81a, CSHM04]. Atomic [BKNS91, CL62, FL57, GÖL70, KSM87, KS88, Kor79, LBH04, Löw50b, Löw63c, Löw63d, Löw63g, Löw65f, Löw65g, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw75e, LÖ76, LÖ77, LÖ78, LÖ79, Löw80c, LÖ80, LÖ81a, LÖ81b, LÖM83, LSZ84a, LZS84, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, Löw92p, LÖSZ93, Löw94e, Löw95i, Löw96e, Löw97b, LÖSZ98, LÖSZ99, LÖS00, ML83b, Öö83, VBHG95, Löw53c, Löw54d, Löw54e, LSZ^{+}98a, LSZ^{+}00, MRH10, Suz92, TSH^{+}19]. Atoms [Frö76, HL65, Hof67, KD04b, Löw58g, Sla76, Van76, CSHM04, HL59, LR59, Löw66h]. atoms-in-molecules [CSHM04]. att [Löw59c]. attraction [MC09, MÇ10a, MÇ10b, MÇ12]. Aufbau [Sce09a]. Aufl [Kle66b, Kut67b]. August [FH58, Löw58f, LGC78a, LGC78b, LCG82a, Löw89a]. Augustine [LSZÖ89, Löw89f, Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q, LÖSZ93, Löw93g, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00]. av [Löw91i]. Axioms [Löw82d]. Aziridines [JSBL04].

B [Ahl83, Kle66b, Kle66a, MY04]. Back [Ano76a]. Background [Löw78c, Löw83d]. Band [Kön68, Löw62b, MV81, WW76, Löw62a]. Band-Structure [MV81, WW76]. Bände [Rei04a]. Barnett [Bou05, BS06, SPH14]. Base [Abd76, KJL87, Kry02]. Based [LFM89a, MRH10]. Bases [Lat76, KKS67]. Basic [Löw77g, Car88]. Basis [ACL73, KS88, Löw67b, Löw67c, Mak81, MV81, FWMZ11, HL83, JH15, LCK76, Pan83, YE81]. Basis-Sets [Mak81]. Be [Löw94g, Löwxxb, Löwxxa, Löw77b]. Beach [LS78, LÖ78, LÖ79, LS79, LÖ80, LÖ81a, LÖ81b, LÖ82c, LÖ82a, LÖM83, LSZ84a, LZS84, Löw94e, Löw94f, ML83b, Öö83]. Begreppet [Löw91i]. Behaviour [LCH72]. Belgium [Boy04, Boy05]. Belysning [Löw58c]. Benzene [Löw53b]. Bernard [Buc65, Gra65, LÖ79, LS79, Sel65]. Bessel [BS06]. Between [BKNS91, Lin04, Löw65j, Löw84, Löw85e, Löw95g, Löw99a, Par09, AM12, FH58, Löw86l]. Betydelse [Löw91i]. Bi [LFM89a]. Bi-Variational [LFM89a]. bias [FWMZ11]. Bibliographies [SYLxx]. Binary [Löwxxa]. Binding [Lin04, Löw62a, Löw81a, Löw62b]. Bindningen [Löw58c]. binomial [MÇ11a]. Binuclear [VBHG95]. Biochemistry [Löw66j, Löw67k, PP76]. Biographical [Lin03]. Biological [Lad76, Löw63e, Löw65h, LÖSZ87b, Löw07b]. Biology [Kle66b, Kle66a, LP64, Löw64i, Löw68d, Löw68g, Löw69e, Löw69f, Löw71f, Löw71g, Löw72d, Löw73e, Löw74f, Löw75f, LS76, LS77, LS78, LS79, LS80, LS81a, LS81b, LS82b, LS82a, LZ83c, LZ83a, Löw83e, Löw83d, LSZ84b, LSZ84c, LÖ85, LÖSZ85d, LÖ86, LÖSZ86d, LÖSZ87d, LÖSZ87b, Löw88g, LÖSZ88c, Löw89f, Löw90e, Löw90j, Löw91d, Löw91p, Löw92e, Löw92c, Löw92q, Löw93g, Löw94a, Löw94f, Löw95a, Löw95c, Löw95j, Löw96b, Löw96f, Löw97c, Löw98b, Löw99b, Löw91e, Löw91f, Löw91g, Löw92g, Löw92h, Löw92f, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c, Löw93c, Löw93d, Buc65, Gra65, Sel65]. Biomagnetic [Baj75]. Biorthonormal [Lat76]. Birthday [BDL^{+}94, LGBC89]. Bivariational [LFM89b]. Bloch [Löw98c]. Blumenfeld [Baj75]. Blumenfeld-Kalmanson-Löwdin [Baj75]. Body [Har04, Lin04, Löw80c, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, Mic76, AM12]. Bogoliubov [Boy04, Boy05]. Bond [CPR04, Dau76, HNLD63, KSM87, KS88, Löw58c, Löw61b, Löw62a, Löw66j, Löw69f, Löw91k, Löw62b]. Bonded [KSM87]. Bonding [CC04, Gos76]. Bonds [Löw67k, RS76a]. Book [Ahl83, Boy04, Boy05, Buc65, Gra65, Hof67, Kle66b, Kle66a, Kön68, Kor79, Kut67b, Kut67a, Löw69c, McL81, Par66, Rei04a, Rei04b, Sel65]. Boolean [Löw92k]. Borelius [LCH72]. Boston [Ahl83]. both [Löw69c]. Bounds [Ahl76, LL68, Löw65a, Löw65j, Löw65k, Löw65m, Löw66b, LL70, Rei76, Gra77, Löw65l, Löw66c]. Bracketing [Brä76, ST14]. Brändas [Boy04, Boy05, Rei04a, Rei04b]. Bridged [TTK04]. Brief [Löw68d, Löw80e, Löw85a, Löw85i]. Brillouin [Gra77, Löw64k, Löw86d]. Bromine [JP76]. Brueckner [Lin04].

C [Hof67, Löw66d, Löw66h, VBHG95]. C-13 [VBHG95]. Calais [McL81, Löw96a]. Calcul [Löw58b]. Calculated [KSM87]. Calculating [Lin04, Löw54c, Suz90]. Calculation [Har76, JP76, Löw47a, Löw51b, Löw51d, Löw51e, LPdH60, Löw65a, Löw65m, Löw66b, LL70, Löw85c, MK08, MK09, MC09, MÇ10a, MC11b, MÇ12, TTK04, Löw48c, Löw48d, LS51, Löw66c, Löw69a, LL51, MÇ11a, Suz84a, Suz84b, Suz85, Suz90, Suz92, Löw53c, MÇ10b]. Calculations [ABD04, BCA04, DDH75, Dog77, HL76, JSBL04, LBH04, Löw54a, Löw62a, Löw80d, Löw80e, Löw86f, LN89, Löw91h, Mal04, MV81, Sla76, Löw62b, WW76]. Call [Löw88a, LZ91]. can [CSHM04]. Cancer [JL87, Löw62e, Löw62h, Löw77d]. Canonical [Har76, Mor78, JH15, NS04]. Carcinogenesis [Löw68h, Löw77h, Löw86c]. Case [Löw86d]. Cases [Löw92m]. catalysis [LGC78a]. Catalytic [TTK04]. Catalyzed [TTK04]. Causality [Löw61b]. Caused [Löw62e]. Causes [BKNS91]. Cavity [Tul04]. Cellular [Frö76, Sla76]. Center [JW89, Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q, MC11b, Jon84, Jon91, Löw53a, MK08, MK09, MC09, MÇ10a, MÇ10b, MÇ12]. centre [LL51]. Certain [JL87, LL51]. Chairman [Löw89d]. challenge [AK02, Sce06]. Change [Löw85d, Löw88e, LFM89a]. Changes [Löw65c]. Changing [Par03]. Character [Gos76]. characteristic [CD04]. Charge [ABD04, JBW87, MC09, MÇ10a, MÇ10b]. Charges [BKNS91, KSM87, KS88, VBHG95, MRH10]. Chemical [Brä14, Dau76, HNLD63, Löw58c, Löw61b, LGC78a, Löw86f, Löw94g, Nye93, Qua04, RS76b, Sch04, May02a]. Chemistry [Ahl83, And04, BK76, Ber04, BH07, Bos83, Boy04, Boy05, BK04, CL78, COM89, Dav76, DFKS05, FH58, FL07, GS12, Gor84, Het00, Kle66b, Kle66a, Lar04, Löw57, Löw58f, Löw61b, Löw64c, LP64, Löw65b, Löw67a, Löw67e, Löw68c, Löw68i, Löw68l, Löw70a, Löw71a, Löw72b, Löw73a, Löw74a, Löw75c, Löw76a, Löw76c, Löw76d, Löw77c, Löw77d, Löw78a, LC78, Löw80b, Löw80f, Löw80a, Löw80j, Löw81c, Löw81d, LÖ81a, LÖ81b, Löw82a, Löw82b, LÖ82c, LÖM83, Löw83e, LSZ84a, LZS84, Löw85b, LÖSZ85e, LÖSZ85c, Löw86a, LÖSZ86e, LÖSZ86c, Löw86g, LÖSZ87c, Löw87c, LSZ88, Löw88c, Löw88h, LÖSZ88d, Löw88i, LSZ89, Löw89a, LSZÖ89, Löw89d, Löw89b, Löw89h, LSZT90, Löw90b, Löw90e, Löw90h, Löw90g, Löw90k, LSZ91, Löw91d, Löw91l, Löw91q, Löw91r, LSZ92a, LSZ92b, Löw92e, Löw92c, LZ92, Löw92l, Löw93c]. Chemistry [Löw93d, LSZ94, Löw94a, LZ94, LSZ95, Löw95a, Löw95c, LSZ96, Löw96b, LSZ^{+}97b, LSZ^{+}97a, LSZ^{+}98c, LSZ^{+}98d, LSZ^{+}98b, LSZB99, LSZ^{+}99b, LSZ^{+}99a, LSZ^{+}00, LSZB00, Mal04, McL81, ML83b, Nye93, Öö83, ÖSZL87, Par03, Par09, Rei04a, Rei04b, SB02, SB17, Sce09a, ZL94, Brä17, Buc65, Car88, Gra65, LÖS02, LCS^{+}70, LGC78b, LP83, Löw88j, Löw88l, LZ91, Löw91e, Löw91f, Löw91g, Löw92g, Löw92h, Löw92f, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c, LSZ^{+}98a, PL83, Sce08, Sel65, SG14, Löw90e, Löw91d, Löw91e, Löw91f, Löw91g, Löw91r, Löw92g, Löw92h, Löw92f, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c, SG14, Ahl83, Kön68, Kut67b, Kut67a, Par66]. Citation [Löw87a]. Clarke [Löw67g]. Classic [Het00, Löw87a]. Clebsch [SW92, WS92]. Clementi [Löw92d]. Closed [CPR04, PdHL62a, GLA^{+}08]. Closed-Shell [CPR04, PdHL62a, GLA^{+}08]. Closer [BKNS91]. Cluster [Mal04, Uzi04]. Clusters [KJL87, Löw81a, LSZ^{+}97a]. CNF [Dog77]. CNH [Dog77]. Coast [LS80, LS81a, LS81b, LS82b, LS82a, LÖ82a, LZ83c, LZ83a, LSZ84b, LSZ84c]. Code [Löw66j, Löw67k, FMPM^{+}14]. coefficient [JW78]. Coefficients [Löw83b, SW92, MÇ11a, Suz92, WS92]. Cognition [SS17]. Cohesion [Löw66d]. Cohesive [FL62, Löw47a, Löw51d, Löw51e, Löw48c, Löw48d, Löw48e, Löw56b, LCS^{+}70, Low01]. Coleman [ES87, Löw87d]. Collected [Sce08]. Collection [Löw76a]. Collision [Kor79, LÖ77, LÖ78, LÖ79, LÖ80, LÖ81a, LÖ81b, LÖM83, ML83b, Öö83]. Collisions [KD04b, Mic76]. Colloquium [Löw81a, LGC78a]. Combined [LR59]. Comment [Suz87]. Commentary [Sce06]. Comments [ACL73, Cal76, GÖL70, Löw66d, Löw67j, Löw68j, Löw71b, LM72b, LM72a, LSZ83b, Löw85h, Löw86c, Löw86k, Löw86l, Löw89b, Löw96g, ML93, Löw69g, LSZ83a, LSZ83c, LSZ83d, Löw98d]. Committee [Löw47b, Löw48b, Löw49]. common [FWMZ11]. Compact [Ant85, Suz87]. Company [Ahl83]. Comparison [KSM87, Lin04, Löw98c]. completeness [LCK76]. Complex [Löw85d, Löw86k, Löw88e, LFM89a, LFM89b, LAS99]. Complexes [KSM87, TTK04, CSHM04]. Compounds [Löw53b, VBHG95, LAS99]. Compressed [Frö76]. compression [Löw48e]. Computation [ČVV87]. Computational [BH07, DFKS05, Kor79, LÖ77, LÖ78, LÖ79, LÖ80, LÖ81a, LÖ81b, LÖM83, LSZ84a, LZS84, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, LÖSZ87c, Löw87c, Löw88h, LÖSZ88d, Löw90k, Löw91q, Löw92p, LÖSZ93, Löw94e, Löw95i, Löw96e, Löw97b, ML83b, Öö83, ÖSZL87, Par09, Duf71, LSZ^{+}97b, HDK72]. Computations [Löw90e, Löw91d, Löw92e, Löw92c, Löw93c, Löw93d, Löw94a, Löw95a, Löw95c, Löw96b, Löw91e, Löw91f, Löw91g, Löw92g, Löw92h, Löw92f, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c]. Computer [Löw92k, Jon84, Jon92]. Computers [Löw87c, Par03]. Concept [Ber04, Löw91i]. Concepts [FH58, Löw88i, Löw90b]. Concluding [Löw90i, Löw91c]. Condensed [LÖ82c, Löw90e, Löw91d, Löw92e, Löw92c, Löw92p, Löw93c, Löw93d, LÖSZ93, Löw94a, Löw94e, Löw95a, Löw95c, Löw95i, Löw96b, Löw96e, Löw97b, LÖSZ98, LÖSZ99, LÖS00, Löw91e, Löw91f, Löw91g, Löw92g, Löw92h, Löw92f, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c]. Conditions [LCC81, Löw86d]. Conference [Löw47b, Löw48b, Löw49, Löw63g, Löw80f, Löw90i, Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q]. conferences [Meh75]. configuration [JH15]. configuration-interaction [JH15]. Configurational [Löw87a, Löw55b]. Configurations [SL55, SL58b, SL59, AK02, LR59]. Conformation [Lad76]. Congress [Ahl83, Löw80a, LP83, Löw89a, Löw89d, Löw91c, PL83]. Conjugated [Löw53b]. Connected [Löw50b, Löw59a]. Connection [Löw63g, Löw85e, Löw95g, Löw96g, Löw99a, Löw65j]. Conqueror [Ohn76]. Consciousness [Löw86j]. Consistency [Löw80i]. Consistent [Löw51a, LA56, Löw60c, Löw62k, Löw53c, Löw54d, Löw54e]. Constant [JP76]. Constants [LL72, Löw47a, Löw62f, Löw68k, LG71, Löw48c, Löw48d, Löw48e]. Constructed [Löw64d]. Construction [ACL73, Mak81, VB75, Pan82a, Pan82b, Pan83, Pan84]. Contact [Löw64b]. Containing [BCA04, CL62]. Continuum [Löw85c, SL55]. Contraction [Löw64m]. Contributed [Löw80a]. Contributions [Löw47b, Löw48b, Löw49, Löw50a, Löw88a, LZ91]. Conventional [BK76]. Convergence [Löw87a, Löw90b, Löw55b]. Convolution [Löw66e, Löw67d]. Coordinate [Las76, MK08, MK09, MC09, MÇ10a, MÇ10b]. corrected [Mei03, SHL^{+}18]. correction [ASW20, TSH^{+}19]. correction-scaling [ASW20]. Correlated [Har04]. Correlation [GÖL70, Hal76, LB76, Löw55a, Löw56a, Löw59b, Löw68a, Löw68b, Löw71d, Löw72f, Löw72g, Löw75b, Löw95b, Löw07a, Löw07c, Löw07d, SL56, GLA^{+}08, Löw55d, LR59, Löw62d]. Correlations [Lun76]. Correspondence [FH58]. Cosmic [Löw78c]. Coulombic [Löw89c, Löw90i]. Coulson [SPH14, Bou05, BS06]. Coulson/Löwdin [Bou05, BS06]. Coupled [Mal04, Uzi04]. Coupled-Cluster [Mal04]. Coupling [Ber04, JP76, Kil69, Löw64b]. Covalent [Löw58c, Kry02]. creation [Pan82b]. Crick [Baj75]. criterion [LCK76]. Crystalline [VB75]. Crystals [ACL73, Ber76, Cal76, FL62, Löw47a, Löw48a, Löw50b, Löw51d, Löw51e, Löw62a, Löw71d, Löw72f, Löw72g, Löw48c, Löw48d, Löw48e, Löw53a]. Cubic [MV81]. Cure [Kry02]. Current [CL78, Löw59b, Löw77h, Löw91r, Löw94g, Löw07a, Per12, dA12]. curriculum [Ano02]. Cycle [TTK04]. Cyclic [LPdH60, PdHL62b].

D [Ahl83, VB75]. dans [LS54]. Dedicated [Löw65f, Löw65g, LGBC89, OS02]. Deduced [JP76]. Defects [Cal76]. Definition [Löw91j, Löw91i, Löw86e, Sut02]. Definitionen [Löw91i]. Deformed [Van76]. Degeneracies [LG71]. Degeneracy [Löw58h, LG68, LG69]. degenerate [Löw55d]. Del [LSZ^{+}99a]. d’Energie [LS54]. densities [Lev76]. Density [CRE79, Cha90, CSHM04, CLS63, Dav76, ES87, JP76, KSM87, Löw63a, LL70, LCG82b, Löw87a, Löw87b, LSZT90, LSZ^{+}98b, Mez97, Pat62, GLA^{+}08, JBW87, Lev76, Löw55b, Löw86i, MC09, MÇ10a, MÇ10b, SHL^{+}18]. Density-Matrices [KSM87, Löw87a]. Density-Matrix [Cha90].

Density-of-States [JP76]. Deoxyribonucleic [BS71a]. Dependent [BL69, Löw63b, Löw65i, Löw67b, Löw67c, Löw67i, Löw71b, LM72b, LM72a, Löw73b, LSZ^{+}99a, LSZB00]. Deras [Löw65c, Löw91i]. Derivation [LCC81, YE81]. descents [Löw50a]. Described [Löw91r]. design [WJ12]. Determinantal [Wei76]. Determinants [Löw83b]. Determination [Lad76]. Determining [Löw51a]. Deuterium [VBHG95]. Deuterium-Isotope [VBHG95]. Development [BH07, Löw58g, Löw65j, Löw87b, Löw87c, Löw89g, Löw91k, Löw91l, Löw91s, Löw95b, Löw97d, Löwxxc, Meh75]. Development-Evaluation [Löw65j]. Developments [Löw83e, Löw89h, Löw07d]. Devoted [Löw90e, Löw91d, Löw92e, Löw92c, Löw93c, Löw93d, Löw94a, Löw95a, Löw95c, Löw96b, Löw91e, Löw91f, Löw91g, Löw92g, Löw92h, Löw92f, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c]. DFT [AM12]. Diatomic [Dog77, Gos76]. Different [Löw59b, Löw07a]. Differential [McI76, Löw52]. diffraction [Löw50a]. Digital [Par03]. Dilemma [Löw63a]. Diplatinum [TTK04]. Dipole [Löw51b]. Dirac [Löw64b]. Directions [Löw86f]. Discipline [Par03]. Disciplines [Nye93]. Discrepancy [BKNS91]. Discussion [CD04, CLS63, FG76, HNLD63, Löw59b, Löw84, Löw07a, LPLH63, Yos20]. Discussions [Löw89b]. Disease [Löw62e]. Dissociation [CC04, CPR04]. Distance [Löw47a, Löw48c, Löw48d, Löw48e]. DNA [Abd76, Baj75, BS71b, FWMZ11, KKS67, KJL87, Kry02, Löw63e, Löw64e, Löw64j, Löw65h, Löw68h, Löw07b, ML68]. DNK [KKS67]. Do [SS17]. d’Onde [Löw58b]. Dordrecht [Ahl83, Boy04, Boy05]. Double [Baj75]. Dubrovnik [Löw89b]. DV [LSZ^{+}00]. DV-X [LSZ^{+}00]. Dynamic [Laz04, Lun76]. Dynamics [LSZÖ89, Nye93]. Dyson [Lin04].

Early [Par03]. Easier [Suz85]. Edited [Boy04, Boy05, Kle66a, Kut67a, Löw69c, Rei04b, Buc65, Gra65, McL81, Sel65]. Editor [Löw51b, Löw53a, LS55, LCÖ84, Löw89a, SL55, SL58a, WW76]. Editorial [Brä14, LSZ83a, LSZ83b, LSZ83c, LSZ83d, Löw90c, Löw90a, Löw90d, Löw90e, Löw91a, Löw91d, Löw91e, Löw91f, Löw91g, Löw92a, Löw92b, Löw92e, Löw92g, Löw92h, Löw93a, Löw93b, Löw93c, Löw93d, Löw94c, Löw95e, Löw95f, Löw96c, LBG^{+}97]. Editorial-Program [Löw90e]. Educational [Ohn76, Kar02]. Effect [Baj75, KS88, Löw64e, Löw65d]. Effects [ABD04, Ber76, Ber04, CPR04, GL04, JSBL04, Löw55a, Löw71d, Löw72f, Löw72g, Löw75b, Löw55d, Löw62d]. efficient [WJ12]. Egenvardesproblem [Löw59c]. Egil [Löw63c, Löw63d, Löw69c, MTW68]. EHT [CD04]. Eigenfunctions [VB75]. Eigenvalue [Löw59c, Löw62j, Löw67b, Löw67c]. Eigenvalues [Löw65a, Löw65j, Löw65k, Löw65m, Löw66b, LL70, Rei76, Gra77, Löw65l, Löw66c, Löw69a]. Elastic [Löw47a, Löw48c, Löw48d, Löw48e]. Electric [Löw51b, Gra77]. Electrical [Cho76]. Electromagnetic [Laz04, Löw50a]. Electron [Dau76, Koł76, LCLL73, Lin04, LS56, Löw59b, Löw62i, Löw95b, Löw07a, Löw07d, Lun76, MY04, AM12, GLA^{+}08, Jon91, Lev76, Löw61a, NA12, Per12, dA12]. electron-repulsion [Jon91]. Electronic [COM89, LBH04, Löw56a, Löw58g, Löw60b, Löw73b, Löw89c, Löw89b, LSZ^{+}97a, Löw99a, ML68, Mak81, Par03, TTK04, KKS67, LSZ^{+}99a, LSZB00, MRH10]. Electronic-Structure [COM89, Mak81]. Electronique [LS54]. Electrons [Har76]. Elektronnaia [KKS67]. element [Pan84]. Elementary [Kat04, Löw58d, Löw62c, Löw63h]. Elements [GL04, Mal04, Sch04, Uzi04, Löw69g]. embedding [MRH10]. Empirical [Löw99a, LZ91]. Energetics [CC04, RG76]. Energies [FL62, GÖL70, Lin04, MY04, MRH10]. Energy [AJA75, FL58, Löw47a, Löw51d, Löw51e, Löw54c, Löw56a, Löw65a, Löw65j, Löw65k, Löw65m, Löw66b, LL70, Löw73b, Löw88i, Löw89g, LN89, Löw90i, ML68, PdHL62a, Rei76, GLA^{+}08, Gra77, Löw48c, Löw48d, Löw48e, Löw65l, Löw66c, Löw69a, Mei03, LÖT76]. energy-corrected [Mei03]. Enrico [Löw92d]. enthalpies [SHL^{+}18]. Entropy [Löw88i]. Environment [Löw97d]. Equation [BC04, Löw58d, Löw62c, Löw63h, Löw64l, Löw64b, Löw67b, Löw67c, Löw85g, Löw88c, Löw98c]. Equations [Löw64m, McI76, Wei76, Löw52]. equilibrium [GLA^{+}08]. Equivalent [Har76]. Erkki [Boy04, Boy05, Rei04a, Rei04b]. Erratum [MÇ10b, Suz84a]. Errors [LB76, Löw64j]. especially [LL51]. Ethic [Löw85h]. Ethylene [LS54]. EUCHEM [Löw90i]. Eugene [Boy04, Boy05, Rei04a, Rei04b]. Evaluate [Cha90]. evaluating [Suz85, Suz92]. Evaluation [Jon88, Löw65j, Bou05, BS06, Pan84, ST05, Pan83]. Evolution [Löw63b, Löw65i, Löw67i]. Evolutions [Cha90]. EWMO [LÖT76]. Exact [Löw60c, Löw62k, Löw64g, NA12, WW76]. Exchange [ABD04, BJ76, Har76, Löw62e, Löw62d, LG68, LG69, Löw71d, Löw72f, Löw72g, Löw76b, Jon84, Löw77f]. Excited [HL76, KKS67, Löw65l, SL58a]. Expanded [Jon88, JBW87]. Expansion [Dog77, Löw60a, Löw83b, MC11b, Gra77, LD89, MC09, MÇ10a, MÇ10b, MÇ12, Suz92]. Expansions [HNLD63, SPH14, NA12]. Expectation [LB76]. Explanation [Sch04]. Explicit [ČVV87]. Exploring [FWMZ11, TTK04]. Exponentially [Har04]. Expressed [HL65, Löw68i, HL59, Löw88j]. Expression [PdHL62a, Ant85, Suz87]. extended [Löw60a]. Extension [Löw55a, Löw66g, Löw75b, Löw91n, Löw69b, Löw55d]. Extensions [Löw68a, Löw68b, Löw07c]. External [Cho76, KD04b]. Eyring [Löw69e].

factor [LR59]. fast [MÇ11a]. father [Brä17]. FCN [DDH75, Dog77]. February [Löw94e, Löw94f, Löw95i, Löw95j, Löw96e, Löw96f, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00]. Fermi [ASW20, Löw64b, SHL^{+}18, TSH^{+}19]. Fermion [CLS63, LSZT90]. Festschrift [CG77]. Field [Cho76, Löw60c, Löw62k, VB75, Gra77]. Fields [KD04b, Löw51a, LA56, Löw53c, Löw54d, Löw54e]. filling [AK02]. Find [Löw86k]. First [DFKS05, HL76, LA56, Lev76, Löw52, Löw86i]. first-order [Lev76, Löw86i]. five [Löw85i]. Flagler [LS78, LÖ78, LÖ79, LS79, LÖ80, LÖ81a, LÖ81b, LÖ82c, LÖ82a, LÖM83, LSZ84a, LZS84, ML83b, Öö83].

Florida [Löw63c, Löw63d, Löw65f, Löw65g, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw74f, Löw75e, Löw75f, LÖ76, LS76, LÖ77, LS77, LS78, LÖ78, LÖ79, LS79, LS80, LÖ80, LÖ81a, LS81a, LS81b, LÖ81b, LÖ82c, LS82b, LS82a, LÖ82a, LÖM83, LZ83c, LZ83a, LSZ84a, LSZ84b, LZS84, LSZ84c, LÖSZ85e, LÖ85, LÖSZ85c, LÖSZ85d, LÖSZ86e, LÖ86, LÖSZ86c, LÖSZ86d, LÖSZ87d, LÖSZ87c, LÖSZ87b, Löw88h, Löw88g, LÖSZ88d, LÖSZ88c, LSZÖ89, Löw89f, Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q, LÖSZ93, Löw93g, Löw94e, Löw94f, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00, ML83b, Öö83, ÖSZL87, Ano59, Löw72e, Löw76b, Löw77f]. Fock [LPLH63, FL83, HL76, Löw55a, Löw55c, Löw55d, Löw58e, Löw60a, Löw63a, Löw66f, Löw66a, Löw66g, Löw69b, Löw75b, LCC81, LFM89a, LFM89b, LM92, ML93, May02a, Uzi04, Van76]. folding [WW76]. Following [BS07]. Fonctions [Löw58b, Löw58a]. Förändringar [Löw65c]. Forces [Mez97]. Form [ČVV87, Duf71, JW78]. Formalism [Löw62j, Löw64k, Löw85g, MH04]. Format [Löw88b]. formation [SHL^{+}18]. Formed [Tul04]. Formula [Löw51d, JW78, YE81]. Formulae [MC11b, MC09, MÇ10a, MÇ10b]. Formulas [Jon84, Löw53a, MÇ12]. Formulated [Löw82d]. Formulation [Hal76]. Forskargruppen [Löw76c]. Forty [DFKS05]. Foundation [LZ94, ZL94]. Foundations [Löw91s, Löw98d]. Four [Har04, LL51]. Four-Body [Har04]. four-centre [LL51]. Fourth [Ahl83, LP83, PL83]. från [LC78]. Fractional [ASW20]. Free [Löw86j, Löw88i, Löw89g, LN89]. French [Löw58a]. Front [Ano76b]. Fukui [Löw83a]. Fulfilling [May02a]. Full [DDH75]. Fully [Mal04]. Function [Ahl76, Brä76, JP76, JW89, Lin04, MC11b, McI76, Ant85, CSHM04, De 76, Duf71, Jon84, JBW87, Jon91, Jon92, Löw60a, Löw86i, MC09, MÇ10a, MÇ10b, MÇ12, NA12, ST05, Suz84a, Suz84b, Suz85, Suz87, Suz90, Suz92, ST14]. Functional [LCG82b, LSZT90, LSZ^{+}98b, SHL^{+}18]. Functionals [ES87, Löw96h, Löwxxd, Lev76]. Functions [ACL73, CL62, Jon88, Löw50b, LA56, Löw58a, Löw86d, Pau76, SPH14, Bou05, BS06, JW78, Löw53c, MK08, MK09, MÇ11a, Pan82a, Pan82b, Pan83, Pan84, Tch99]. Fundamental [BK04, Löw92q, Löw93g, Löw94f, Löw95j, Löw96f, Löw97c, Löw98b, Löw99b, Rei04a, Boy04, Boy05, Rei04a, Rei04b]. Fundamentals [Löw95g]. Fundaments [Löw64f]. Future [And04, Löw68l, Löw76a, Löw88i, Löw89g, Löw90h, Löw91l, Löw88l]. fuzzy [CSHM04].

Gamma [Löw88k]. Gas [LCLL73]. Gauss [BS06]. geb [Kle66b, Kut67b]. General [Har04, Löw51d, Löw83e, Löw86d, Löw88c, LM92, Löw93f, ML93, Löw85d, Löw88e, PdHL62a]. Generalization [KS73, Löw86d, Ste77, Löw64k]. Generalizations [Löw58e]. Generated [FL57, AM12]. Generator [Las76]. Genetic [Löw64e, Löw65c, Löw66j, Löw67k]. Genetics [Löw62g, Löw63f, Löw65h]. geometry [GLA^{+}08]. Germany [HDK72]. Gets [Ano59]. Getting [NS04].

Girona [Car88, Löw68i, Löw88j]. Giuseppe [LSZ^{+}99a]. Global [Löw97d]. goals [Brä14]. Gordan [SW92, WS92]. Goscinski [McL81]. Got [SS17]. gradients [MRH10, TSH^{+}19]. Gram [NSS12]. Graphs [Löw85c]. Green [McI76]. Green’s [Lin04]. Green’s-Function [Lin04]. Ground [HL76, Löw51d, LR59, Löw65l, Löw65k]. Group [Löw66e, Löw67d, LG68, LG69, Löw76c, PP76, Tch99, LCS^{+}70, LGC78b]. Grunder [Löw91s]. Guseinov [MÇ10b, MK08, MK09, MC09, MÇ10a, MC11b, MÇ12].

H [Löw71g]. Halides [LCH72, Löw48e]. Hamiltonian [CD04, Löw85d, Löw85g, Löw88e, LFM89a, Löw89c, Löw90i, May02a, Uzi04]. Hamiltonians [Löw82c]. Harald [Löw69c]. Harmoni [Löw91s]. Harmonilärans [Löw91s]. Harmony [Löw91m, Löw91s]. Hartree [HL76, LPLH63, Van76, FL83, Löw55a, Löw55c, Löw55d, Löw58e, Löw60a, Löw63a, Löw66f, Löw66a, Löw66g, Löw69b, Löw75b, LCC81, LFM89a, LFM89b, LM92, ML93]. Having [Löw96g]. HCN [DDH75, Dog77]. Heavy [GL04, Mal04, Uzi04]. Hebrew [Löw89a]. Held [FH58, HDK72, Löw71f, Löw72d, Löw73e, LGC78b, Löw80a, Löw81a, LS81b, LÖ81b, LS82a, LÖ82a, LCG82a, LZ83a, LZS84, LSZ84c, LÖSZ85c, LÖSZ86c, LÖSZ86d, LÖSZ87c, LÖSZ87b, LÖSZ88d, LÖSZ88c, ML83b, Öö83, Löw58f, Löw67g, Löw68g, Löw69e, Löw71g, Löw74e, Löw74f, Löw75e, Löw75f, LÖ76, LS76, LÖ77, LS77, LGC78a, LS78, LÖ78, LÖ79, LS79, LS80, LÖ80, LÖ81a, LS81a, LÖ82c, LS82b, LP83, LÖM83, LZ83c, LSZ84a, LSZ84b, LÖSZ85e, LÖ85, LÖSZ86e, LÖ86, LÖSZ87d, Löw88h, Löw88g, LSZÖ89, Löw89f, Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q, LÖSZ93, Löw93g, Löw94e, Löw94f, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00, ÖSZL87, PL83]. Helical [Baj75]. Helium [LS55, SL56, LR59]. Helium-Like [SL56, LR59]. Helmholtz [CD04]. hennes [Löw74c]. Henry [Löw69e]. Herausgeg [Kle66b, Kut67b]. Herausgegeben [Rei04a]. Heredity [Löw65h, Löw07b]. Hermitian [May02a]. Heterocyclics [Löw51b]. High [LCH72, SPH14, Löw48e]. Hilbert [Lat76]. Historic [Löw85a, Löw85i]. Historical [Löw91k, Löw95b]. History [GS12, Löw87c]. Homolytic [CC04]. Honor [Löw63c, Löw63d, Löw76a, Löw91b, Löw92d, Löw67g, Löw69e, Löw71g, LÖ79, LS79, LSZ^{+}99a]. Honorary [Lin93]. Honoring [CG77]. Honour [Löw63g]. Horizons [Ahl83, LP83, PL83]. Hubbard [Lar04]. Hulthén [LS51]. Human [Löw77g, Löw80j, Löw84]. hybrid [Tch99]. Hydrido [TTK04]. Hydrido-Bridged [TTK04]. Hydrogen [BK76, CC04, HL65, HL76, KSM87, Las76, Löw66j, Löw67k, Löw69f, RS76a, HL59]. Hydrogen-Bonded [KSM87]. Hydrosilylation [TTK04]. Hylleraas [Löw63c, Löw63d, Löw63g, MTW68, Löw69c]. Hyperbola [Par03].

ICQC [Löw89d]. Ideas [Löw59b, Löw07a]. Identity [Par03]. II [Rei04b, Roo00, HM01b, Löw48d, Löw51e, Löw54d, Löw54e, Löw55c, Löw63a, Löw64k, LFM89a, LSZ^{+}98d, Löwxxa, ÖL83, Öö83, OS02, Pan82b, PdHL62b, Suz85, WS82]. III [Boy05, BK04, Löw55d, LA56, Löw63a, Löw64l, Pan83, Suz90]. Illustrated [Las76]. Impact [Heh18]. Imperatives [Par09]. implementation [Jon84, Pon19, Yos20]. Implications [Löw63e]. Importance [BJ76, Löw62h, Löw68h, Löw72a, Löw77g, Löw88c, Löw90h, Löw91l, Löw92k, Löw97d, Löwxxc]. impulse [WJ12]. Inception [FL07]. Include [Löw55a, Löw75b, Löw55d]. Incorporation [Löw64j]. Indefinite [Löw96g]. independence [NS04]. Independent [Löw66g, Löw69b, Löw68a, Löw68b, Löw07c]. Independent-Particle [Löw07c]. Independent-Particles [Löw66g]. Individual [RS76a]. Induced [RG76]. Inelastic [Ber76]. Inequalities [Löw88d, Löw93e]. Infinite [BS76]. Infinite-Order [BS76]. Influence [Löw65d, Kar02]. Information [Löw64e]. Inheritance [Löw65c]. inhomogeneous [AM12, GLA^{+}08]. Initio [Löw99a, Sce09a]. Inn [LÖ82a]. Inner [Ahl76, ČVV87, Gos76, Löw68l, Löw71c, Löw70e, Löw88l]. Inner-Projections [Ahl76]. Inner-Shell [Gos76]. input [Lin10]. insight [LSZ^{+}99b]. Instability [LCC81]. Institute [Boy04, Boy05, Löw76c]. Institutionen [Löw76c]. Integral [DDH75, Dog77, JW89, Jon92, MK08, MK09, MC09, MÇ10a, MÇ10b, YE81]. Integrals [CL62, Har04, JP76, Jon88, Löw54a, MC11b, Jon84, Jon91, Löw53a, LL51, MÇ12, Pan83]. Integration [Löw51a, Löw52]. Intelligence [Löw86l]. Intensities [Laz04]. Interaction [HL65, Koł76, Löw87a, ASW20, HL59, JH15, Löw55b, SHL^{+}18, TSH^{+}19]. Interactions [Luq04, BJ76]. Interionic [Löw47a, Löw48c, Löw48d, Löw48e]. Intermediate [Uzi04]. Intermolecular [Löw90i]. International [Ahl83, Car88, HDK72, Kor79, Löw63c, Löw65f, Löw65g, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw74f, Löw75e, Löw75f, LÖ76, LS76, LÖ77, LS77, LS78, LÖ78, LÖ79, LS79, Löw80f, Löw80a, LS80, LÖ80, Löw81a, LÖ81a, LS81a, LS81b, LÖ81b, LÖ82c, LS82b, LS82a, LÖ82a, LP83, LÖM83, LZ83c, LZ83a, LSZ84a, LSZ84b, LZS84, LSZ84c, LÖSZ85e, LÖ85, LÖSZ85c, LÖSZ85d, LÖSZ86e, LÖ86, LÖSZ86c, LÖSZ86d, LÖSZ87d, LÖSZ87c, LÖSZ87b, Löw88h, Löw88g, LÖSZ88d, LÖSZ88c, Löw89a, LSZÖ89, Löw89f, Löw89d, Löw90k, Löw90j, Löw91q, Löw91p, Löw92e, Löw92c, Löw92p, Löw92q, LÖSZ93, Löw93g, Löw94a, Löw94e, Löw94f, Löw95a, Löw95c, Löw95i, Löw95j, Löw96b, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00, ML83b, Öö83, ÖSZL87, PL83]. International [Brä14, Löw63d, Löw63g, Löw76c, Löw90e, Löw91d, Löw91e, Löw91f, Löw91g, Löw91r, Löw92g, Löw92h, Löw92f, Löw93c, Löw93d, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c, SG14]. internationalization [Kar02]. Internationella [Löw76c]. Interpolation [Löw54d, Löw54e]. Interpretations [Löw87a, Löw55b]. Intersecting [Tul04]. Intersection [TL81, TL82]. Introduction [Löw63d, Low69d, Löw71e, Löw71d, Löw72f, Löw78b, LÖ82b, LZ83b, Löw84, LÖSZ85a, LÖSZ85b, LÖSZ86a, LÖSZ86b, LÖSZ87a, LÖSZ88a, LÖSZ88b, Löw90f, Löw91b, Löw92c, LZ92, Löw92i, Löw92j, Löw92d, LZ94, LBÖ^{+}97, ML83a, ML83b, ÖL83, Öö83, OS02, ZL94, Löw65g, LS81b, LÖ81b, LS82a, LÖ82a, LZ83a, LZS84, LSZ84c, LÖSZ85c, LÖSZ85d, LÖSZ86c, LÖSZ86d, LÖSZ87c, LÖSZ87b, LÖSZ88d, LÖSZ88c]. Introductory [Löw72g, Löw80c, Löw89d]. invariance [BDMC06]. Inverse [LPdH60, Ste77]. Inversion [KS76]. Invertibility [WS81]. Investigation [Löw48a, Löw48e]. Investigations [Pau76]. Ionic [Ber76, Cal76, FL62, Löw47a, Löw48a, Löw48c, Löw48d, Löw48e]. Ionization [MY04, SHL^{+}18]. Ions [LA56, SL56]. ISBN [Boy04, Boy05]. Island [Löw63c, Löw63d, Löw63g, Löw65f, Löw65g, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw74f, Löw75e, Löw75f, LÖ76, LS76, LÖ77, LS77]. Isomer [Dog77]. Isotope [JSBL04, Löw65d, VBHG95]. Israel [Löw89a]. Issue [COM89, Lin93, LGBC89, Löw91b, Löw92d, OS02]. Italy [Löw81a]. Iteration [Löw58d, Löw59c, Löw62c, Löw63h]. Iteration-Variation [Löw58d, Löw59c, Löw62c, Löw63h]. Iterations [Löw59c]. Iterations-Variationsmetod [Löw59c]. IV [Löw62j, Pan84, Suz92]. Ivar [LGJLCB89, LGBC89]. IX [Löw65j].

J [Boy04, Boy05, Rei04a, Rei04b, Suz84a, Suz87]. Jahn [Ber04]. January [Löw63c, Löw63d, Löw63g, Löw65f, Löw65g, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw74f, Löw75e, Löw75f, LÖ76, LS76, LÖ77, LS77]. Japan [Löw80a]. Jauch [Löw74b]. Jean [McL81, Löw96a]. Jean-Louis [McL81, Löw96a]. Jerusalem [Löw89a]. John [ES87, Hof67, Löw66d, Löw67g, Löw69c, Löw87d, Löw66h, Löw71g]. Josef [Löw74b]. Josef-Maria [Löw74b]. Journal [Löw90e, Löw91d, Löw91e, Löw91f, Löw91g, Löw91r, Löw92e, Löw92c, Löw92g, Löw92h, Löw92f, Löw93c, Löw93d, Löw94c, Löw94a, Löw95e, Löw95a, Löw95c, Löw95f, Löw95d, Löw96b, Löw96c, SG14]. July [Löw58f, Löw91c]. June [Car88, LP83, Löw90i, PL83, LGBC89, LGJLCB89].

kaleidoscopic [SG14]. Kalmanson [Baj75]. Kasha [Löw91b]. Kashiwagi [Ste77]. Kemiska [Löw58c]. Kenichi [Löw83a]. Key [Löw91r]. Kiev [Boy04, Boy05]. Kinematic [Löw39]. Kinematiska [Löw39]. Kinetic [FL58, JSBL04]. Kinetics [Qua04]. Kluwer [Boy04, Boy05]. Knight [Sce06]. Knut [Löw69c]. Kohn [AM12]. Kovalenta [Löw58c]. Kryachko [Boy04, Boy05, Rei04a, Rei04b]. Kvantgenetisk [Löw65c]. Kvantkemi [Löw77d, LC78]. Kvantkemiska [Löw76c]. Kvantmekanisk [Löw58c]. kvantteorins [Löw74c]. Kyoto [Löw80a, Löw80f].

Laboratory [LÖSZ87b, Löw88h, Löw88g]. Laguerre [Réd76]. Lanczos [HL83]. Lapland [LCG82a]. Large [Löw80i, LÖ82a, Löw87c, Tch99]. Large-Scale [Löw87c]. Law [LCH72]. LCAO [ACL73]. leading [HM01b]. Learning [SS17]. Least [Löw92l]. lectures [GC84]. LEDO [Dog77]. Legacy [Bun17]. Leon [Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00]. Letter [Löw51b, Löw53a, LS55, SL55, SL58a, WW76]. Leuven [Boy04, Boy05]. Levdinovskii [KKS67]. Levels [Löw54c, Löw56a]. LiBr [JP76]. Lie [BC04]. Life [Löw84, Löw85c, SB02]. Life-Times [Löw85c]. Light [Löw58c]. Like [LA56, SL56, LR59]. Limited [Dog77]. limits [Löw69a]. Lindgren [Löw98c]. Linear [Löw64f, Löw77b, Löw92l, Löw93f, Löw96g, Löw96h, Löw98a, Löwxxd, Löwxxb, Löwxxa, McI76, TL81, NS04, TL82]. Linearly [BL69, Löw67b, Löw67c]. Liouville [Löw88c]. Liouvillian [Löw85g]. Liouvillians [Löw82c]. liquid [GLA^{+}08]. liquids [AM12]. Literature [Man76]. Ljunggren [Löw50a]. Local [Pan82b]. Localization [Lar04, CRE79]. Localized [RS76b, Löw64a]. Logic [Löw95g]. London [Ahl83, Kle66b, Kön68, Kut67b]. Long [HL59, HL65, Koł76, Löw89c, Löw90i, VBHG95]. Long-Range [HL65, Koł76, VBHG95, HL59]. Look [BKNS91]. Lorentz [Löw39]. Lorentz-Transformationen [Löw39]. Lösa [Löw59c]. Lost [Heh18]. Louis [McL81, Löw96a]. Low [Löw78c, MY04]. Low-Lying [MY04]. Löwdin [Ahl83, Buc65, Gra65, Kle66b, Kle66a, Kön68, Kut67b, Lin03, McL81, Rei04a, Roo00, Sch77, Sel65, Suz84a, FMPM^{+}14, KKS67, AJA75, AEG80, AK02, AM12, Ano59, Ano02, Ant85, ASW20, Baj75, BS71a, BS71b, BKNS91, BM86, BS07, Bos83, Bou05, BS06, Brä76, BK04, Brä17, BDMC06, Bun17, CGLÖ76a, CGLÖ76b, CG77, COM89, CD04, Cha90, ČVV87, CSHM04, Cro77, De 76, DDH75, Dog77, Duf71, FG76, FWMZ11, Gor84, GC84, GLA^{+}08, Gra77, Hal04, HM01a, HM01b, Heh18, HL83, Hof67, Jan77, JH15, JW78, Jon84, JBW87, Jon88, JW89, Jon91, Jon92, KSM87, KS88, Kar02, KS73, Kil69, Kor79, Kry02, Kut67a, LBH04, LCK76, LAS99, LD89, Lin93, LÖS02, Lin10, LÖBS17, Löw76a, Mak81, MK08, MK09, MC09, MÇ10a, MÇ10b, MÇ11a]. Löwdin [MC11b, MÇ12, Man76, MV81, May02a, May02b, May10, MRH10, Mei03, Mez97, Mor78, NSS12, NA12, NS04, Ohn76, ÖSB01, OS02, Pan82a, Pan82b, Pan83, Pan84, Par66, Per12, Pon19, RSS08, SB02, SB17, Sce06, Sce09a, SHL^{+}18, Sco73, SPH14, SG14, SW92, ST05, Ste77, Sut02, Suz84b, Suz85, Suz87, Suz90, Suz92, ST14, SM13, Tch99, TSH^{+}19, VB75, VBHG95, WJ12, WW76, WS81, WS82, Yos20, dA12, WS92, Boy04, Boy05, Rei04a, Rei04b]. Löwdin- [MÇ12]. Löwdin-alpha [MÇ11a]. Löwdin’s [SS17]. Lower [LL68, Löw65a, Löw65m, Löw66b, LL70, Rei76, Löw66c, Löw69a, Gra77, Löw65k, Löw65l]. Lying [MY04].

Macromolecular [Mez97]. Macromolecules [Lad76]. Madelung [AK02]. Magnetic [JL87, KD04b]. Man [Löw74c]. Manageability [Par09]. Manifolds [TL81, LAS99, TL82]. Mäniskan [Löw74c]. Many [Dau76, Lin04, Löw59b, Löw62i, Löw80c, Löw85d, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, Löw87a, Löw88e, LFM89a, LFM89b, Löw07a, Löw07d, Lun76, Mic76, Löw53a, Löw55b, Löw55c, Löw55d, Löw63i, LSZT90, Roo00]. Many-Body [Lin04, Löw80c, LÖSZ85c, LÖSZ86c, Mic76]. many-center [Löw53a]. Many-Electron [Dau76, Löw59b, Löw62i, Löw07a, Löw07d, Lun76]. many-fermion [LSZT90]. Many-Particle [Löw85d, Löw87a, Löw88e, LFM89a, LFM89b, Löw55b, Löw55c, Löw55d, Löw63i, Roo00]. March [FHL55, LS78, LÖ78, LÖ79, LS79, LS80, LÖ80, LÖ81a, LS81a, LS81b, LÖ81b, LÖ82c, LS82b, LS82a, LÖ82a, LÖM83, LZ83c, LZ83a, LSZ84a, LSZ84b, LZS84, LSZ84c, LÖSZ85e, LÖ85, LÖSZ85c, LÖSZ85d, LÖSZ86e, LÖ86, LÖSZ86c, LÖSZ86d, LÖSZ87d, LÖSZ87c, LÖSZ87b, Löw88h, Löw88g, LÖSZ88d, LÖSZ88c, Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q, LÖSZ93, Löw93g, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, Löw99b, LÖS00, ML83b, Öö83, ÖSZL87]. Maria [Löw74b]. Marine [LÖSZ87b, Löw88h, Löw88g]. Marine-Biological [LÖSZ87b]. Marineland [LÖSZ85e, LÖ85, LÖSZ85c, LÖSZ85d, LÖSZ86e, LÖ86, LÖSZ86c, LÖSZ86d, LÖSZ87d, LÖSZ87c, LÖSZ87b, Löw88h, Löw88g, LÖSZ88d, LÖSZ88c, ÖSZL87]. mark [LGC78b]. Marriott [Löw94e]. Materials [Löw86f, LSZ^{+}00]. Math [Suz84a, Suz87]. Mathematical [Löw77b, Löw90b, Löw90g, Löwxxb, Löwxxa, Löw86e, Yos20]. Mathematician [Löw87d]. Mathematics [Löw90g]. Matrices [CLS63, Dav76, ES87, KSM87, LL70, Löw87a, Löw87b, Mez97, CD04, Lev76, Löw55b, Löw86i]. Matrix [Cha90, LBH04, LPdH60, Löw63a, Pat62, CRE79, JW78, JBW87, May02a, Sco73, Pan84]. Matter [Ano76a, Ano76b, COM89, LÖ82c, Löw86l, Löw90e, Löw91d, Löw92e, Löw92c, Löw92p, Löw93c, Löw93d, LÖSZ93, Löw94a, Löw94e, Löw95a, Löw95c, Löw95i, Löw96b, Löw96e, Löw97b, LÖSZ98, LÖSZ99, LÖS00, Nye93, Löw91e, Löw91f, Löw91g, Löw92g, Löw92h, Löw92f, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c]. Maximum [LÖT76]. May [Löw81a]. Mayer [NSS12]. MCSCF [RS76b]. Meaning [Löw91i]. Means [Löw65a, Löw65m, Löw66b, LG71, Löw87a, Löw55b, Löw66c]. Measures [Mez97]. Mechanical [BCA04, Cal76, FL58, Löw47a, Löw51d, Löw51e, Löw58c, Löw86d, Löw48c, Löw48d, Löw48e, Löw51c]. Mechanics [KD04a, Löw59b, Löw59a, Löw62f, Löw63c, Löw63d, Löw63g, Löw66e, Löw67d, Löw72a, Löw07a]. Mechanism [BS71b, Löw68h, RG76, BS71a, KKS67, Kry02]. Meeting [FH58]. mekhanizm [KKS67]. melting [LGC78b, McL81]. membranes [LGC78a]. Memoir [Lin03]. Memoriam [Löw74b]. Memory [Boy04, Boy05, BK04, Löw96a, Rei04a, Rei04b, SS17]. Mesomeric [Löw58a]. Mésomerie [Löw58a]. Metals [LA56, Löw66d, Löw51e]. Method [CL62, LCLL73, Las76, Löw54b, Löw55a, Löw59c, Löw66f, Löw66a, Löw66g, Löw68k, Löw75b, Löw85d, Löw86k, Löw87a, Löw88a, Löw88e, LFM89a, Löw91k, Löw92l, LM92, MY04, ML93, Pan82a, Pan82b, Pan83, Pan84, PdHL62a, PdHL62b, Van76, CRE79, JH15, Jon84, Jon91, Löw50a, Löw55b, Löw69b, May02b, MRH10, Pon19]. Methods [BK76, Luq04, BJ76, Dog77, LBH04, LB76, Löw51a, LÖ77, LÖ78, LÖ79, LÖ80, LÖ82c, Löw85e, Löw85g, LÖSZ87c, Löw88h, LÖSZ88d, Löw90k, Löw91q, LZ92, Löw92p, LÖSZ93, Löw94e, Löw95i, Löw96e, Löw97b, Löw99a, ÖSZL87, Sla76, TTK04, Uzi04, ASW20, CGLÖ76b, LR59, LZ91, ST05, Tch99, Kor79]. Metric [LBH04, Löw96g]. Michael [Löw91b]. Microelectronics [Löw89g]. Microwave [Löw78c]. Midtdal [Löw69c]. Mind [Löw80j, Löw84, Löw86j, Löw86l]. Minimum [AJA75]. mixing [LAS99]. MM [Tch99]. Mobilities [RS76a]. Mobility [Löw73b]. Model [Frö76, Tul04, FWMZ11, Löw69b, Löw68a, Löw68b, Löw07c]. Modeling [VBHG95]. Models [Löw77h, Löw86c]. Modern [Löw61b, Löw74c, Löw75a, Löw75g, Löw80j, Löw92r, Löw97d, LSZ^{+}98a, Löwxxc]. moderna [Löw74c]. modified [JW78]. Moléculaire [Löw58b]. Molecular [DDH75, Dog77, Har76, Jon88, JW89, Kor79, LCLL73, Löw53b, Löw54a, Löw54b, Löw54c, Löw56a, Löw62e, Löw63c, Löw63d, Löw63g, LP64, Löw64g, Löw65f, Löw65g, Löw67g, Löw68d, Löw68g, Löw69e, Löw69f, Löw71f, Löw71g, Löw71a, Löw72d, Löw73e, Löw74e, Löw75e, LÖ76, LÖ77, LÖ78, LÖ79, Löw80d, Löw80e, LÖ80, Löw81a, LÖ81a, LÖ81b, Löw83a, LÖM83, Löw83e, Löw83d, LSZ84a, LZS84, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, LSZÖ89, Löw91h, Löw91j, Löw92p, LÖSZ93, Löw94e, Löw95i, Löw95k, Löw96e, Löw97b, LÖSZ98, Löw99a, LÖSZ99, LÖS00, Mic76, ML83b, Öö83, PdHL62a, PdHL62b, FMPM^{+}14, FWMZ11, JBW87, Löw62d, Löw86e, MK08, MK09, MC09, MÇ10a, MÇ10b, Tch99, Buc65, Gra65, Kle66b, Kle66a, Sel65]. Molecular-Orbital [DDH75, Dog77, Löw53b]. Molecule [HL76, JP76, Las76, Laz04, Löw65h, Löw07b, Löw86e, Sut02]. Molecules [BCA04, FHL55, Gos76, Hof67, KS88, KD04b, Löw50b, Löw51d, Löw51e, LFH55, Löw60b, Löw62h, Löw80i, Löw89c, Löw89b, Mak81, Mal04, Sla76, CSHM04, GLA^{+}08, Löw53a, Löw66h, LL51]. Moments [Löw51b]. Momentum [KS76, Löw64d, Löw73b, Cro77]. Monovalent [Löw66d]. Moral [Löw85h]. Motion [LL72, Löw62f, Löw68k, LG71, Löw85g, Löw91j]. Mott [Lar04]. Mountains [Ohn76]. MR [Suz87]. Mulliken [Löw65f, Löw65g, BKNS91, CSHM04, KSM87, KS88, LP64, Kle66b, Kle66a]. Multi [JW89]. Multi-Center [JW89]. Multicenter [Jon88]. Multiconfigurational [MY04]. Multidimensional [LG71]. Mutagenesis [Abd76]. mutasii [KKS67]. Mutation [Löw65d, FWMZ11, KKS67]. Mutations [Löw64e, Löw65h, Löw07b, RG76, Kry02].

N [MY04, MY04]. Något [Löw91s]. National [Löw47b, Löw48b, Löw49]. Natural [HNLD63, HL65, LS55, LS56, Löw70d, Löw84, Löw86l, Löw87a, Löw95g, Löw97d, Löwxxc, RS76b, SL58b, SL59, AM12, CSHM04, HL59, Lev76, Löw55b, NA12]. Nature [Löw58a, Löw67e, Löw81a, Löw58a]. near [JBW87]. Neither [GS12]. Netherlands [Boy05]. News [FMPM^{+}14]. Niveaux [LS54]. NMR [VBHG95]. NMR-Spectra [VBHG95]. no [Suz87]. Nobel [Löw78c, Löw80c]. Non [Del76, Löw50b, Löw68e, Löw91k]. Non-Orthogonality [Del76, Löw50b, Löw68e, Löw91k]. Nonadditivity [MH04]. Noninteger [MC11b, MÇ12]. Nonmetallic [Löw62a]. nonorthogonal [YE81]. Nonorthogonality [Löw70b]. nor [GS12]. Normal [Löw62f, Löw72a]. Norway [Löw69c]. Note [Löw61a, RSS08, Löw50a, LS51, Löw51c, Sco73, LCÖ84]. November [Löw80a]. NSO’s [Löw63a]. NTH [Löw69c]. Nuclear [JP76, Löw91j, MC09, MÇ10a, MÇ10b, MÇ12]. Nuclei [Laz04]. Numbers [MC11b, ASW20, MÇ12]. Numerical [BS76, Heh76, Löw51a, Löw51e, ST05, Bou05, Löw48e, LS51, Löw52].

O [Ahl83, Kle66b, Kle66a, Kön68, Kut67b, Kut67a, Roo00, CPR04, Sut02]. O. [Bos83, CG77, Gor84, Jan77, SG14]. Obituary [Löw73c]. Objectivity [Löw75a, Löw75g, Löw92r]. observations [Sut02]. Occasion [BDL^{+}94]. occupation [ASW20]. Occurrence [Löw85d, Löw88e, LFM89a]. occurring [LL51]. October [HDK72, Lin03, Löw80a]. Öhrn [BDL^{+}94, LSZ^{+}99b]. Olov [Boy04, Boy05, Buc65, Gra65, Hof67, Kor79, McL81, Par66, Rei04a, Rei04b, Sel65, Ano02, BK04, Brä17, Bun17, CGLÖ76a, CGLÖ76b, COM89, FG76, Hal04, Heh18, Kar02, Lin93, LÖS02, Lin03, Lin10, LÖBS17, Löw76a, Man76, Ohn76, ÖSB01, OS02, SB02]. Om [Löw91i, Löw91s]. One [MC11b, AM12, MC09, MÇ12]. one-body [AM12]. One-Center [MC11b, MC09, MÇ12]. ONIOM [MRH10]. ONIOM-based [MRH10]. Open [CPR04]. Open- [CPR04]. Opening [Löw80f]. Operator [BL69, Löw62j, Löw63b, Löw65i, Löw67i, Löw93e, VB75, WS81, Yos20]. Operators [FL57, FL83, Kil69, Löw64d, Löw72a, Löw82c, Löw83c, LFM89b, Löw93f, Löw96h, Löwxxd, Löwxxa, Cro77, Löw63i, Pan82a, Pan82b, WS82]. Optimal [Mak81]. Orbit [Kil69, Löw64b]. Orbital [ABD04, DDH75, Dog77, LB76, LCLL73, LAS99, Löw53b, Löw54c, Löw65f, Löw65g, PdHL62a, PdHL62b, AK02, ASW20, CRE79, FMPM^{+}14, JH15, MK08, MK09, MC09, SHL^{+}18, Suz92, TSH^{+}19]. Orbitals [Gos76, Har76, HL65, Jon88, JW89, Kle66b, Kle66a, Lin04, Löw54b, LS55, LS56, LP64, Löw64g, Löw87a, MC11b, RS76b, SL58b, SL59, AM12, Bou05, Buc65, Gra65, HL59, Jon91, Lev76, Löw55b, MÇ10a, MÇ10b, MÇ12, NA12, Sel65]. Order [BS76, LL68, LÖ82a, SPH14, Lev76, Löw52, Löw86i]. Orders [KSM87, KS88]. Ordinary [McI76, Löw52, Löw55c]. Organic [Löw53b, Löw89h]. Origin [Löw84, Löw91m, JBW87, Kry02, AK02]. Orthogonal [LD89, WJ12]. Orthogonality [Del76, Löw50b, Löw68e, Löw91k]. Orthogonalization [AJA75, Del76, KS73, Mak81, RSS08, AEG80, BM86, JH15, LAS99, May02b, NSS12, NS04, Sco73, Ste77, SM13, WJ12]. Orthogonalizations [SS17]. Orthonormal [ACL73]. Orthonormalization [Löw93h, Mor78, Sch77]. Osäkerhetsrelationerna [Löw91i]. osnovanii [KKS67]. Osvaldo [McL81]. Other [Löw92l, Jan77]. Our [Löw97d]. Overlap [ABD04, Ber76, BJ76, Dog77, HM01a, HM01b, JP76, LÖT76, LPdH60, MC11b, Jon92, MK08, MK09, Pan83]. overlapping [WJ12].

P [Ahl83, Kle66b, Kle66a, Kön68, Kut67b, Kut67a, Roo00]. P-O [Roo00]. P. [Bos83, CG77, Gor84, Jan77, SG14, Sut02]. P.-O [Ahl83, Kle66b, Kle66a, Sut02]. P.-O. [Bos83, CG77, Gor84, Jan77, SG14]. pages [Ahl83]. Pair [FL83, Löw83c, KKS67]. Pairing [Löw91n, HM01a, HM01b, Kry02, May10]. Pairs [Abd76, Dog77, KJL87, Kry02, Löw61a]. Palm [LS80, LS81a, LS81b, LS82b, LS82a, LÖ82a, LZ83c, LZ83a, LSZ84b, LSZ84c]. Panel [FH58, Löw84]. Paper [Löw66d, Löw88a, LZ92, LZ94, ZL94, Löw50a, LZ91]. Papers [Löw80a, MTW68, Het00, Löw69c, Sce08, Sce09b]. Parallel [Löw87c]. Parameter [Har76, Löw64l, SW92, CD04, WS92]. Pariser [Löw88a]. Parr [Löw88a]. Part [FH58, ML83b, Öö83, LSZ^{+}98c, LSZ^{+}98d, Löw51d, Löw51e, Löw85d, Löw88e, LFM89a, Löwxxa, ML83a, ÖL83, OS02]. Partial [DDH75, Dog77]. Participants [Löw76a]. Particle [Löw68a, Löw68b, Löw85d, Löw87a, Löw88e, LFM89a, MH04, FMPM^{+}14, Löw55b, Löw55c, Löw55d, Löw63i, Löw69b, LFM89b, Roo00, Löw07c]. Particles [Löw66g, Löw50a]. Particular [Löw66j, Löw67k]. Particularly [FL62, Löw76b, Löw86f, Löw77f]. partition [Tch99]. Partitioning [BS07, Löw63h, Löw65a, Löw65m, Löw66b, Löw68k, LG71, Löw82e, Löw85e, Löw98c, SS04, Wei76, HL83, Löw66c, Löw69a, WS81, WS82]. Past [And04]. Per-Olov [Boy04, Boy05, Buc65, Gra65, Hof67, Kor79, McL81, Par66, Rei04a, Rei04b, Sel65, Ano02, BK04, Brä17, Bun17, CGLÖ76a, CGLÖ76b, COM89, FG76, Hal04, Heh18, Kar02, Lin93, LÖS02, Lin03, Lin10, LÖBS17, Löw76a, Man76, Ohn76, ÖSB01, OS02, SB02]. Periodic [Sch04, AK02, Löw69g, Sce09b]. permutation [Pan83]. Permutational [Kat04]. Perspective [Roo00]. Perturbation [AJA75, BS76, BJ76, LL68, Löw60c, Löw62j, Löw62k, Löw63h, Löw64k, Löw64l, Löw64m, Löw64a, Löw64b, Löw65a, Löw65j, Löw65k, Löw65m, Löw66b, Löw67j, Löw68j, Löw68k, Löw69a, LG71, LÖ82a, Löw82e, Löw85e, LN89, MV81, MH04, SS04, Gra77, Löw51c, Löw65l, Löw66c]. Perturbative [Lin04]. Perturbed [Löw53b]. pg [MK09]. Pharmacology [Löw74f, Löw75f, LS76, LS77, LS78, LS79, LS80, LS81a, LS81b, LS82b, LS82a, LZ83c, LZ83a, LSZ84b, LSZ84c, LÖ85, LÖSZ85d, LÖ86, LÖSZ86d, LÖSZ87d, LÖSZ87b, Löw88g, LÖSZ88c, Löw89f, Löw90j, Löw91p, Löw92q, Löw93g, Löw94f, Löw95j, Löw96f, Löw97c, Löw98b, Löw99b]. Phases [Löw81a, LS51]. Phenomena [Kor79, Löw63b, Löw65i, Löw67i, LÖ77, LÖ78, LÖ79, LÖ80, LÖ81a, LÖ81b, LÖM83, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, ML83b, Öö83]. Phenomenon [LG68, LG69]. Philosopher [Sce09a]. Philosophy [Nye93, Sce08]. Phosphate [PP76]. Photodetachment [MY04]. Photons [Ber76]. Phys [Löw87a, Roo00, Suz84a, Suz87]. Physical [Löw86k, Löw87a, Löw88c, Sch04, Löw55b]. Physics [Boy04, Boy05, GS12, HDK72, Kle66b, Kle66a, LP64, Löw71a, Löw78c, Löw83e, Löw90e, Löw90k, Löw91d, Löw91q, Löw92e, Löw92c, Löw93c, Löw93d, Löw94a, Löw94g, Löw95a, Löw95c, Löw96b, Buc65, Gra65, LGC78a, Löw91e, Löw91f, Löw91g, Löw92g, Löw92h, Löw92f, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c, LSZ^{+}98a, LSZ^{+}98c, LSZ^{+}98d, Löw98d, Meh75, Sel65, Brä14, Löw47b, Löw48b, Löw49]. PLA [Dog77]. Plane [ML68]. plicies [Löw95e]. point [Kry02]. Polar [KS88]. Policies [Löw90e, Löw91d, Löw91e, Löw91f, Löw92b, Löw92e, Löw93c, Löw93d, Löw91g, Löw92g, Löw92h, Löw94c, Löw95f, Löw96c]. polynomial [LD89]. Polynomials [Réd76]. Ponce [Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00]. Pons [Yos20]. Ponte [Löw94e, Löw94f]. Pople [Löw88a]. population [BDMC06, CSHM04, Per12, dA12]. positive [Sco73]. Positrons [BCA04]. Possible [Löw68a, Löw68b, Löw68h, Löw07c]. pot [LGC78b, McL81]. Potential [BGC76, AM12]. Potentials [MY04, SHL^{+}18]. Power [Sla76]. pp [Boy04, Boy05]. PPM [WJ12]. practical [Pon19]. Practice [LC78, Par03]. praktik [LC78]. prediction [May02a]. Preface [CG77, Lin93, Löw64h, Löw65e, Löw67f, Löw68f, Löw70c, Löw72c, Löw73d, Löw74d, Löw75d, Löw77e, Löw78d, Löw79, Löw80g, Löw80h, Löw81e, Löw81f, Löw81g, Löw82f, Löw82g, LCG82c, LCG83, Löw85f, Löw86h, Löw88f, LGBC89, Löw89e, Löw91o, Löw92n, Löw92o, Löw94d, Löw95h, Löw96d, Löw81a, Löw89a]. Preis [Ahl83, Kön68]. Present [And04, Löw57, Löw68i, Löw89d, Löw89b, Löw88j]. Press [Kle66b, Kön68, Kut67b, Löw69c]. Pressure [LCH72]. pressures [Löw48e]. Price [Löw69c]. Principal [MC11b, MÇ12]. Principle [LL72, Löw39, Löw61b, Löw71b, LM72b, LM72a, Löw86d, LFM89a, Sce09a, LS51]. Principles [Löw83e, LZ94, ZL94]. prizes [Löw78c]. Probability [Per12, dA12]. Problem [Dau76, Del76, Hal76, JW89, Löw58h, Löw59b, Löw59c, Löw59a, Löw62j, Löw67b, Löw67c, LG68, Löw68e, Löw68a, Löw68b, LG69, LL70, Löw70b, Löw80i, Löw83c, Löw87b, Löw91k, Löw95b, Löw97a, Löw07a, Löw07c, Löw07d, SL58b, SL59, SS04, Lev76, Löw50b, Pan84, Ste77]. Problems [Löw54e, Löw64e, Löw65h, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, Löw87a, Löw91r, Löw92q, Löw93g, Löw94f, Löw94g, Löw95j, Löw96f, Löw97c, Löw98b, Löw99b, Löw07b, Löw54d, Löw55b]. Procedure [Löw58d, Löw62c, Löw63h, LAS99, Suz85, WW76, Suz90, Suz92]. Procedures [Del76, Löw93h]. Proceedings [Ahl83, FH58, HDK72, Löw63c, Löw63d, Löw65f, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw74f, Löw75e, Löw75f, LÖ76, LS76, LÖ77, LS77, LS78, LÖ78, LÖ79, LS79, LS80, LÖ80, LÖ81a, LS81a, LÖ82c, LS82b, LÖM83, LZ83c, LSZ84a, LSZ84b, LÖSZ85e, LÖ85, LÖSZ86e, LÖ86, LÖSZ87d, Löw88h, Löw88g, LSZÖ89, Löw89f, Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q, LÖSZ93, Löw93g, Löw94e, Löw94f, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00, ÖSZL87, Car88, LGC78a, LP83, PL83, ES87, Löw65g, Löw80a, Löw81a, LS81b, LÖ81b, LS82a, LÖ82a, LCG82a, LZ83a, LZS84, LSZ84c, LÖSZ85c, LÖSZ85d, LÖSZ86c, LÖSZ86d, LÖSZ87c, LÖSZ87b, LÖSZ88d, LÖSZ88c, Löw89a, ML83b, Öö83, Kor79]. Process [Löw70d]. Processes [Löw73b, LSZ^{+}99a, LSZB00]. Product [Löwxxa]. Prof [Löw63g]. Professor [Löw63c, Löw63d, Löw65f, Löw65g, Löw66d, LGJLCB89]. Program [Löw67h, Löw90e, Löw91d, Löw91e, Löw91f, Löw92b, Löw92e, Löw93c, Löw93d, Löw91g, Löw92g, Löw92h, Löw94c, Löw95e, Löw95f, Löw96c]. Progress [Löw62i, Löw95k]. Project [Löw76b, Löw77f, Lin10]. Projected [Löw66f, Löw66a, Löw66g, LCC81, Löw83b, Pau76, Löw69b, HL76, Van76]. Projection [BL69, ČVV87, FL57, Kil69, Löw62f, Löw62j, Löw64d, VB75, Cro77, Mei03, Pon19, Yos20]. Projections [Ahl76, Löw68l, Löw71c, Löw70e, Löw88l]. Projector [TL81, TL82]. Projectors [Cha90]. Projects [Löw72e]. Propagating [LSZ^{+}99b]. Propagator [LÖ82c, Löw85g, MY04]. Propagators [Löw94h]. Properties [Frö76, HNLD63, JL87, LL72, Löw48a, Löw63a, Löw66j, Löw67j, Löw67k, Löw68j, Löw71c, LG71, LFM89b, Löw93f, Löw96g, Löw96h, Löwxxd, Pau76, RSS08, Löw48e, Löw56b, LCS^{+}70, Löw70e, Low01]. property [De 76]. protein [Per12, dA12]. Proton [Abd76, Cho76, KJL87, Löw62e, Löw63e, Löw64e, Löw64j, RS76a, Kry02]. Pseudoquadrupole [JP76]. Psi [Löw88k]. Psyche [Löw74c]. psyke [Löw74c]. PT [Sch04]. Publications [CGLÖ76a, Löw66i]. Published [Löw69c]. Publishers [Boy04, Boy05]. Publishing [Ahl83]. Pullman [Kle66b, LÖ79, LS79, Buc65, Gra65, Kle66a, Sel65]. Pullmann [Ahl83].

QC [Tch99]. QC/MM [Tch99]. QED [GL04]. QM [Luq04, MRH10]. quadrature [BS06]. Quantization [McI76]. Quantum [Ahl83, BK76, Bos83, Boy04, Boy05, BK04, BCA04, CGLÖ76b, Cal76, CL78, COM89, Car88, Dav76, FHL55, FL58, FH58, FL07, GS12, Gor84, Het00, KD04a, Kön68, Kut67b, Kut67a, Lar04, Löw47a, Löw51d, Löw51e, Löw55b, Löw55c, Löw55d, LFH55, LS56, Löw56b, Löw57, Löw58c, Löw58f, Löw59b, Löw59a, Löw60b, Löw61b, Löw62f, Löw62g, Löw62h, Löw62a, Löw62i, Löw63c, Löw63d, Löw63f, Löw63b, Löw63g, Löw64c, Löw64f, Löw64i, Löw65c, Löw65b, Löw65f, Löw65g, Löw65h, Löw65i, Löw66e, Löw66h, Löw67a, Löw67d, Löw67e, Löw67i, Löw68c, Löw68g, Löw68i, Löw68l, Löw69e, Löw70a, Löw71f, Löw71g, Löw71a, Löw72b, Löw72d, Löw72e, Löw72a, Löw73a, Löw73e, Löw74a, Löw74c, Löw74e, Löw74f, Löw75c, Löw75e, Löw75f, Löw76a, LÖ76, LS76, Löw76d, Löw76b, Löw77c, LS77, Löw77a, Löw77b, Löw78a, LS78]. Quantum [LGC78b, Löw78e, LÖ79, LS79, Löw80b, Löw80f, Löw80a, LS80, LÖ80, Löw80j, Löw81c, Löw81d, Löw81b, LÖ81a, LS81a, LS81b, LÖ81b, Löw82a, Löw82b, Löw82d, LÖ82c, LS82b, LS82a, Löw83a, LÖM83, LZ83c, LZ83a, LSZ84a, LSZ84b, LZS84, LSZ84c, Löw85b, LÖSZ85e, LÖ85, LÖSZ85c, LÖSZ85d, Löw86a, Löw86d, Löw86f, LÖSZ86e, LÖ86, LÖSZ86c, LÖSZ86d, Löw86g, Löw86j, LÖSZ87d, LÖSZ87c, LÖSZ87b, Löw87c, LSZ88, Löw88c, Löw88h, Löw88g, LÖSZ88d, LÖSZ88c, Löw88i, LSZ89, Löw89a, LSZÖ89, Löw89f, Löw89d, Löw89g, Löw89b, LN89, LSZT90, Löw90e, Löw90h, Löw90k, Löw90j, LSZ91, Löw91d, Löw91e, Löw91f, Löw91g, Löw91q, Löw91p, Löw91r, LSZ92a, LSZ92b, Löw92e, Löw92c, Löw92g, Löw92h, Löw92f, LZ92, Löw92l, Löw93c, Löw93d, Löw93e, LSZ94, Löw94c, Löw94a, LZ94, LSZ95, Löw95e, Löw95a, Löw95c, Löw95f]. Quantum [Löw95d, Löw95k, LSZ96, Löw96b, Löw96c, LSZ^{+}97b, LSZ^{+}97a, Löw97a, LSZ^{+}98c, LSZ^{+}98d, LSZ^{+}98b, Löw98a, LSZB99, LSZ^{+}99b, LSZ^{+}99a, Löw99a, Löwxxb, LSZ^{+}00, LSZB00, Low01, Löw07a, Löw07b, Löw07d, Mal04, MC11b, Mez97, ML83b, Öö83, ÖSZL87, Par03, Par09, PP76, Qua04, Rei04a, Rei04b, Roo00, SB02, SB17, Sce09a, SG14, SW92, ZL94, Brä17, LÖS02, Löw48c, Löw48d, Löw48e, Löw51c, Löw63i, LCS^{+}70, Löw77f, LP83, Löw88j, Löw88l, LZ91, LSZ^{+}98a, MÇ12, Par66, PL83, WS92, WS81, WS82, Lin10, Löw76c, Löw77d, LC78, LGC78b, Löw87a, Hof67, McL81]. Quantum-Chemistry [Bos83, COM89, Gor84]. Quantum-Mechanical [Cal76, Löw51d, Löw51e, Löw58c, Löw86d, Löw51c]. Quasi [MH04]. Quasi-Particle [MH04].

R [Kle66b, Kle66a, AM12, CPR04]. R. [LP64]. Radial [MC11b, MK08, MK09, MC09, MÇ10a, MÇ10b, MÇ11a, MÇ12]. Radiation [Löw78c]. radio [WJ12]. Range [HL65, Koł76, VBHG95, HL59]. Rates [Löw65d]. Rational [ČVV87, Löw68l, Löw82e, Löw85e, Löw88l]. Re [LSZ^{+}99a]. Reach [Löw88k]. Reaction [RS76b, Löw63i]. Reactions [CPR04, RS76b]. Reading [Löw68h]. Real [Löw85c]. Reality [Löw75a, Löw75g, Löw92r]. recurrence [JW78]. recurrent [Pan82a, Pan82b, Pan83, Pan84]. Reduced [Dav76, LL70, Löw87b, Löw86i]. Reduction [Löw86l]. Reference [Löw66j, Löw67k]. Reformulation [Ber04]. Regression [Löw92l]. Reidel [Ahl83]. Relation [AM12, Löw83e, Löw84, Löw86l, NSS12]. Relations [Löw59a, Löw91i, Löw92l, JW78]. Relativistic [Mal04, Uzi04]. Relativitetsprincipen [Löw39]. Relativity [Löw39]. remark [Löw50a]. Remarks [JL87, Löw63g, Löw72g, Löw80c, Löw86i, Löw89d, LN89, Löw90i, Löw91c, Löw93h, Löw94h, Sce09a, TL81, TL82]. Reminiscences [Réd76]. Renormalized [ČVV87, NA12]. Replication [BS71b, Löw64j, ML68, BS71a]. Report [FHL55, Löw47b, Löw48b, Löw49, LFH55, Löw58f]. Reports [Löw66i]. Representability [LL70, Löw87b, Löw97a, Lev76, Löw86i]. representation [Pan82b]. Representations [Pat62]. repulsion [HM01a, HM01b, Jon91]. Research [Löw62h, LCS^{+}70, Löw70d, Löw72e, Löw76a, Löw76c, Löw77g, Löw85h]. Resemblance [Löw93h]. Resolvent [Löw68k, Löw85e]. Resolvents [Löw94h]. Resonance [Löw85c, Löw85d, Löw86k, Löw88e, LFM89a, YE81]. Resonant [BGC76]. Resort [LÖ82a, Löw94e, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00]. Response [Sce06]. Responsibilities [Löw85h]. restriction [NS04]. Results [BS07]. Rev [Löw87a, Roo00]. Reversal [Qua04]. Review [Ahl83, Boy04, Boy05, Buc65, Gra65, Hof67, Kle66b, Kle66a, Kön68, Kor79, Kut67b, Kut67a, Löw69c, Löw80e, Löw91h, Löw07a, McL81, Par66, Rei04a, Rei04b, Sel65, Löw59b]. Reviews [SG14]. Revisited [FG76, ST14]. Revival [FWMZ11]. rigid [Gra77]. Ring [Löw53b]. Robert [Löw65f, Löw65g]. Rocky [Ohn76]. Role [Löw88i, Löw89g, LZ94, SL55, ZL94]. root [Sco73]. rotation [MK08, MK09]. rotation-angular [MK08, MK09]. rotational [BDMC06]. Rotations [Löw92m]. rotator [Gra77]. round [NS04]. Row [LA56]. Rule [Ahl76, AK02]. Russian [KKS67].

S [Boy04, Boy05, Kle66b, Kle66a, Kut67b, LP64, Rei04a, Rei04b]. Sanderson [Löw65f, Löw65g]. Sanibel [Löw63c, Löw63d, Löw63g, Löw65f, Löw65g, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw74f, Löw75e, Löw75f, LÖ76, LS76, LÖ77, LS77, Löw83b, Löw85i, Löw85a]. Sasaki [Ste77]. Sawgrass [Löw94e]. Scale [Löw87c]. Scaling [Löw59a, Löw85d, Löw86k, Löw88e, LFM89a, ASW20]. Scattering [ABD04, Ber76, BGC76, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, HL83]. SCF [KSM87, Löw64g]. Scheme [FL83, Löw63a, LFM89a, LFM89b, Löw55d, Löw58e, NSS12]. Schemes [Löw66g, LCC81, RSS08, Löw60a]. Schmidt [NSS12]. School [Löw58f]. Schrödinger [BC04, Löw58d, Löw62c, Löw63h, Löw64l, Löw65j]. Science [Löw75a, Löw75g, Löw92r, CGLÖ76b, LSZ^{+}00]. Sciences [Löw70d, Löw76b, Löw83a, Löw84, Löw86l, Löw92k, Löw92l, Löw95g, Löw97d, Löwxxc, Löw77f]. Scientific [Bun17, FG76, Het00, Lin10, Löw85a, Man76, MTW68, Ohn76, Par03, Jan77, Kar02, Löw69c, LGC78b, Löw85i, McL81]. Scientists [Löw85h]. Sciten [Kön68]. SCRF [Luq04]. Search [Löw92l]. Second [LL68, Löw91c]. Second-Order [LL68]. Secular [Löw64m, Löw67b, Löw67c]. Selected [MTW68, Sce09b, Löw69c]. Self [Löw51a, LA56, Löw60c, Löw62k, ASW20, Löw53c, Löw54d, Löw54e, SHL^{+}18, Sco73, TSH^{+}19]. self-adjoint [Sco73]. Self-Consistent [Löw51a, LA56, Löw60c, Löw62k, Löw53c, Löw54d, Löw54e]. self-interaction [ASW20, SHL^{+}18, TSH^{+}19]. Semi [Löw99a, LZ91]. Semi-Empirical [Löw99a, LZ91]. Semiconductors [MV81]. Semiempirical [LZ92]. separability [Löw61a]. Separation [Löw64b]. September [LGC78b]. Series [JSBL04, NA12, Sla76, JBW87]. Set [ACL73, BL69, KS88, Löw77b, Löwxxb, Löwxxa, Boy04, JH15, YE81]. Sets [Mak81, LCK76]. several [SM13]. Sham [AM12]. Shapiro [WS92, SW92]. Sharma [JW78]. Shell [CPR04, Gos76, PdHL62a, GLA^{+}08]. Sheraton [LÖ82a]. Shieldings [Laz04]. Shifts [VBHG95]. Short [Löw76a]. Silico} [And04]. Similar [Löw88i]. Similarity [Löw85d, Löw88e, LFM89a, Mez97]. Simple [BS07, CL62, LCC81, Suz84a, Suz84b, Suz85, Suz90, Suz92]. Simplifications [Löw54c]. simplified [CD04]. since [Meh75]. Situation [Löw57, Löw68i, Löw89d, Löw89b, Löw88j]. Size [Löw80i]. Size-Consistency [Löw80i]. Slater [Löw67g, MÇ10b, AM12, Bou05, JH15, Jon88, JW89, Jon91, Löw53c, Löw66d, Löw66h, Löw83b, MK08, MK09, MC09, MÇ10a, MC11b, MÇ12, Suz92, Hof67]. Slater-Type [Jon88, JW89, JH15, Jon91, Suz92]. Society [Brä14, Löw97d, Löwxxc]. Sodium [Löw51e]. Software [FMPM^{+}14]. Solid [HDK72, Hof67, Kor79, Löw62g, Löw65h, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw75e, LÖ76, LÖ77, LÖ78, LÖ79, LÖ80, LÖ81a, LÖ81b, LÖM83, LSZ84a, LZS84, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, LÖSZ87c, Löw88h, LÖSZ88d, LSZÖ89, Löw90k, Löw91q, LZ92, ML83b, Öö83, ÖSZL87, Löw62d, Löw66h, LZ91]. Solid-State [Kor79, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw75e, LÖ76, LÖ77, LÖ78, LÖ79, LÖ80, LÖ81a, LÖ81b, LÖM83, LSZ84a, LZS84, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, LÖSZ87c, Löw88h, LÖSZ88d, LSZÖ89, Löw90k, Löw91q, LZ92, ML83b, Öö83, ÖSZL87, Löw62d, LZ91]. Solids [FL62, Löw71a, LCH72, Sla76, Löw56b, LCS^{+}70, Low01]. Solute [Luq04]. Solution [CC04, Lev76, Löw62j, Löw64l]. Solvay [Meh75]. Solved [Löw94g]. Solvent [Luq04]. Solving [Löw58d, Löw59c, Löw62c, Löw63h]. Some [ACL73, BS07, Cal76, JL87, KSM87, KS88, Koł76, Löw47a, Löw48a, Löw51b, Löw58g, Löw59b, Löw62i, Löw64j, Löw64i, Löw66j, Löw67j, Löw67k, Löw68a, Löw68b, Löw68h, Löw68i, Löw68j, Löw69f, Löw69g, Löw70d, Löw70e, Löw71b, Löw71c, LM72b, LM72a, Löw75a, Löw75g, Löw76b, Löw77f, Löw77g, Löw77h, Löw77b, Löw80j, Löw83e, Löw84, Löw85g, Löw85h, Löw86c, Löw86j, Löw86k, Löw86l, Löw87b, Löw87c, Löw88c, Löw88j, Löw88d, Löw89b, LFM89b, Löw89h, LN89, Löw91r, Löw92m, Löw92r, LM92, Löw93f, Löw93h, Löw94g, Löw94h, Löw95k, Löw96g, Löw96h, Löw97d, Löw98c, Löw98d, Löwxxc, Löwxxd, Löwxxb, Löwxxa, Löw07a, Löw07b, Löw07c, Löw07d, ML93, Sut02, TL81, TL82, GLA^{+}08, Löw48c, Löw48d, Löw48e, May10, Löw60c, Löw62k, Löw65h]. Something [Löw91s]. Son [Heh18]. Sorrento [Löw81a]. sostoianiakh [KKS67]. Space [BL69, KS76, Lat76, Uzi04]. Spaces [Löw96g, Löwxxa]. Spain [Car88]. Spanned [BL69]. Special [COM89, Löw85g, Löw92m, Löw92d, OS02, Suz85]. Specifics [Bun17]. Spectra [Löw85d, Löw88e, LFM89a, Löw95k, Löw99a, VBHG95]. Spectral [LBH04]. spectrally [WJ12]. Spectroscopies [Laz04]. spectroscopy [LSZ^{+}00]. Spheres [Tul04]. Spherical [Frö76, Löw50a]. Spin [HL65, Kat04, Kil69, LS55, Löw58h, Löw64b, LG68, LG69, Löw83b, Löw87a, Löw92m, Löw96g, MY04, Pau76, SL58b, SL59, HL59, Lev76, Löw55b, Löw62d, Pan82a, Pan82b, Pan83, Pan84, Pon19, Yos20]. spin-adapted [Pan82a, Pan82b, Pan83, Pan84]. Spin-Orbit [Kil69, Löw64b]. Spin-Orbitals [HL65, LS55, Löw87a, HL59, Lev76, Löw55b]. Spin-Projected [Löw83b]. Spins [Kat04]. Splitting [SL56]. sponsored [LGC78b]. spontaneous [Kry02]. Square [Löw92l, Sco73]. St [LSZÖ89, Löw89f, Löw90k, Löw90j, Löw91q, Löw91p, Löw92p, Löw92q, LÖSZ93, Löw93g, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00]. Stability [Löw66j, Löw67k, Löw72a, Löw83c, Löw86d]. Stabilization [Löw85c]. State [FL57, HDK72, Hof67, Kor79, Löw51d, Löw65k, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw75e, LÖ76, LÖ77, LÖ78, LÖ79, LÖ80, LÖ81a, LÖ81b, LÖM83, LSZ84a, LZS84, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, Löw86g, LÖSZ87c, Löw88h, LÖSZ88d, LSZÖ89, Löw90k, Löw91q, LZ92, ML83b, Öö83, ÖSZL87, Löw62d, Löw65l, Löw66h, LZ91]. statement [Brä14]. States [HL76, JP76, Kat04, Löw85c, Löw85d, Löw86k, Löw88e, LFM89a, HL83, KKS67, LR59, Löw65l, SL58a, BH07]. Static [Lun76]. Statistics [Löw74e, Löw75e, LÖ76, LÖ79, LÖ80, Löw88c]. steepest [Löw50a]. STO [JW78]. Stockholm [LFH55, FHL55]. Strong [VB75]. Structure [COM89, CLS63, LBH04, Löw58g, Löw60b, Löw68i, Löw80d, Löw80e, Löw89c, Löw89b, Löw91h, Löw91j, Mak81, MV81, PdHL62a, TTK04, CGLÖ76b, CD04, KKS67, Löw86e, Löw88j, LSZ^{+}97a, LSZ^{+}99a, LSZB00, WW76]. Structures [Baj75]. struktura [KKS67]. Studies [KJL87, Löw53c, Löw54d, Löw54e, LA56, Löw60c, Löw62j, Löw62k, Löw63h, Löw64k, Löw64l, Löw64m, Löw64a, Löw64b, Löw65j, Löw65k, Löw65l, Löw65m, Löw66b, Löw68k, LG71, LM92, PdHL62a, PdHL62b, RS76b, WS81, WS82, SS04]. Study [Baj75, BS71b, JL87, Löw64b, LCH72, Uzi04, BS71a, LCK76, LS51, Löw55c]. Sub [Löw62e]. Sub-Molecular [Löw62e]. Substituent [CPR04]. Substituted [JSBL04]. subtraction [Pan82a]. successively [Suz90]. Sum [Ahl76]. Sum-Rule [Ahl76]. Summaries [Löw76a]. Summary [Löw68d]. Summer [Löw58f]. Super [Löw67b, Löw67c]. Supercomputers [Löw87c]. Superconductivity [Lar04]. Superheavy [Mal04]. Superoperators [Löw82c]. Superposition [SL55, SL58b, SL59, LR59]. Supersymmetric [KD04a]. Surface [ABD04, Löw81a, Tul04]. Surfaces [Löw90i, LGC78a]. Survey [Löw85a, Löw85i]. Sweden [Boy05, FH58, Löw58f, LCG82a, LP83, PL83]. Swedish [Löw39, Löw47b, Löw48b, Löw49, Löw58c, Löw59c, Löw65c, Löw74c, Löw76c, Löw76b, Löw77d, Löw77f, LC78, Löw91i, Löw91s]. Symbolic [ČVV87]. Symmetric [LG68, LG69, LFM89b, May02b, Sch77]. symmetrical [NSS12]. Symmetry [De 76, Har04, KS76, Kat04, Löw63a, Löw67j, Löw68j, LG71, Qua04, RSS08, Pan83, LL72]. Symposia [Löw85i, Löw85a]. Symposium [ES87, FHL55, HDK72, Kor79, Löw63c, Löw63d, Löw65f, Löw65g, Löw67g, Löw68g, Löw69e, Löw71f, Löw71g, Löw72d, Löw73e, Löw74e, Löw74f, Löw75e, Löw75f, Löw76a, LÖ76, LS76, LÖ77, LS77, LS78, LÖ78, LÖ79, LS79, Löw80c, LS80, LÖ80, LÖ81a, LS81a, LS81b, LÖ81b, LÖ82c, LS82b, LS82a, LÖM83, LZ83c, LZ83a, LSZ84a, LSZ84b, LZS84, LSZ84c, LÖSZ85e, LÖ85, LÖSZ85c, LÖSZ85d, LÖSZ86e, LÖ86, LÖSZ86c, LÖSZ86d, LÖSZ87d, LÖSZ87c, LÖSZ87b, Löw88a, Löw88h, Löw88g, LÖSZ88d, LÖSZ88c, LSZÖ89, Löw89f, Löw90k, Löw90j, Löw91q, Löw91p, LZ92, Löw92p, Löw92q, LÖSZ93, Löw93g, LZ94, Löw94e, Löw94f, Löw95i, Löw95j, Löw96e, Löw96f, Löw97b, Löw97c, LÖSZ98, Löw98b, LÖSZ99, Löw99b, LÖS00, ML83b, Öö83, ÖSZL87, ZL94, LGC78b, LZ91, LFH55]. Synpunkt [Löw65c]. System [Löw91m, PdHL62a, Löw69g, MK08, MK09, MC09, MÇ10a, MÇ10b].

Systems [CPR04, Har04, Koł76, LS56, LPdH60, Löw62i, Löw87a, LN89, Löw91s, Löw07d, Lun76, McI76, PdHL62b, Löw55b, Löw55c, Löw55d, Löw63i, LSZT90, LSZ^{+}98c, LSZ^{+}98d, Roo00, Tch99, Löw80c].

T. [Löw50a]. Tab [Kle66b]. Table [Sch04, AK02, Sce09b]. Tame [Löw87d]. Tarfala [LCG82a]. Taylor [JBW87]. Technique [Baj75, BS07, Löw62f, Löw63b, Löw63h, Löw65a, Löw65i, Löw67i, Löw68k, Löw82e, Löw85e, Löw98c, MV81, Wei76, Löw69a, Mei03, WW76, WS81, WS82]. Techniques [Löw65m, Löw66b, Löw66c]. tell [CSHM04]. Teller [Ber04]. Temperature [Löw78c]. Tensor [MY04]. teori [LC78]. term [HM01b]. Terms [HL65, Löw64b, HL59, NA12]. Test [BK76, LCLL73]. Thalberg [Löw69c]. Their [Löw91i, Löw91r, Löw97d, Löwxxc, Löwxxa]. Theorem [FL62, Löw59a, Löw72a, Löw86d, Löw91n, Bou05, HM01a, HM01b, Löw61a, May10, SL58a]. Theorems [Löw93h, Löw60a]. Theoretical [BS71b, Ber04, Boy04, Boy05, Brä14, JSBL04, Lad76, Löw48a, LCH72, Löw83e, Löw83d, Löw89h, Löw94g, Nye93, BS71a, Löw48e]. Theories [Mak81]. Theory [BS76, DFKS05, FHL55, Heh76, Hof67, Kil69, Kor79, Lin10, LL68, Löw50b, Löw53b, Löw54c, LFH55, Löw56a, LS56, Löw58g, Löw60b, Löw60c, Löw62h, Löw62a, Löw62j, Löw62k, Löw62i, Löw63b, Löw63h, Löw64f, Löw64g, Löw64k, Löw64l, Löw64m, Löw64a, Löw64b, Löw65a, Löw65f, Löw65g, Löw65h, Löw65i, Löw65j, Löw65k, Löw65m, Löw66b, Löw67g, Löw67i, Löw67j, Löw68g, Löw68j, Löw68k, Löw69e, Löw71f, Löw71g, LG71, Löw72d, Löw72e, Löw73e, Löw74c, Löw74e, Löw75e, LÖ76, LÖ77, Löw77a, Löw77b, LC78, LÖ78, Löw78e, LÖ79, Löw80c, LÖ80, LÖ81a, LÖ81b, Löw82d, LCG82b, LÖ82c, LÖ82a, Löw82e, LÖM83, LSZ84a, LZS84, Löw85e, Löw85d, LÖSZ85e, LÖSZ85c, LÖSZ86e, LÖSZ86c, Löw86j, Löw87a, LÖSZ87c, Löw87b, Löw88e, Löw88h, LÖSZ88d, LSZÖ89, Löw89g, Löw90e, Löw91d, Löw91l, Löw91m, Löw92e]. Theory [Löw92c, LZ92, Löw92p, Löw92q, Löw92m, Löw93c, Löw93d, Löw93e, LÖSZ93, Löw93g, Löw94a, Löw94e, Löw94f, Löw95a, Löw95c, Löw95i, Löw95j, Löw95k, Löw96b, Löw96e, Löw96f, Löw96g, Löw97b, Löw97c, Löw97a, Löw98a, LÖSZ98, Löw98b, Löw99a, LÖSZ99, Löw99b, Löwxxb, Löwxxa, LÖS00, Löw07b, Löw07d, Mic76, ML83b, Öö83, ÖSZL87, Löw50a, Löw51c, Löw53a, Löw55b, Löw55c, Löw55d, Löw56b, Löw62b, Löw62d, Löw63i, Löw65l, Löw66c, Löw66h, Löw81b, LSZT90, LZ91, Löw91e, Löw91f, Löw91g, Löw92g, Löw92h, Löw92f, Löw94c, Löw95e, Löw95f, Löw95d, Löw96c, LSZ^{+}98b, Low01, LL51, Roo00, SHL^{+}18, WS81, WS82, SS04]. Thermodynamical [Löw88d]. three [Jon84, LL51, MÇ10a, MÇ10b]. three-center [Jon84, MÇ10a, MÇ10b]. three-centre [LL51]. Tight [Löw62a, Löw62b]. Tight-Binding [Löw62a, Löw62b]. till [LC78]. Time [Löw63b, Löw65i, Löw67i, Löw71b, LM72b, LM72a, Löw73b, Qua04, LSZ^{+}99a, LSZB00]. Time-Dependent [Löw63b, Löw65i, Löw67i, Löw71b, LM72b, LM72a, Löw73b, LSZ^{+}99a, LSZB00]. Times [Löw85c]. Ton [Löw91i]. Tone [Löw91i, Löw91m, Löw91s, Löw91s]. Toner [Löw91s]. Tonsystemets [Löw91s]. tool [BS06]. Tools [Löw77b, Löw93e, Löwxxb, Löwxxa]. Toronto [Löw91c]. total [Löw60a]. Trace [Löw77a, Löw78e, Löw82d, Löw81b]. Transactinide [Mal04]. Transfer [Cho76, Löw64j, Löw73b, Per12, dA12]. Transfers [KJL87]. Transform [Mez97]. Transformation [FL58, Löw39, CD04, LD89]. Transformationen [Löw39]. Transformations [Löw85d, Löw88e, LFM89a, Löw92m, Löw96g]. transformed [CD04]. Transition [LA56]. Treated [Löw62f, Löw63b, Löw65i, Löw67i]. Treating [CL62, MH04]. Treatment [BC04, Cal76, LL72, Löw67j, Löw68j, Löw71d, Löw72f, Löw72g, Tul04, JBW87, Löw48e, Löw51e, Löw68k, LG71]. Trends [CL78, Löw88c, Car88, LSZ^{+}98a]. Trial [Löw86d]. Tribute [Boy04, Boy05, BK04, COM89, Hof67, Kle66b, Kle66a, LP64, Löw83a, Löw87d, Rei04a, Rei04b, CGLÖ76b, Löw66h, LSZ^{+}99b, SB02]. Tromsø [Löw90i]. Trondheim [Löw69c]. Tumors [Löw64e, Löw65h]. Tumours [Löw07b]. Tunneling [Löw63e, Löw64e, Löw65d, Kry02]. Tunnelling [Abd76]. Twenty [Löw85i]. Twenty-five [Löw85i]. Two [Har76, HL65, Koł76, LS56, MC11b, HL59, Jon91, MK08, MK09, MC09, MÇ10a, MÇ10b, MÇ12, NA12, TL82, Tul04, WS92]. Two-Center [MC11b, Jon91, MK08, MK09, MC09, MÇ10a, MÇ10b, MÇ12]. Two-Electron [Koł76, LS56, NA12]. Type [Baj75, Jon88, JW89, MC11b, Bou05, JH15, Jon91, MK08, MK09, MC09, MÇ10a, MÇ10b, MÇ12, Suz92].

u [Kle66b]. Ukraine [Boy05]. Unbounded [Löw85d, Löw88e, LFM89a]. Uncertainty [Löw91i]. Understanding [CPR04]. Unified [Laz04]. Unitary [Löw92m, Löw96g]. United [BH07]. Universal [Lev76, FWMZ11]. Universitet [Löw76c]. University [Boy04, Boy05, Löw63c, Löw63d, LGC78a, Löw88h, Löw88g, Löw89a, Löw76c, Löw76d, LGC78b]. Unreachable [Löw88k]. Unsöld [LL68]. Updates [FMPM^{+}14]. Upper [LL68, Löw65a, Löw65j, Löw65m, Löw66b, Löw66c, Löw69a]. Uppkomst [Löw91s]. Uppsala [Boy04, Boy05, FHL55, LFH55, LGC78a, LP83, PL83, FL07, Löw54a, LCS^{+}70, Löw72e, Löw76c, Löw76d, Löw76b, Löw77f, LGC78b, Réd76]. Uppsala-Florida [Löw76b]. Ur [Löw65c]. Use [DDH75, Dog77, Löw50b, Löw68l, MÇ11a, MC11b, LR59, Löw88l]. Used [Löw51a, Löw77b, Löwxxb, Löwxxa]. Using [Baj75, Har76, JP76, Jon88, MY04, TTK04, LAS99, MK08, MK09, MC09, MÇ10a, MÇ10b, MÇ12, MRH10]. UWB [WJ12].

V [AM12, KKS67, Löw60c, Löw62k]. Vålådalen [Löw58f, FH58]. Vålådalensia [FG76]. Valence [Löw62a, Löw91k, Löw62b]. Valencies [KSM87, KS88]. Valley [LCG82a]. Values [LB76, Suz85]. vanishing [Suz85]. variable [CD04]. Variation [LL72, Löw58d, Löw59c, Löw62c, Löw63h, Löw64l, Löw71b, LM72b, LM72a, Löw86d, LN89, Mor78, SL58a, Löw63a]. Variational [LL72, LFM89a, Löw97a, Lev76, LS51]. Variationsmetod [Löw59c]. Various [Löw65j, Löw93h]. världsbild [Löw74c]. Vedra [Löw94e, Löw94f]. VI [Löw64m]. VIA [Laz04, Kry02]. Vibrational [Laz04, Ste77]. Vibronic [Ber04]. vid [Löw76c, Löw91i]. View [Löw65h, Löw74c, Löw80j, Löw87c, Löw89b, Löw07b, Sce09a]. Viewpoint [Löw65c]. VII [Löw64a]. VIII [Kle66b, Löw64b]. Virial [FL62, Löw59a]. vitae [Ano02]. Viver [Yos20]. Vleck [Löw71g]. Vol [Bos83, Gor84, Kut67b, Kut67a, MK09]. Vols [Rei04b]. Volume [Boy04, Boy05, LSZ^{+}99a, SB17, BK04]. Volumes [Boy04, Löw69c]. vozbuzhdennykh [KKS67].

Waller [LGBC89, LGJLCB89]. Water [JL87, KJL87]. WATOC [Löw91c]. Watson [Baj75]. Wave [BC04, Löw50b, LA56, Löw63i, Löw86d, CSHM04, Löw60a, Löw86i, NA12, Pan82a, Pan82b, Pan83, Pan84]. wave-function [Löw86i]. Wavefunctions [FL57, LL72, Löw64d]. waves [Löw50a]. Way [Löw89c, Löw90i]. Weak [BJ76]. Weighted [LÖT76, Mak81, BM86]. Weights [Kat04]. Welfare [Löw77g]. Wergeland [Löw69c]. Weyl [Heh76]. Whitney [LÖSZ87b, Löw88h, Löw88g]. Wigner [Löw64k]. Wildbad [HDK72]. Will [Löw86j]. without [BDMC06]. Wolfsberg [CD04]. Words [Löw91r]. work [SB02]. Workshop [Car88, LÖ82a, LCG82a, Löw89b, Löw88j, Löw68i]. World [Boy04, Boy05, BK04, Löw74c, Löw91c, Rei04a, Rei04b].

X [Löw65k, LSZ^{+}00]. XI [Kut67b, Löw65l]. XII [Löw65m, Löw66b]. XIII [Löw68k]. XIV [LG71]. xlvi [Boy05]. xxxii [Boy04].

Years [DFKS05, Löw85a, Par03, Löw85i, LGJLCB89]. Yngve [BDL^{+}94, LSZ^{+}99b]. York [Kle66b, Kön68, Kut67b]. York-London [Kle66b, Kut67b].

zahlr [Kle66b].

References

Abdulnur:1976:PTD

[Abd76] Suheil F. Abdulnur. Proton tunnelling in DNA base pairs and mutagenesis. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 567–575. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

           QC174.125 .Q371. URL http://link.springer.com/content/
           pdf/10.1007/978-1-4757-1659-7_42.pdf.

Amos:2004:EOO

[ABD04] A. T. Amos, B. L. Burrows, and S. G. Davison. Effects of orbital overlap on calculations of charge exchange in atom–surface scattering. In Brändas and Kryachko [BK04], pages 295–321. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http://link.

springer.com/book/10.1007/978-94-017-0448-9.

Ahlenius:1973:SCC

[ACL73] Tor Ahlenius, Jean-Louis Calais, and Per-Olov Löwdin. Some comments on the construction of an orthonormal set of LCAO basis functions for crystals. Journal of Physics C: Solid State Physics, 6(11):1896–1908, June 11, 1973. CODEN JPSOAW. ISSN 0022-3719.

Aiken:1980:LO

[AEG80] John G. Aiken, John A. Erdos, and Jerome A. Goldstein. On Löwdin orthogonalization. International Journal of Quantum Chemistry, 18(4):1101–1108, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ahlberg:1976:BSR

[Ahl76] Ragnar Ahlberg. Bounds to the sum-rule function from innerprojections. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 349–355. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_22/.

Ahlrichs:1983:BRB

[Ahl83] R. Ahlrichs. Book review: New Horizons of Quantum Chemistry. Proceedings of the Fourth International Congress of Quantum Chemistry, by P.-O. Löwdin, B. Pullmann. D. Reidel Publishing Company, Dordrecht, Boston, London 1983. 457 pages, Preis: $67.50. Berichte der Bunsen-Gesellschaft für Physikalische Chemie, 1963–1990, 87(9):836–837, September 1983. CO-DEN BBPCAX. ISSN 0005-9021. URL http://onlinelibrary.

wiley.com/doi/10.1002/bbpc.19830870928/abstract.

Aiken:1975:LOM

[AJA75] J. G. Aiken, H. B. Jonassen, and H. S. Aldrich. Löwdin orthogonalization as a minimum energy perturbation. Journal of Chemical Physics, 62(7):2745–2746, ???? 1975. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Allen:2002:LCO

[AK02] Leland C. Allen and Eugene T. Knight. The Löwdin challenge: Origin of the n+ℓ, n (Madelung) rule for filling the orbital configurations of the periodic table. International Journal of Quantum Chemistry, 90(1):80–88, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Amovilli:2012:RBS

[AM12] C. Amovilli and N. H. March. Relation between the Slater–Kohn– Sham orbitals generated by the one-body potential v(r) of DFT and the Löwdin natural orbitals for inhomogeneous electron liquids. Physics and Chemistry of Liquids, 50(1):121–127, ???? 2012. CODEN PCLQAC. ISSN 0031-9104, 0031-9104, 1026-7727, 1029-0451.

Andre:2004:SCP

[And04] Jean-Marie André. In Silico chemistry: Past, present and future. In Brändas and Kryachko [BK04], pages 3–21. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http://link.springer.

com/book/10.1007/978-94-017-0448-9.

Anonymous:1959:FGL

[Ano59] Anonymous. Florida gets Löwdin. Chemical and Engineering News, 37(31):68–??, August 3, 1959. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605x (electronic). URL http://pubs.

acs.org/doi/pdf/10.1021/cen-v037n031.p068.

Anonymous:1976:BM

[Ano76a] Anonymous. Back matter. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 589– 595. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.springer.

com/content/pdf/bbm:978-1-4757-1659-7/1.pdf.

Anonymous:1976:FM

[Ano76b] Anonymous. Front matter. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages i– ix. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.springer.

com/content/pdf/bfm:978-1-4757-1659-7/1.pdf.

Anonymous:2002:POL

[Ano02] Anonymous. Per-Olov Löwdin curriculum vitae. International Journal of Quantum Chemistry, 90(1):1–11, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Antone:1985:CEL

[Ant85] A. A. Antone. Compact expression for Löwdin’s alpha function. Journal of Mathematical Physics, 26(5):940–942, May 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [Suz87].

Aquino:2020:FON

[ASW20] Fredy W. Aquino, Ravindra Shinde, and Bryan M. Wong. Fractional occupation numbers and self-interaction correction-scaling methods with the Fermi–Löwdin orbital self-interaction correction approach. Journal of Computational Chemistry, 41(12):1200– 1208, May 5, 2020. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bajaj:1975:SBK

[Baj75] M. M. Bajaj. Study of Blumenfeld-Kalmanson-Löwdin effect in Watson–Crick type of double helical DNA structures using biomagnetic technique. Bulletin of the American Physical Society, 20(12):1501, ???? 1975. CODEN BAPSA6. ISSN 0003-0503.

Burrows:2004:SWE

[BC04] B. L. Burrows and M. Cohen. Schrödinger’s wave equation — a Lie algebra treatment. In Brändas and Kryachko [BK04], pages 43–65. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http:

//link.springer.com/book/10.1007/978-94-017-0448-9.

Bubin:2004:QMC

[BCA04] Sergiy Bubin, Mauricio Cafiero, and Ludwik Adamowicz. Quantum mechanical calculations on molecules containing positrons. In Brändas and Kryachko [BK04], pages 521–545. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http://link.springer.

com/book/10.1007/978-94-017-0448-9.

Bartlett:1994:OYO

[BDL^{+}94] Rodney J. Bartlett, Erik Deumens, Per-Olov Löwdin, David A. Micha, Hendrik J. Monkhorst, John R. Sabin, Samuel B. Trickey, and Michael C. Zerner. On the occasion of Yngve Öhrn’s 60th birthday. International Journal of Quantum Chemistry, 28:3–5, 1994. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bruhn:2006:LPA

[BDMC06] George Bruhn, Ernest R. Davidson, Istvan Mayer, and Aurora E. Clark. Löwdin population analysis with and without rotational invariance. International Journal of Quantum Chemistry, 106(9): 2065–2072, ???? 2006. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berggren:1976:ISP

[Ber76] K.-F. Berggren. Inelastic scattering of photons from ionic crystals and effects of overlap. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 141– 169. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.springer.

com/chapter/10.1007/978-1-4757-1659-7_10/.

Bersuker:2004:RCJ

[Ber04] I. B. Bersuker. Reformulation of the concept of Jahn–Teller vibronic coupling effects in theoretical chemistry. In Brändas and Kryachko [BK04], pages 257–271. ISBN 90-481-6687-X (print), 94-

017-0448-1 (e-book). URL http://link.springer.com/book/
10.1007/978-94-017-0448-9.

Berrondo:1976:RPS

[BGC76] M. Berrondo and G. García-Calderón. On resonant potential scattering. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 295–304. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_18/.

Bolcer:2007:DCC

[BH07] John D. Bolcer and Robert B. Hermann. The development of computational chemistry in the United States. In Reviews in Computational Chemistry, volume 5, pages 1–63. John Wiley and Sons, Inc., New York, NY, USA, January 2007. ISSN 1069-3599 (print), 1934-5372 (electronic). LCCN QD39.3.E46 R48. URL http://

onlinelibrary.wiley.com/bookseries/10.1002/SERIES6143.

Block:1976:IOA

[BJ76] R. Block and L. Jansen. Importance of overlap in the analysis of weak exchange interactions by perturbation methods. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 123–140. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_9/.

Bailey:1976:TCQ

[BK76] T. L. Bailey and J. L. Kinsey. Test of the conventional quantum chemistry methods on the hydrogen atom. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 249–263. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_15/.

Brandas:2004:FWQ

[BK04] Erkki J. Brändas and Eugene S. Kryachko, editors. Fundamental World of Quantum Chemistry: A Tribute to the Memory of Per-Olov Löwdin. Volume III. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). xlvi +

677 pp. URL http://link.springer.com/book/10.1007/978-
94-017-0448-9.

Behera:1991:CLC

[BKNS91] L. Behera, T. Kar, P. K. Nandi, and A. B. Sannigrahi. A closer look into the causes of discrepancy between Löwdin and Mulliken atomic charges. Theochem — Journal of Molecular Structure, 82 (1–2):15–21, November 1991. ISSN 0166-1280 (print), 1872-7999 (electronic).

Berrondo:1969:POS

[BL69] Manuel Berrondo and Per-Olov Löwdin. The projection operator for a space spanned by a linearly dependent set. International Journal of Quantum Chemistry, 3(6):767–780, November 1969. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bhatia:1986:WLO

[BM86] Rajendra Bhatia and Kalyan K. Mukherjea. On weighted Löwdin orthogonalization. International Journal of Quantum Chemistry, 29(6):1775–1778, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bosanac:1983:AQC

[Bos83] S. D. Bosanac. Advances in quantum-chemistry, vol 14 — P.-O. Löwdin. Journal of the Chemical Society — Faraday Transactions II, 79(??):761, ???? 1983. CODEN JCFTBS. ISSN 0300-9238 (print), 2050-5663 (electronic).

Bouferguene:2005:ATS

[Bou05] Ahmed Bouferguene. Addition theorem of Slater type orbitals: a numerical evaluation of Barnett–Coulson/Löwdin functions. Journal of Physics A (Mathematical and General), 28(13):2899–2916, 2005. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Boyd:2004:BRB

[Boy04] Russell J. Boyd. Book review: Fundamental World of Quantum Chemistry: A Tribute to the Memory of Per-Olov Löwdin, Volumes 1 and 2 Edited by Erkki J. Brändas (Uppsala University) and Eugene S. Kryachko (Bogoliubov Institute for Theoretical Physics, Kiev, and University of Leuven, Belgium). Kluwer Academic Publishers: Dordrecht. 2003. xxxii + 678 pp (Volume 1) and xxxii + 696 pp (Volume 2). $450.00 (set). ISBN: 1-4020-1286-1 (set). Journal of the American Chemical Society, 126(21):6836, April 28, 2004. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984.

Boyd:2005:BRB

[Boy05] Russell J. Boyd. Book review: Fundamental World of Quantum Chemistry: A Tribute to the Memory of Per-Olov Löwdin, Volume III Edited by Erkki J. Brändas (Uppsala University, Sweden) and Eugene S. Kryachko (Bogoliubov Institute for Theoretical Physics, Kiev, Ukraine, and University of Leuven, Belgium). Kluwer Academic Publishers: Dordrecht, The Netherlands. 2004. xlvi + 677 pp. $225.00. ISBN 1-4020-2583-1. Journal of the American Chemical Society, 127(41):14526, September 9, 2005. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984.

Brandas:1976:LBF

[Brä76] Erkki Brändas. On the Löwdin bracketing function. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 381–392. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_25/.

Brandas:2014:ESG

[Brä14] Erkki Brändas. Editorial: The statement of goals of the International Society for Theoretical Chemical Physics. International Journal of Quantum Chemistry, 114(15):961–962, August 5, 2014. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandas:2017:POL

[Brä17] Erkki J. Brändas. Per-Olov Löwdin — father of quantum chemistry. Molecular Physics, 115(??):1–30, March 2017. CO-DEN MOPHAM. ISSN 0026-8976 (print), 1362-3028 (elec-

tronic). URL https://www.researchgate.net/publication/
315799822_Per-Olov_Lowdin_-_Father_of_Quantum_Chemistry.

Bass:1971:DAR

[BS71a] G. E. Bass and L. J. Schaad. Deoxyribonucleic acid replication. A theoretical study of Löwdin’s mechanism. Journal of the American Chemical Society, 93(18):4585–4591, September 8, 1971. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984.

Bass:1971:DRT

[BS71b] G. E. Bass and L. J. Schaad. DNA replication: a theoretical study of Löwdin’s mechanism. Journal of the American Chemical Society, 93(18):4585–4591, ???? 1971. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984.

Bartlett:1976:NIO

[BS76] Rodney J. Bartlett and David M. Silver. Numerical infinite-order perturbation theory. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 393–408. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_26/.

Bouferguene:2006:GBQ

[BS06] Ahmed Bouferguene and Hassan Safouhi. Gauss–Bessel quadrature: a tool for the evaluation of Barnett–Coulson/Löwdin functions. International Journal of Quantum Chemistry, 106(11): 2398–2407, ???? 2006. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bochevarov:2007:SSR

[BS07] Arteum D. Bochevarov and C. David Sherrill. Some simple results following from Löwdin’s partitioning technique. Journal of Mathematical Chemistry, 42(1):59–64, July 2007. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (electronic). URL http://

link.springer.com/article/10.1007/s10910-006-9097-6.

Buckingham:1965:BRB

[Buc65] David Buckingham. Book review: Molecular orbitals in chemistry, physics, and biology, edited by Per-Olov Löwdin and Bernard Pullman. American Scientist, 53(1):116A–116A, March 1965. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL

http://www.jstor.org/stable/27835959.

Bunge:2017:SSL

[Bun17] Carlos F. Bunge. Specifics on the scientific legacy of Per-Olov Löwdin. Advances in Quantum Chemistry, 74:53–82, 2017. CODEN AQCHA9. ISSN 0065-3276. URL http://www.

sciencedirect.com/science/article/pii/S0065327616300338.

Calais:1976:SCQ

[Cal76] Jean-Louis Calais. Some comments on the quantum-mechanical treatment of defects in ionic crystals. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 171–178. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http:

//link.springer.com/chapter/10.1007/978-1-4757-1659-7_
11/.

Carbo:1988:QCB

[Car88] Ramón Carbó, editor. Quantum chemistry: basic aspects, actual trends: proceedings of an International Workshop on Quantum Chemistry, Girona, Spain, 13–18 June 1988, volume 62 of Studies in physical and theoretical chemistry. Elsevier, Amsterdam, The Netherlands, 1988. ISBN 0-444-87494-1. LCCN QD450.S7 no.62; QD462.A1 I58 1988. NL FL375.

Canuto:2004:HBE

[CC04] Sylvio Canuto and Kaline Coutinho. Hydrogen bonding and the energetics of homolytic dissociation in solution. In Brändas and Kryachko [BK04], pages 581–599. ISBN 90-481-6687-X (print), 94-

017-0448-1 (e-book). URL http://link.springer.com/book/
10.1007/978-94-017-0448-9.

Carbo-Dorca:2004:DVW

[CD04] Ramon Carbó-Dorca. Discussion on the variable Wolfsberg– Helmholtz parameter, a new simplified Löwdin transformation and the characteristic structure of the transformed EHT Hamiltonian matrices. International Journal of Quantum Chemistry, 98 (1):26–32, ???? 2004. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Calais:1977:FHP

[CG77] J. L. Calais and O. Goscinski. Festschrift honoring P.-O. Löwdin — preface. International Journal of Quantum Chemistry, 11(6): 895, ???? 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Calais:1976:PPO

[CGLÖ76a] Jean-Louis Calais, Osvaldo Goscinski, Jan Linderberg, and Yngve Öhrn. Publications of Per-Olov Löwdin 1939–1976. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 13–23. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_2/.

Calais:1976:QSM

[CGLÖ76b] Jean-Louis Calais, Osvaldo Goscinski, Jan Linderberg, and Yngve Öhrn. Quantum science: methods and structure: a tribute to Per-Olov Löwdin. Plenum Press, New York, NY, USA; London, UK, 1976. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). ix + 595 pp. LCCN QC174.125 .Q371. URL http://

www.springerlink.com/content/978-1-4757-1659-7.

Chandrakumar:1990:ALP

[Cha90] N. Chandrakumar. Application of Löwdin projectors to evaluate density-matrix evolutions. Journal of Magnetic Resonance, 88(1): 86–92, June 1990. CODEN JOMRA4. ISSN 0022-2364 (print), 1557-8968 (electronic).

Chojnacki:1976:EEF

[Cho76] H. Chojnacki. External electrical field and proton transfer. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 561–565. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371.

URL http://link.springer.com/content/pdf/10.1007/978-
1-4757-1659-7_41.pdf.

Calais:1962:SMT

[CL62] Jean-Louis Calais and Per-Olov Löwdin. A simple method of treating atomic integrals containing functions of r_{12}. Journal of Molecular Spectroscopy, 8(3):203–211, March 1962. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Calais:1978:CTQ

[CL78] J. L. Calais and P.-O. Lowdin. Current trends in quantum chemistry. Kemisk Tidskrift, 90(11):18–21, ???? 1978. ISSN 0039-6605.

Coleman:1963:DSF

[CLS63] A. J. Coleman, P.-O. Löwdin, and Darwin W. Smith. Discussion on “Structure of Fermion Density Matrices”. Reviews of Modern Physics, 35(3):687–689, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Calais:1989:SIQ

[COM89] J. L. Calais, W. J. Orvillethomas, and Z. B. Maksic. Special issue — quantum-chemistry — electronic-structure of matter — a tribute to Per-Olov Löwdin. Theochem — Journal of Molecular Structure, 58(??):ix, September 1989. ISSN 0166-1280 (print), 1872-7999 (electronic).

Coote:2004:UAS

[CPR04] Michelle L. Coote, Addy Pross, and Leo Radom. Understanding alkyl substituent effects in R–O bond dissociation reactions in open- and closed-shell systems. In Brändas and Kryachko [BK04], pages 563–579. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-

book). URL http://link.springer.com/book/10.1007/978-
94-017-0448-9.

Caldwell:1979:DMM

[CRE79] Dennis Caldwell, Patrick Redington, and Henry Eyring. Density matrix method for orbital localization. Proceedings of the National Academy of Sciences of the United States of America, 76(7):3042– 3045, ???? 1979. CODEN PNASA6. ISSN 0027-8424 (print),

1091-6490 (electronic). URL http://www.pnas.org/content/
76/7/3042.full.pdf+html.

Crossley:1977:LPO

[Cro77] Richard J. S. Crossley. On Löwdin’s projection operators for angular momentum. I. International Journal of Quantum Chemistry, 11(6):917–929, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Clark:2004:DWF

[CSHM04] A. E. Clark, J. L. Sonnenberg, P. J. Hay, and R. L. Martin. Density and wave function analysis of actinide complexes: What can fuzzy atom, atoms-in-molecules, Mulliken, Löwdin, and natural population analysis tell us? Journal of Chemical Physics, 121 (6):2563–2570, August 2004. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Cizek:1987:RIP

[ČVV87] Jiří Čížek, Francine Vinette, and Edward R. Vrscay. Renormalized inner projection, symbolic computation, and Löwdin rational approximants in explicit form. International Journal of Quantum Chemistry, 32(S21):757–758, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

deAndrade:2012:PCP

[dA12] Paulo Cesar Peres de Andrade. Probability current in protein electron transfer: Löwdin population analysis. International Journal of Quantum Chemistry, 112(20):3325–3332, October 15, 2012. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Daudel:1976:CBM

[Dau76] R. Daudel. The chemical bond as a many-electron problem. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 445–458. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_31/.

Davidson:1976:RDM

[Dav76] Ernest R. Davidson. Reduced Density Matrices in Quantum Chemistry, volume 6 of Theoretical chemistry: a series of monographs. Academic Press Inc., New York, USA, 1976. ISBN 0-12-205850-X. viii + 135 pp. LCCN QD462 .D4. URL http://www.

sciencedirect.com/science/book/9780122058509.

Dixon:1975:UPF

[DDH75] M. Dixon, G. Doggett, and G. Howat. Use of partial (and full) Löwdin integral approximations in molecular-orbital calculations on HCN and FCN. Journal of the Chemical Society — Faraday Transactions II, 71(??):452–462, ???? 1975. CODEN JCFTBS. ISSN 0300-9238 (print), 2050-5663 (electronic).

DeAViccaro:1976:SPL

[De 76] M. H. De A. Viccaro. Symmetry property of Löwdin’s “alpha” function. International Journal of Quantum Chemistry, 10(6): 1081–1082, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DelRe:1976:NOP

[Del76] Giuseppe Del Re. The non-orthogonality problem and orthogonalization procedures. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 53–75. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_5/.

Dykstra:2005:TAC

[DFKS05] Clifford E. Dykstra, Gernot Frenking, Kwang S. Kim, and Gustavo E. Scuseria, editors. Theory and Applications of Computational Chemistry: the First Forty Years. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 0-08-044493-8 (print), 0-444-51719-7, 0-444-51904-1 (flexicover). xxvii + 1308 pp. LCCN QD39.3.E46

T45 2005. URL http://www.loc.gov/catdir/enhancements/
fy0625/2005049899-d.html;   http://www.loc.gov/catdir/
enhancements/fy0625/2005049899-t.html.

Doggett:1977:UPL

[Dog77] G. Doggett. Use of partial Löwdin (PLA) and limited expansion of diatomic overlap (LEDO) integral approximation methods in molecular-orbital calculations of isomer pairs HCN, CNH and FCN, CNF. Journal of the Chemical Society — Faraday Transactions II, 73(??):1153–1168, ???? 1977. CODEN JCFTBS. ISSN 0300-9238 (print), 2050-5663 (electronic).

Duff:1971:CFL

[Duf71] K. J. Duff. A computational form for Löwdin’s alpha function. International Journal of Quantum Chemistry, 5(1):111–113, January 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Erdahl:1987:DMD

[ES87] Robert Erdahl and Vedene H. Smith, Jr., editors. Density Matrices and Density Functionals. Proceedings of the A. John (Albert John) Coleman Symposium. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1987. ISBN 90-277-2477-6. LCCN QC174.17.D44 D47 1987.

Froman:1976:AVR

[FG76] Anders Fröman and Osvaldo Goscinski. Acta Vålådalensia revisited: Per-Olov Löwdin in scientific discussion. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 33–51. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_4/.

Froman:1958:CBC

[FH58] Anders Fröman and George Hall, editors. Correspondence between Concepts in Chemistry and Quantum Chemistry: Proceedings of a Panel Meeting Held at Vålådalen, Sweden, August 26–30, 1958, …, Part 1, Acta Vålådalensia. Uppsala University, Uppsala, Sweden, 1958. Proceedings of a panel meeting held at Vålådalen, Sweden, 1958.

Fischer-Hjalmars:1955:RSQ

[FHL55] Inga Fischer-Hjalmars and Per-Olov Löwdin. Report from the symposium on quantum theory of molecules: Stockholm and Uppsala, March 21–25, 1955. Svensk kemisk tidskrift, 67(8):365–398, 1955. CODEN SKTIAF. ISSN 0371-0742.

Fieschi:1957:ASW

[FL57] R. Fieschi and P.-O. Löwdin. Atomic state wavefunctions generated by projection operators. Technical Report TU-4, Kvantkemi, Upsala Universitet, Uppsala, Sweden, 1957. ?? pp.

Freeman:1958:QMK

[FL58] A. J. Freeman and Per-Olov Löwdin. Quantum mechanical kinetic energy transformation. Physical Review, 111(5):1212–1213, September 1, 1958. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v111/i5/p1212_1.

Froman:1962:VTC

[FL62] A. Fröman and Per-Olov Löwdin. Virial theorem and cohesive energies of solids, particularly ionic crystals. Journal of Physics and Chemistry of Solids, 23:75–84, 1962. CODEN JPCSAW. ISSN 0022-3697 (print), 1879-2553 (electronic).

Froelich:1983:HFS

[FL83] Piotr Froelich and Per-Olov Löwdin. On the Hartree–Fock scheme for a pair of adjoint operators. Journal of Mathematical Physics, 24(1):88–96, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Froman:2007:IQC

[FL07] A. Fröman and J. Linderberg. Inception of quantum chemistry at Uppsala. Acta Universitatis Upsaliensis C, Organisation och historia, 78:1–46, 2007. ISBN 91-554-6754-7. ISSN 0502-7454.

Flores-Moreno:2014:SNU

[FMPM^{+}14] Roberto Flores-Moreno, Edwin Posada, Félix Moncada, Jonathan Romero, Jorge Charry, Manuel Díaz-Tinoco, Sergio A. González, Néstor F. Aguirre, and Andrés Reyes. Software news & updates: LOWDIN: the any particle molecular orbital code. International Journal of Quantum Chemistry, 114(1):50–56, January 5, 2014. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Froman:1976:PCA

[Frö76] Anders Fröman. Properties of compressed atoms from a spherical cellular model. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 179–199. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_12/.

Fu:2011:ECM

[FWMZ11] Liang-Yu Fu, Guang-Zhong Wang, Bin-Guang Ma, and Hong-Yu Zhang. Exploring the common molecular basis for the universal DNA mutation bias: Revival of Löwdin mutation model. Biochemical and Biophysical Research Communications, 409(3):367– 371, June 10, 2011. CODEN BBRCA9. ISSN 0006-291x (print), 1090-2104 (electronic).

Goscinski:1984:LL

[GC84] Osvaldo Goscinski and Jean-Louis Calais. Löwdin lectures. International Journal of Quantum Chemistry, 26(2):295, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goidenko:2004:QEH

[GL04] Igor Goidenko and Leonti Labzowsky. QED effects in heavy elements. In Brändas and Kryachko [BK04], pages 407–421. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http://link.

springer.com/book/10.1007/978-94-017-0448-9.

Grassi:2008:LCE

[GLA^{+}08] A. Grassi, G. M. Lombardo, G. G. N. Angilella, G. Forte, N. H. March, C. Van Alsenoy, and R. Pucci. Löwdin correlation energy density of the inhomogeneous electron liquid in some closedshell molecules at equilibrium geometry. Physics and Chemistry of Liquids, 46(5):484–490, ???? 2008. CODEN PCLQAC. ISSN 0031-9104, 0031-9104, 1026-7727, 1029-0451.

Gruninger:1970:CAA

[GÖL70] John Gruninger, Yngve Öhrn, and Per-Olov Löwdin. Comments on the analysis of atomic correlation energies. Journal of Chemical Physics, 52(11):5551–5554, June 1, 1970. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Gordon:1984:AQC

[Gor84] M. S. Gordon. Advances in quantum-chemistry, vol 15 — P.-O. Löwdin. Journal of the American Chemical Society, 106(10):3067, ???? 1984. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984.

Goscinski:1976:BCI

[Gos76] Osvaldo Goscinski. Bonding character of inner-shell orbitals in diatomic molecules. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 427–432. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_29/.

Gray:1965:BRB

[Gra65] Harry B. Gray. Book review: Molecular orbitals in chemistry, physics, and biology, edited by Per-Olov Löwdin and Bernard Pullman. Journal of Chemical Education, 42(3):A224–??, March 1965. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic). See [LP64].

Grau:1977:ABL

[Gra77] D. Grau. An application of the Brillouin–Löwdin perturbation expansion: Lower bounds to the energy eigenvalues of the rigid rotator in an electric field. International Journal of Quantum Chemistry, 11(6):931–941, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gavroglou:2012:NPN

[GS12] Kostas Gavroglou and Ana Simões. Neither Physics nor Chemistry: a History of Quantum Chemistry. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01618-4 (hardcover). xiv + 351 pp. LCCN QD462 .G38 2012.

Hall:1976:NFC

[Hal76] G. G. Hall. A new formulation of the correlation problem. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 433–444. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_30/.

Hall:2004:POL

[Hal04] G. G. Hall. Per-Olov Löwdin. In Brändas and Kryachko [BK04], pages 1–2. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book).

URL http://link.springer.com/book/10.1007/978-94-017-
0448-9.

Harcourt:1976:CEP

[Har76] R. D. Harcourt. The calculation of the exchange parameter J = (1/2)(E_{singlet} − E_{triplet}) for two equivalent electrons using canonical molecular orbitals. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 105– 122. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.springer.

com/chapter/10.1007/978-1-4757-1659-7_8/.

Harris:2004:IEC

[Har04] Frank E. Harris. Integrals for exponentially correlated four-body systems of general angular symmetry. In Brändas and Kryachko [BK04], pages 115–127. ISBN 90-481-6687-X (print), 94-017-0448-

1 (e-book). URL http://link.springer.com/book/10.1007/
978-94-017-0448-9.

Herman:1972:CSS

[HDK72] Frank Herman, Norris W. Dalton, and Thomas R. Koehler, editors. Computational Solid State Physics: Proceedings of an International Symposium Held October 6–8, 1971, in Wildbad, Germany, The IBM Research Symposia Series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. ISBN 0-306-30583-6, 1-4684-1979-X (print), 1-4684-1977-3 (e-book). LCCN QC176-176.9. URL http://link.springer.

com/book/10.1007/978-1-4684-1977-1.

Hehenberger:1976:NAW

[Heh76] Michael Hehenberger. Numerical aspects of Weyl’s theory. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 265–275. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_16/.

Hehenberger:2018:POL

[Heh18] Michael Hehenberger. Per-Olov Löwdin’s impact on a “Lost Son”. Advances in Quantum Chemistry, 77:1–18, 2018. CODEN

AQCHA9. ISSN 0065-3276. URL https://www.sciencedirect.
com/science/article/pii/S0065327617300680.

Hettema:2000:QCH

[Het00] Hinne Hettema, editor. Quantum Chemistry: Classic Scientific papers, volume 8 of World scientific series in 20th century chemistry. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2000. ISBN 981-02-2771-X, 981-279-576-6 (e-book). xxxix + 478 pp. LCCN QD462 .Q34613 2000.

Hirschfelder:1959:LRI

[HL59] J. O. Hirschfelder and P.-O. Löwdin. Long-range interaction of two 1s-hydrogen atoms expressed in terms of natural spin-orbitals. Molecular Physics, 2(3):229–258, July 1959. CODEN MOPHAM. ISSN 0026-8976 (print), 1362-3028 (electronic).

Hirschfelder:1965:LRI

[HL65] J. O. Hirschfelder and Per-Olov Löwdin. Long-range interaction of two 1s-hydrogen atoms expressed in terms of natural spin-orbitals. Molecular Physics, 9(5):491–496, 1965. CODEN MOPHAM. ISSN 0026-8976 (print), 1362-3028 (electronic).

Howat:1976:PHF

[HL76] G. Howat and S. Lunell. Projected Hartree–Fock calculations on the ground and first excited ^{1}Σ_{g}^{+} states of the hydrogen molecule. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 491–498. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_35/.

Hermann:1983:LPL

[HL83] M. R. Hermann and P. W. Langhoff. Löwdin partitioning in a Lanczos basis: Applications to scattering states. International Journal of Quantum Chemistry, 23(1):135–146, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hamza:2001:ORLa

[HM01a] A. Hamza and I. Mayer. Overlap repulsion with Löwdin’s pairing theorem. International Journal of Quantum Chemistry, 82(2):53–59, ???? 2001. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract/77502751/
START;   http://www3.interscience.wiley.com/cgi-bin/fulltext/
77502751/FILE?TPL=ftx_start;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=77502751&PLACEBO=IE.pdf.

Hamza:2001:ORLb

[HM01b] A. Hamza and I. Mayer. Overlap repulsion with Löwdin’s pairing theorem II. The leading term. International Journal of Quantum Chemistry, 82(3):105–112, ???? 2001. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract/77004286/
START;   http://www3.interscience.wiley.com/cgi-bin/fulltext/
77004286/FILE?TPL=ftx_start;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=77004286&PLACEBO=IE.pdf.

Harris:1963:DNE

[HNLD63] Frank E. Harris, Robert K. Nesbet, Per-Olov Löwdin, and Ernest R. Davidson. Discussion on “Natural Expansions and Properties of the Chemical Bond”. Reviews of Modern Physics, 35(3):629–630, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.35.629;   http://
rmp.aps.org/abstract/RMP/v35/i3/p629_1.

Hoffman:1967:BRB

[Hof67] Joseph G. Hoffman. Book review: Quantum Theory of Atoms, Molecules and the Solid State: A Tribute to John C. Slater, by Per-Olov Löwdin. Physics Today, 20(11):91–92, November 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/?PTO/20/91/1;
http://www.physicstoday.org/resource/1/phtoad/v20/i11/
p91_s1.

Jansen:1977:PLS

[Jan77] L. Jansen. P.-O. Löwdin’s scientific and other activities. International Journal of Quantum Chemistry, 11(6):897–905, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1987:MLF

[JBW87] H. W. Jones, B. Bussery, and C. A. Weatherford. An “E matrix” for the Löwdin α function, expanded in a Taylor series: an analytic treatment of molecular charge density near the origin. International Journal of Quantum Chemistry, 32(S21):693–698, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Jiao:2015:ALC

[JH15] Li Guang Jiao and Yew Kam Ho. Application of Löwdin’s canonical orthogonalization method to the Slater-type orbital configuration-interaction basis set. International Journal of Quantum Chemistry, 115(7):434–441, April 5, 2015. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jhon:1987:SRC

[JL87] Mu Shik Jhon and Per-Olov Löwdin. Some remarks on certain magnetic properties of water in the study of cancer. International Journal of Quantum Chemistry, 32(S14):9–14, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Jones:1984:FTC

[Jon84] Herbert W. Jones. Formulas for three-center exchange integrals by computer implementation of the Löwdin α-function method. International Journal of Quantum Chemistry, 26(S18):61–66, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Jones:1988:AEM

[Jon88] H. W. Jones. Analytical evaluation of multicenter molecular integrals over Slater-type orbitals using expanded Löwdin alphafunctions. Physical Review A (Atomic, Molecular, and Optical Physics), 38(2):1065–1068, July 15, 1988. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Jones:1991:ALA

[Jon91] Herbert W. Jones. Analytic Löwdin alpha-function method for two-center electron-repulsion integrals over Slater-type orbitals. Journal of Computational Chemistry, 12(10):1217–1222, December 1991. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jones:1992:LFO

[Jon92] Herbert W. Jones. Löwdin α-function, overlap integral, and computer algebra. International Journal of Quantum Chemistry, 41 (5):749–754, March 5, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jansen:1976:CBN

[JP76] H. B. Jansen and P. Pyykkö. Calculation of the bromine nuclear pseudoquadrupole coupling constant in the LiBr molecule using a density-of-states function deduced from overlap integrals. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 409–416. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_27/.

Jansson:2004:TCK

[JSBL04] C. M. Jansson, N. Salhi-Benachenhou, and S. Lunell. Theoretical calculations of kinetic isotope effects for a series of substituted aziridines. In Brändas and Kryachko [BK04], pages 601–617. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http:

//link.springer.com/book/10.1007/978-94-017-0448-9.

Jones:1978:MFS

[JW78] Herbert W. Jones and C. A. Weatherford. A modified form of Sharma’s formula for STO Löwdin alpha functions with recurrence relations for the coefficient matrix. International Journal of Quantum Chemistry, 14(S12):483–488, ???? 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1989:LAF

[JW89] H. W. Jones and C. A. Weatherford. The Löwdin alpha-function and its application to the multi-center molecular integral problem over Slater-type orbitals. Theochem — Journal of Molecular Structure, 58(??):233–243, September 1989. ISSN 0166-1280 (print), 1872-7999 (electronic).

Karlsson:2002:SEI

[Kar02] Göran Karlsson. The scientific, educational and internationalization influence of Per-Olov Löwdin. Advances in Quantum Chemistry, 41:1–8, 2002. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S006532760241043X.

Katriel:2004:WSP

[Kat04] Jacob Katriel. Weights of spin and permutational symmetry adapted states for arbitrary elementary spins. In Brändas and Kryachko [BK04], pages 23–41. ISBN 90-481-6687-X (print), 94-

017-0448-1 (e-book). URL http://link.springer.com/book/
10.1007/978-94-017-0448-9.

Kibler:2004:SQM

[KD04a] M. R. Kibler and M. Daoud. On supersymmetric quantum mechanics. In Brändas and Kryachko [BK04], pages 67–96. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http://link.

springer.com/book/10.1007/978-94-017-0448-9.

Krems:2004:CAM

[KD04b] Roman Krems and Alexander Dalgarno. Collisions of atoms and molecules in external magnetic fields. In Brändas and Kryachko [BK04], pages 273–294. ISBN 90-481-6687-X (print), 94-017-0448-

1 (e-book). URL http://link.springer.com/book/10.1007/
978-94-017-0448-9.

Killingbeck:1969:ALP

[Kil69] J. Killingbeck. An application of Löwdin projection operators to spin-orbit coupling theory. Physica Status Solidi. B, Basic Research, 36(1):K67–K69, ???? 1969. CO-DEN PSSBBD. ISSN 0370-1972 (print), 1521-3951 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
pssb.19690360159/abstract.

Kong:1987:SPT

[KJL87] Young Shik Kong, Mu Shik Jhon, and Per-Olov Löwdin. Studies on proton transfers in water clusters and DNA base pairs. International Journal of Quantum Chemistry, 32(S14):189–209, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kanilov:1967:ESP

[KKS67] V. I. Kanilov, V. A. Kuprievich, and O. V. Shramko. Elektronnaia struktura par osnovanii dnk v vozbuzhdennykh sostoianiakh i levdinovskii mekhanizm mutasii. (Russian) [The electronic structure of pair bases of DNA in excited states, and Lowdin’s mutation mechanism]. Doklady Akademii Nauk SSSR, 177(6):1465–1468, December 21, 1967. CODEN DANKAS. ISSN 0002-3264.

Klessinger:1966:BRBb

[Kle66a] M. Klessinger. Book review: Molecular Orbitals in Chemistry, Physics, and Biology. A Tribute to R. S. Mulliken. Edited By P.- O. Löwdin and B. Pullman. Angewandte Chemie, International Edition in English, 5(1):146, January 1966. CODEN ACIEAY.

ISSN 0570-0833. URL http://onlinelibrary.wiley.com/doi/
10.1002/anie.196601461/.

Klessinger:1966:BRBa

[Kle66b] M. Klessinger. Book review: Molecular Orbitals in Chemistry, Physics, and Biology. A Tribute to R. S. Mulliken. Herausgeg. von P.-O. Löwdin und B. Pullman. Academic Press, New York-London 1964. 1. Aufl., VIII, 578 S., zahlr. Abb. u. Tab., geb. $22. Angewandte Chemie, 78(2):154, ???? 21, 1966. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/ange.19660780219.

Kolos:1976:LRI

[Koł76] W. Kołos. Long-range interaction in some two-electron systems. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 465–477. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_33/.

Konig:1968:BRB

[Kön68] Edgar König. Book review: Advances in Quantum Chemistry, by P. O. Löwdin. Band 4. Academic Press, New York und London 1968. 334 Sciten. Preis: $17.00. Berichte der Bunsen-Gesellschaft für Physikalische Chemie, 1963–1990, 72 (7):888, September 1968. CODEN BBPCAX. ISSN 0005-

9021. URL http://onlinelibrary.wiley.com/doi/10.1002/
bbpc.19680720733/abstract.

Kornblum:1979:BRB

[Kor79] Zvi C. Kornblum. Book review: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Methods, by Per-Olov Löwdin. Journal of the Electrochemical Society, 126(??):490C–??, ???? 1979. CODEN JESOAN. ISSN 0013-4651 (print), 1945-7111 (electronic).

Kryachko:2002:OSP

[Kry02] Eugene S. Kryachko. The origin of spontaneous point mutations in DNA via Löwdin mechanism of proton tunneling in DNA base pairs: Cure with covalent base pairing. International Journal of Quantum Chemistry, 90(2):910–923, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kashiwagi:1973:GLO

[KS73] H. Kashiwagi and F. Sasaki. A generalization of the Löwdin orthogonalization. International Journal of Quantum Chemistry, 7(S7):515–520, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaijser:1976:ISM

[KS76] Per Kaijser and Vedene H. Smith, Jr. On inversion symmetry in momentum space. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 417–425. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_28/.

Kar:1988:EBS

[KS88] T. Kar and A. B. Sannigrahi. Effect of basis set on Mulliken and Löwdin atomic charges, bond orders and valencies of some polar molecules. Theochem — Journal of Molecular Structure, 42(1–2): 47–54, April 1988. ISSN 0166-1280 (print), 1872-7999 (electronic).

Kar:1987:CAC

[KSM87] T. Kar, A. B. Sannigrahi, and D. C. Mukherjee. Comparison of atomic charges, valencies and bond orders in some hydrogenbonded complexes calculated from Mulliken and Löwdin SCF density-matrices. Theochem — Journal of Molecular Structure, 38(1–2):93–101, September 1987. ISSN 0166-1280 (print), 1872-7999 (electronic).

Kutzelnigg:1967:BRBb

[Kut67a] W. Kutzelnigg. Book review: Advances in Quantum Chemistry, Vol. 2. Edited by P. O. Löwdin. Angewandte Chemie, International Edition in English, 6(3):278, March 1967. CODEN

ACIEAY. ISSN 0570-0833. URL http://onlinelibrary.wiley.
com/doi/10.1002/anie.196702781/abstract.

Kutzelnigg:1967:BRBa

[Kut67b] W. Kutzelnigg. Book review: Advances in Quantum Chemistry, Vol. 2. Herausgeg. von P. O. Löwdin, Academic Press, Inc., New York-London 1965. 1. Aufl., XI, 372 S., 79 Abb., geb $14.50. Angewandte Chemie, 79(6):282, March 21, 1967. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
ange.19670790624/abstract.

Lowdin:1956:SSC

[LA56] Per-Olov Löwdin and K. Appel. Studies of self-consistent fields. III. analytic wave functions for the argon-like ions and for the first row of the transition metals. Physical Review, 103(6):1746–1755, September 15, 1956. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v103/i6/p1746_1.

Ladik:1976:TTD

[Lad76] János J. Ladik. Towards the theoretical determination of the conformation of biological macromolecules. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 539–548. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_39/.

Larsson:2004:QCL

[Lar04] Sven Larsson. Quantum chemistry, localization, superconductivity, and Mott–Hubbard U. In Brändas and Kryachko [BK04], pages 247–256. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-

book). URL http://link.springer.com/book/10.1007/978-
94-017-0448-9.

Laskowski:1976:GCM

[Las76] Bernard Laskowski. The generator coordinate method illustrated on the hydrogen molecule. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 479– 489. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.springer.

com/chapter/10.1007/978-1-4757-1659-7_34/.

Lepadatu:1999:OMD

[LAS99] C. Lepadatu, H. Adamsky, and H. H. Schmidtke. s-orbital mixing into d-orbital manifolds in complex compounds analyzed using Löwdin’s orthogonalization procedure. Journal of Molecular Structure. Theochem, 488(??):241–245, October 1999. CODEN THEODJ. ISSN 0166-1280 (print), 1872-7999 (electronic).

Lathouwers:1976:BBH

[Lat76] Luc Lathouwers. Biorthonormal bases in Hilbert space. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 77–91. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_6/.

Lazzeretti:2004:UAI

[Laz04] Paolo Lazzeretti. Unified approach to intensities in vibrational spectroscopies via dynamic electromagnetic shieldings at the nuclei of a molecule. In Brändas and Kryachko [BK04], pages 199–213. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book).

URL http://link.springer.com/book/10.1007/978-94-017-
0448-9.

Larsson:1976:OMC

[LB76] S. Larsson and R. E. Brown. Orbital methods and correlation errors in expectation values. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 459– 463. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.springer.

com/chapter/10.1007/978-1-4757-1659-7_32/.

Lowdin:1997:E

[LBG^{+}97] P.-O. Löwdin, E. Brändas, O. Goscinski, S. Lunell, J. R. Sabin, M. C. Zerner, and N. Y. Öhrn. Editorial. International Journal of Quantum Chemistry, 61(1):v, 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Langhoff:2004:ALM

[LBH04] P. W. Langhoff, J. A. Boatz, and R. J. Hinde. Applications of Löwdin’s metric matrix: Atomic spectral methods for electronic structure calculations. In Brändas and Kryachko [BK04], pages 97–114. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-

book). URL http://link.springer.com/book/10.1007/978-
94-017-0448-9.

Lowdin:1997:I

[LBÖ^{+}97] Per-Olov Löwdin, Erkki Brändas, Yngve Öhrn, Osvaldo Goscinski, John R. Sabin, Sten Lunell, and Michael C. Zerner. Introduction. International Journal of Quantum Chemistry, 63(1):

1–2, 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X

(electronic). URL http://www3.interscience.wiley.com/
cgi-bin/abstract?ID=42552;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=42552&PLACEBO=IE.pdf. Spe-

cial issue in memory of Jean-Louis Calais (Part I of III). [Parts II and III appear as numbers 2 and 3 of this volume, 63.].

Lowdin:1978:KFT

[LC78] Per-Olov Löwdin and Jean-Louis Calais. Kvantkemi — från teori till praktik. (Swedish) [Quantum chemistry – from theory to practice]. Svensk kemisk tidskrift, ??(??):??, 1978. CODEN KETIAL. ISSN 0039-6605.

Lowdin:1981:SDC

[LCC81] Per-Olov Löwdin, Jean Louis Calais, and Jacques M. Calazans. Simple derivation of conditions for instability in the Hartree– Fock and projected Hartree–Fock schemes. International Journal of Quantum Chemistry, 20(6):1201–1215, ???? 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:PTW

[LCG82a] P.-O. Löwdin, J. L. Calais, and O. Goscinski. Proceedings of the Tarfala Workshop — Held in the Tarfala Valley Lapland, Sweden, August 1978. International Journal of Quantum Chemistry, 21 (1):215, ???? 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:DFT

[LCG82b] Per-Olov Löwdin, Jean-Louis Calais, and Osvaldo Goscinski. Density functional theory. International Journal of Quantum Chemistry, 21(2):357, ???? 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:Pc

[LCG82c] Per-Olov Löwdin, Jean-Louis Calais, and Osvaldo Goscinski. Preface. International Journal of Quantum Chemistry, 21(1):215, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:P

[LCG83] Per-Olov Löwdin, Jean-Louis Calais, and Osvaldo Goscinski. Preface. International Journal of Quantum Chemistry, 23(1):v, 1983.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1972:TSB

[LCH72] Per-Olov Löwdin, J.-L. Calais, and M. R. Hayns. A theoretical study of the behaviour of solids under high pressure and the Borelius’ law, with applications to the alkali halides. Journal of Nonmetals, 1(1):63–78, 1972. CODEN JNMTA3. ISSN 0090-3477.

Leopold:1976:SLC

[LCK76] J. G. Leopold, M. Cohen, and J. Katriel. A study of Löwdin’s criterion for completeness of basis sets. International Journal of Quantum Chemistry, 10(6):1049–1056, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laskowski:1973:EGT

[LCLL73] Bernard Laskowski, Jean-Louis Calais, Per-Olov Löwdin, and Piet V. Leuven. Electron gas test for the alternant molecular orbital method. Journal of Physics C: Solid State Physics, 6(18): 2777–2787, September 18, 1973. CODEN JPSOAW. ISSN 0022-3719.

Lowdin:1984:EN

[LCÖ84] Per-Olov Löwdin, Jean-Louis Calais, and Yngve Öhrn. Editor’ Note. International Journal of Quantum Chemistry, 26(4):i, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1970:RCP

[LCS^{+}70] P.-O. Löwdin, J.-L. Calais, G. Sperber, M. Hayns, and T. Ahlenius. Research on cohesive properties of solids in the Uppsala quantum chemistry group. 4th Nordic solid state conference (abstracts), page 17, 1970.

Leschber:1989:OPE

[LD89] Y. Leschber and J. P. Draayer. Orthogonal polynomial expansion for the Löwdin transformation. Journal of Physics A (Mathematical and General), 22(14):2675–2685, ???? 1989. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Levy:1976:UVF

[Lev76] Mel Levy. Universal variational functionals of electron densities, first-order density matrices, and natural spin-orbitals and solution of the v-representability problem. Proceedings of the National Academy of Sciences of the United States of America, 76 (12):6062–6065, ???? 1976. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.pnas.

org/content/76/12/6062.full.pdf+html.

Lowdin:1955:RSQ

[LFH55] Per-Olov Löwdin and I. Fischer-Hjalmars. Report from the Symposium on Quantum Theory of Molecules, Stockholm and Uppsala, 1955. Svensk kemisk tidskrift, 67(??):365–398, 1955. CODEN KETIAL. ISSN 0039-6605. See especially pp. 369, 370, 373, 375, 379, 380 and 383.

Lowdin:1989:CSA

[LFM89a] Per-Olov Löwdin, Piotr Froelich, and Manoj Míshra. On the change of spectra associated with unbounded similarity transformations of a many-particle Hamiltonian and the occurrence of resonance states in the method of complex scaling. Part II. Applications to the Hartree–Fock scheme based on the bi-variational principle. Advances in Quantum Chemistry, 20:185–237, 1989. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608606279.

Lowdin:1989:SPB

[LFM89b] Per-Olov Löwdin, Piotr Froelich, and Manoj Mishra. Some properties of the bivariational Hartree–Fock scheme for complex symmetric many-particle operators. International Journal of Quantum Chemistry, 36(2):93–103, August 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1968:EPS

[LG68] Per-Olov Löwdin and Osvaldo Goscinski. The exchange phenomenon, the symmetric group, and the spin degeneracy problem. Preprint 166, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, 1968. iv + 99 pp.

Lowdin:1969:EPS

[LG69] Per-Olov Löwdin and Osvaldo Goscinski. The exchange phenomenon, the symmetric group, and the spin degeneracy problem. International Journal of Quantum Chemistry, 4(S3b):533–591, January 13–18, 1969. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1971:SPT

[LG71] P.-O. Löwdin and O. Goscinski. Studies in perturbation theory. XIV. Treatment of constants of motion, degeneracies and symmetry properties by means of multidimensional partitioning. International Journal of Quantum Chemistry, 5(S5):685–705, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Lowdin:1989:IDI

[LGBC89] P.-O. Löwdin, O. Goscinski, E. Brändas, and J. L. Calais. Issue dedicated to Ivar Waller on his 90th birthday, June 11, 1988 — preface. International Journal of Quantum Chemistry, 35(6):703, Spring 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1978:CPS

[LGC78a] Per-Olov Löwdin, Osvaldo Goscinski, and Jean-Louis Calais, editors. Chemical physics of surfaces, catalysis and membranes: proceedings of a colloquium held 28–31 August 1977 at the University of Uppsala, volume 2S of International journal of quantum chemistry. John Wiley and Sons, Inc., New York, NY, USA, 1978. CODEN IJQCB2. ISBN 0-471-05247-7. ISSN 0020-7608 (print), 1097-461X (electronic). LCCN ????

Lowdin:1978:QCS

[LGC78b] Per-Olov Löwdin, Osvaldo Goscinski, and Jean-Louis Calais, editors. Quantum chemistry — a scientific melting pot: a symposium sponsored by the Quantum Chemistry Group to mark the 500th Anniversary of the University of Uppsala. Held 31 August through 4 September, 1977, volume 12 (supplement 1) of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1978. CODEN IJQBDZ. ISSN 0360-8832. LCCN QD462.A1 Q367 1978.

Lowdin:1989:PIW

[LGJLCB89] Per-Olov Löwdin, Osvaldo Goscinski, Jean-Louis-Calais, and Erkki Brändas. Professor Ivar Waller 90-years anniversary June 11, 1988. International Journal of Quantum Chemistry, 35(6): 703, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Linderberg:1993:LPO

[Lin93] J. Linderberg. Per-Olov Löwdin honorary issue — preface. Theoretica Chimica Acta, 86(1–2):1–2, July 1993. CODEN TCHAAM. ISSN 0040-5744.

Linderberg:2003:BMP

[Lin03] Jan Linderberg. Biographical memoir: Per-Olov Löwdin: 28 October 1916–6 October 2000. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 147(2):175–180, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL http://www.amphilsoc.org/sites/default/files/210.
pdf;   http://www.jstor.org/stable/1558261.

Lindgren:2004:CBM

[Lin04] Ingvar Lindgren. Comparison between the many-body perturbative and Green’s-function approaches for calculating electron binding energies and affinities: Brueckner and Dyson orbitals. In Brändas and Kryachko [BK04], pages 215–245. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http://link.springer.

com/book/10.1007/978-94-017-0448-9.

Linderberg:2010:SIP

[Lin10] Jan Linderberg. Scientific input by Per-Olov Löwdin to the Quantum Theory Project. Molecular Physics, 108(21–23):2847–2851, ???? 2010. CODEN MOPHAM. ISSN 0026-8976 (print), 1362-3028 (electronic).

Lundqvist:1951:CCI

[LL51] Stig O. Lundqvist and Per-Olov Löwdin. On the calculation of certain integrals occurring in the theory of molecules, especially three-centre and four-centre integrals. Arkiv för fysik / utgivet av K. Svenska Vetenskapsakademien, 3(10):147–154, 1951. CODEN AFYSA7. ISSN 0365-2440.

Lindner:1968:ULB

[LL68] Peter Lindner and Per-Olov Löwdin. Upper and lower bounds in second-order perturbation theory and the Unsöld approximation. International Journal of Quantum Chemistry. Symposium, 2(S2):

161–173, 1968. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1970:CLB

[LL70] Per-Olov Löwdin and Tiong-Koon Lim. Calculation of lower bounds to energy eigenvalues by reduced density matrices and the representability problem. International Journal of Quantum Chemistry, 3(S3b):697–702, January 13–18, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Laskowski:1972:TCM

[LL72] Bernard Laskowski and Per-Olov Löwdin. Treatment of constants of motion in the variation principle. I. Symmetry properties of variational wavefunctions. Chemical Physics Letters, 16(1):1–4, September 15, 1972. CODEN CHPLBC. ISSN 0009-2614 (print), 1873-4448 (electronic).

Lowdin:1972:SCTb

[LM72a] P.-O. Löwdin and P. K. Mukherjee. Some comments on the timedependent variation principle. Chemical Physics Letters, 14(1): 1–7, 1972. CODEN CHPLBC. ISSN 0009-2614 (print), 1873-4448 (electronic).

Lowdin:1972:SCTa

[LM72b] Per-Olov Löwdin and P. K. Mukherjee. Some comments on the time-dependent variation principle. Report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, January 1972.

Lowdin:1992:SSG

[LM92] Per-Olov Löwdin and István Mayer. Some studies of the general Hartree–Fock method. Advances in Quantum Chemistry, 24:79–114, 1992. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S006532760860101X.

Lowdin:1989:SRP

[LN89] Per-Olov Löwdin and Bengt Nagel. Some remarks on the perturbation and variation perturbation calculations of the free energy for quantum systems. International Journal of Quantum Chemistry, 35(6):839–850, Spring 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Ivar Waller issue.

Lowdin:1976:PISa

[LÖ76] Per-Olov Löwdin and Yngve Öhrn, editors. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics: held at Sanibel Island, Florida, January 19–24, 1976, volume 10 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1976. CODEN IJQBDZ. ISBN ???? ISSN 0360-8832. LCCN QD462 .I61 1976.

Lowdin:1977:PISa

[LÖ77] Per-Olov Löwdin and Yngve Öhrn, editors. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Methods, held at Sanibel Island, Florida, January 16–22, 1977, volume 11 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1977. CODEN IJQBDZ. ISSN 0360-8832. LCCN QD462 .I61 1977.

Lowdin:1978:PISb

[LÖ78] Per-Olov Löwdin and Yngve Öhrn, editors. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Methods, held at Flagler Beach, Florida, March 12–18, 1978, volume 12 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1978. CODEN IJQBDZ. ISBN 0-471-05633-2 (paperback). ISSN 0360-8832. LCCN QD462 .I61 1978.

Lowdin:1979:PISa

[LÖ79] Per-Olov Löwdin and Yngve Öhrn, editors. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods, held in honor of Alberte and Bernard Pullman at Flagler Beach, Florida, March 11–17, 1979, volume 13 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1979. CODEN IJQBDZ. ISSN 0360-8832. LCCN QD462 .I61 1979. US$30.00 (paperback).

Lowdin:1980:PISb

[LÖ80] Per-Olov Löwdin and Yngve Öhrn, editors. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods, held at Flagler Beach, Florida, March 10–20, 1980, volume 14 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1980. CODEN IJQBDZ. ISBN 0-471-09189-8 (paperback). ISSN 0360-8832. LCCN QD462 .I61 1980. Held as two symposia: the first, The International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena and Computational Quantum Chemistry, Mar. 9–15, 1980, in honor of E. Bright Wilson; the second, The International Symposium on Fundamentals of Quantum Statistics and Quantum Theory, particularly Density Matrix Formalism, Mar. 16–20, 1980. 20th Sanibel symposia.

Lowdin:1981:PISa

[LÖ81a] Per-Olov Löwdin and Yngve Öhrn, editors. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry, held at Flagler Beach, Florida, March 8–14, 1981, volume 15 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1981. CODEN IJQBDZ. ISBN 0-471-86671-7. ISSN 0360-8832. LCCN QD462 .I61 1981; QH 345 I609 1981.

Lowdin:1981:PISd

[LÖ81b] Per-Olov Löwdin and Yngve Öhrn. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry Held in Flagler Beach, Florida, 8–14 March 1981 — Introduction. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 15:iii–iv, 1981. CODEN IJQSDI. ISSN 0161-3642.

Lowdin:1982:PIW

[LÖ82a] P.-O. Löwdin and Y. Öhrn. Proceedings of the International Workshop on Perturbation Theory at Large Order — Held at the Sheraton Palm Coast Resort Inn Flagler Beach, Florida 2–4 March 1981 — Introduction. International Journal of Quantum Chemistry, 21(1):1–2, ???? 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:I

[LÖ82b] Per-Olov Löwdin and Yngve Öhrn. Introduction. International Journal of Quantum Chemistry, 21(1):1–2, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:PISa

[LÖ82c] Per-Olov Löwdin and Yngve Öhrn, editors. Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter, held at Flagler Beach, Florida, March 1–3 and 7–13, 1982, volume 16 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1982. CODEN IJQBDZ. ISBN 0-471-89124-X (paperback). ISSN 0360-8832. LCCN QD462 .I61 1982.

Lowdin:1985:PISb

[LÖ85] Per-Olov Löwdin and Yngve Öhrn, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at Marineland, Florida, March 14–16, 1985, volume 12 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1985. CODEN IJQBDZ. ISBN 0-471-84878-6 (paperback). ISSN 0360-8832. LCCN QH345.I53 1985.

Lowdin:1986:PISb

[LÖ86] Per-Olov Löwdin and Yngve Öhrn, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at Marineland, Florida, March 6–8, 1986, volume 13 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1986. CODEN IJQBDZ. ISBN 0-471-84878-6 (paperback). ISSN 0360-8832. LCCN QH345.I53 1986.

Linderberg:2017:POL

[LÖBS17] Jan Linderberg, Yngve Öhrn, Erkki J. Brändas, and John R. Sabin. Per-Olov Löwdin. Advances in Quantum Chemistry, 74:1–7, 2017. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327616300302.

Lowdin:1983:PISa

[LÖM83] Per-Olov Löwdin, Yngve Öhrn, and David Micha, editors. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry, held at Flagler Beach, Florida, March 3–12, 1983, volume 17 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1983. CODEN IJQBDZ. ISBN 0-471-88169-4. ISSN 0360-8832. LCCN QD462 .I61 1983.

Lowdin:2000:PIS

[LÖS00] Per-Olov Löwdin, Yngve Öhrn, and John R. Sabin, editors. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory, held at the Ponce de Leon Resort, St. Augustine, Florida, February 25–March 3, 2000, volume 80(4/5) of International Journal of Quantum Chemistry. John Wiley and Sons, Inc., New York, NY, USA, 2000. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Quantum Chemistry Symposium No. 34.

Linderberg:2002:AQC

[LÖS02] Jan Linderberg, Yngve Öhrn, and John R. Sabin. Advancing quantum chemistry: Per-Olov Löwdin 1916–2000. Advances in Quantum Chemistry, 41:x–xv, 2002. CODEN AQCHA9. ISSN

0065-3276. URL http://www.sciencedirect.com/science/
article/pii/S0065327602410428.

Lowdin:1985:Ia

[LÖSZ85a] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 28(S12): v, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1985:Ib

[LÖSZ85b] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 28(S19):1– 2, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1985:PISc

[LÖSZ85c] P.-O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory, Scattering Problems, Many-Body Phenomena, and Computational Quantum Chemistry, Held at Marineland, Florida, March 18–23, 1985 — Introduction. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 19:1–2, 1985. CODEN IJQSDI. ISSN 0161-3642.

Lowdin:1985:PISd

[LÖSZ85d] P.-O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology, Marineland, Florida, March 14–16, 1985 — Introduction. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 12:v, 1985. CODEN IJQBDZ. ISSN 0360-8832.

Lowdin:1985:PISa

[LÖSZ85e] Per-Olov Löwdin, Yngve Öhrn, John R. Sabin, and Michael C. Zerner, editors. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry: held at Marineland, Florida, March 18–23, 1985, volume 19 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1985. CODEN IJQBDZ. ISBN 0-471-84879-4. ISSN 0360-8832. LCCN QD462 .I61 1985.

Lowdin:1986:Ia

[LÖSZ86a] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 30(S13):v, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1986:Ib

[LÖSZ86b] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 30(S20):1– 2, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Lowdin:1986:PISc

[LÖSZ86c] P.-O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory, Scattering Problems, Many-Body Phenomena, and Computational Quantum Chemistry, Held at Marineland, Florida, March 10–15, 1986 — Introduction. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 20:1–2, 1986. CODEN IJQSDI. ISSN 0161-3642.

Lowdin:1986:PISd

[LÖSZ86d] P.-O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology — Held at Marineland, Florida, March 6–8, 1986 — Introduction. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 13:v, 1986. CODEN IJQBDZ. ISSN 0360-8832.

Lowdin:1986:PISa

[LÖSZ86e] Per-Olov Löwdin, Yngve Öhrn, John R. Sabin, and Michael C. Zerner, editors. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry: held at Marineland, Florida, March 10–15, 1986, volume 20 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1986. CODEN IJQBDZ. ISBN 0-471-85972-9. ISSN 0360-8832. LCCN QD462 .I61 1986.

Lowdin:1987:I

[LÖSZ87a] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 32(S21): 1, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Lowdin:1987:PISb

[LÖSZ87b] P.-O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology Held at Whitney Marine-Biological Laboratory, Marineland, Florida, March 12–14, 1987 — Introduction. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 14:v, 1987. CO-DEN IJQBDZ. ISSN 0360-8832.

Lowdin:1987:PISa

[LÖSZ87c] P.-O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods Held at Marineland, Florida, March 12–21, 1987 — Introduction. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 21:3, 1987. CODEN IJQSDI. ISSN 0161-3642.

Lowdin:1987:PIS

[LÖSZ87d] Per-Olov Löwdin, Yngve Öhrn, John R. Sabin, and Michael C. Zerner, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at Marineland, Florida, March 12–14, 1987, volume 14 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1987. CODEN IJQBDZ. ISSN 0360-8832. LCCN QH345.I53 1987.

Lowdin:1988:Ia

[LÖSZ88a] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 34(S15): v, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1988:Ib

[LÖSZ88b] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 34(S22): ix, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Lowdin:1988:PISd

[LÖSZ88c] P.-O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology Held at Marineland, Florida, March 12–19, 1988 — Introduction. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 15:v, 1988. CODEN IJQBDZ. ISSN 0360-8832.

Lowdin:1988:PISc

[LÖSZ88d] P.-O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods — Held at Marineland, Florida, March 12–19, 1988 — Introduction. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 22:ix, 1988. CODEN IJQSDI. ISSN 0161-3642.

Lowdin:1993:PISa

[LÖSZ93] Per-Olov Löwdin, Yngve Öhrn, John R. Sabin, and Michael C. Zerner, editors. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods: held at St. Augustine, Florida, March 13–20, 1993, volume 27 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1993. CODEN IJQBDZ. ISBN 0-471-02505-4. ISSN 0360-8832. LCCN QD453.A1 I58 S95, no.27.

Lowdin:1998:PISa

[LÖSZ98] Per-Olov Löwdin, Yngve Öhrn, John R. Sabin, and Michael C. Zerner, editors. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory: held at Ponce de Leon Resort, St. Augustine, Florida, February 21–27, 1998, volume 70(4/5) of International journal of quantum chemistry. John Wiley and Sons, Inc., New York, NY, USA, 1998. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). Quantum Chemistry Symposium No. 32.

Lowdin:1999:PISa

[LÖSZ99] Per-Olov Löwdin, Yngve Öhrn, John R. Sabin, and Michael C. Zerner, editors. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory: held at Ponce de Leon Resort, St. Augustine, Florida, February 21–27, 1998, volume 75(4/5) of International journal of quantum chemistry. John Wiley and Sons, Inc., New York, NY, USA, 1999. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). Quantum Chemistry Symposium No. 33.

Linderberg:1976:EWM

[LÖT76] Jan Linderberg, Yngve Öhrn, and Poul W. Thulstrup. Energy Weighted Maximum Overlap (EWMO). In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 93–103. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.

springer.com/chapter/10.1007/978-1-4757-1659-7_7/.

Lowdin:1939:LTO

[Löw39] Per-Olov Löwdin. Lorentz-transformationen och den kinematiska relativitetsprincipen. (Swedish) [The Lorentz transformation and the kinematic relativity principle]. Elementa: tidskrift for matematik, fysik och kemi, 22(??):161–169, 1939. CODEN EMNTAE. ISSN 0013-5933.

Lowdin:1947:QMC

[Löw47a] P.-O. Löwdin. A quantum mechanical calculation of the cohesive energy, the interionic distance, and the elastic constants of some ionic crystals. Arkiv för matematik, astronomi och fysik, utgivet av K. Svenska vetenskapsakademien, 35A(??):1–10, 1947. CODEN AMAFAX. ISSN 0365-4133. This paper introduces Löwdin’s “canonical orthonormalization” scheme. See [Löw56b] for his “symmetric orthonormalization”.

Lowdin:1947:CRC

[Löw47b] Per-Olov Löwdin. Contributions to the report from the conference of the Swedish National Committee for Physics. Technical report, ????, ????, 1947.

Lowdin:1948:TISa

[Löw48a] P.-O. Löwdin. A Theoretical Investigation Into Some Properties of Ionic Crystals. Dissertation, Institute of Mechanics and Theoretical Physics, Upsala Universitet, Uppsala, Sweden, May 1, 1948. 126 pp.

Lowdin:1948:CRC

[Löw48b] Per-Olov Löwdin. Contributions to the report from the conference of the Swedish National Committee for Physics. Technical report, ????, ????, 1948.

Lowdin:1948:QMCa

[Löw48c] Per-Olov Löwdin. A quantum mechanical calculation of the cohesive energy, the interionic distance, and the elastic constants of some ionic crystals. I. Arkiv för matematik, astronomi och fysik, utgivet av K. Svenska vetenskapsakademien, 35A(9):1–10, 1948. CODEN AMAFAX. ISSN 0365-4133.

Lowdin:1948:QMCb

[Löw48d] Per-Olov Löwdin. A quantum mechanical calculation of the cohesive energy, the interionic distance, and the elastic constants of some ionic crystals. II. The elastic constants c_{12} and c_{44}. Arkiv för matematik, astronomi och fysik, utgivet av K. Svenska vetenskapsakademien, 35A(30):1–18, 1948. CODEN AMAFAX. ISSN 0365-4133.

Lowdin:1948:TISb

[Löw48e] Per-Olov Löwdin. A theoretical investigation into some properties of ionic crystals: a quantum mechanical treatment of the cohesive energy, the interionic distance, the elastic constants, and the compression at high pressures, with numerical applications to some alkali halides. Almqvist and Wiksell, Uppsala, Sweden, 1948. xi + 126 + 1 pp. LCCN QB 498 349 (??).

Lowdin:1949:CRC

[Löw49] Per-Olov Löwdin. Contributions to the report from the conference of the Swedish National Committee for Physics. Technical report, ????, ????, 1949.

Lowdin:1950:NMS

[Löw50a] Per-Olov Löwdin. A note on the method of steepest descents with a remark on T. Ljunggren’s paper “Contributions to the theory of diffraction of electromagnetic waves by spherical particles.”. Arkiv för fysik / utgivet av K. Svenska Vetenskapsakademien, 2 (36):367–370, 1950. CODEN AFYSA7. ISSN 0365-2440.

Lowdin:1950:NOP

[Löw50b] Per-Olov Löwdin. On the non-orthogonality problem connected with the use of atomic wave functions in the theory of molecules and crystals. Journal of Chemical Physics, 18(3):365–375, March 1950. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (elec-

tronic). URL http://jcp.aip.org/resource/1/jcpsa6/v18/
i3/p365_s1.

Lowdin:1951:MNI

[Löw51a] P.-O. Löwdin. On the methods of numerical integration used in determining self-consistent fields. Shelter island conference, nasonr report, Office of Naval Research, Washington, DC, USA (??), 1951. 187–194 pp.

Lowdin:1951:LEC

[Löw51b] Per-Olov Löwdin. Letter to the editor: Calculation of electric dipole moments of some heterocyclics. Journal of Chemical Physics, 19(10):1323–1324, October 1951. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://

jcp.aip.org/resource/1/jcpsa6/v19/i10/p1323_s1.

Lowdin:1951:NQM

[Löw51c] Per-Olov Löwdin. A note on the quantum-mechanical perturbation theory. Journal of Chemical Physics, 19(11):1396–1401, November 1951. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://jcp.aip.org/resource/1/
jcpsa6/v19/i11/p1396_s1.

Lowdin:1951:QMCa

[Löw51d] Per-Olov Löwdin. On the quantum-mechanical calculation of the cohesive energy of molecules and crystals. Part I. A general energy formula for the ground state. Journal of Chemical Physics, 19 (12):1570–1578, December 1951. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://jcp.aip.org/

resource/1/jcpsa6/v19/i12/p1570_s1.

Lowdin:1951:QMCb

[Löw51e] Per-Olov Löwdin. On the quantum-mechanical calculation of the cohesive energy of molecules and crystals. Part II. Treatment of the alkali metals with numerical applications to sodium. Journal of Chemical Physics, 19(12):1579–1591, December 1951. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL

http://jcp.aip.org/resource/1/jcpsa6/v19/i12/p1579_s1.

Lowdin:1952:NIO

[Löw52] Per-Olov Löwdin. On the numerical integration of ordinary differential equations of the first order. Quarterly of Applied Mathematics, 10(2):97–111, July 1952. CODEN QAMAAY. ISSN 0033-569x (print), 1552-4485 (electronic).

Lowdin:1953:LEA

[Löw53a] Per-Olov Löwdin. Letter to the editor: Approximate formulas for many-center integrals in the theory of molecules and crystals. Journal of Chemical Physics, 21(2):374–375, February 1953.

CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (elec-

tronic). URL http://jcp.aip.org/resource/1/jcpsa6/v21/
i2/p374_s1.

Lowdin:1953:MOT

[Löw53b] Per-Olov Löwdin. On the molecular-orbital theory of conjugated organic compounds with application to the perturbed benzene ring. Journal of Chemical Physics, 21(3):496–515, March 1953. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (elec-

tronic). URL http://jcp.aip.org/resource/1/jcpsa6/v21/
i3/p496_s1.

Lowdin:1953:SAS

[Löw53c] Per-Olov Löwdin. Studies of atomic self-consistent fields. I. Calculation of Slater functions. Physical Review, 90(1):120–125, April 1, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v90/i1/p120_1.

Lowdin:1954:CMI

[Löw54a] Per-Olov Löwdin. Calculations of molecular integrals in Uppsala. In Symposium on Molecular Physics at Nikko, Japan in 1953, pages 113–117. Maruzen, Tokyo, Japan, 1954. LCCN ????

Lowdin:1954:MAM

[Löw54b] Per-Olov Löwdin. A method of alternant molecular orbitals. In Symposium on Molecular Physics at Nikko, Japan in 1953, pages 13–16. ????, ????, 1954. LCCN ????

Lowdin:1954:RSM

[Löw54c] Per-Olov Löwdin. Recent simplifications in the molecular orbital theory of calculating energy levels. In Proceedings of the International Conference of Theoretical Physics in Japan, pages 599–611. ????, ????, September 1954.

Lowdin:1954:SAS

[Löw54d] Per-Olov Löwdin. Studies of atomic self-consistent fields. II. Interpolation problems. Physical Review, 94(6):1600–1609, June 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v94/i6/p1600_1.

Lowdin:1954:U

[Löw54e] Per-Olov Löwdin. Studies of atomic self-consistent fields. II. Interpolation problems. Quarterly Report 4, Solid-State and Molecular Theory Group, MIT, Cambridge, MA, USA, ???? 1954. 1–?? pp. Published in [Löw53c, Löw54d].

Lowdin:1955:EHF

[Löw55a] Per-Olov Löwdin. An extension of the Hartree–Fock method to include correlation effects. In Les Electrons dans les Métaux. Dixième Conference Solvay, Bruxelles 1954, pages 71–88. ????, ????, 1955. LCCN ????

Lowdin:1955:QTMa

[Löw55b] Per-Olov Löwdin. Quantum theory of many-particle systems. I. Physical interpretations by means of density matrices, natural spin-orbitals, and convergence problems in the method of configurational interaction. Physical Review, 97(6):1474–1489, March 15, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v97/i6/p1490_1.

Lowdin:1955:QTMb

[Löw55c] Per-Olov Löwdin. Quantum theory of many-particle systems. II. Study of the ordinary Hartree–Fock approximation. Physical Review, 97(6):1490–1508, March 15, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v97/i6/p1490_1.

Lowdin:1955:QTMc

[Löw55d] Per-Olov Löwdin. Quantum theory of many-particle systems. III. Extension of the Hartree–Fock scheme to include degenerate systems and correlation effects. Physical Review, 97(6):1509–1520, March 15, 1955. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v97/i6/p1509_1.

Lowdin:1956:ECT

[Löw56a] Per-Olov Löwdin. Electronic correlation in the theory of molecular energy levels. Texas Journal of Science, 8:163–??, 1956. CO-DEN TJSCAU. ISSN 0040-4403. Proceedings of the Molecular Quantum Mechanics Conference, December 7–9, 1955, Texas, p.

Lowdin:1956:QTC

[Löw56b] Per-Olov Löwdin. Quantum theory of cohesive properties of solids. Advances in Physics, 5(17):1–172, January 1956. CODEN AD-PHAH. ISSN 0001-8732 (print), 1460-6976 (electronic). This paper introduces Löwdin’s “symmetric orthonormalization” scheme. See [Löw47a] for his “canonical orthonormalization”.

Lowdin:1957:PSQ

[Löw57] Per-Olov Löwdin. Present situation of quantum chemistry. Journal of Physical Chemistry, 61(1):55–68, January 30, 1957. CO-DEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic).

Lowdin:1958:NFM

[Löw58a] P.-O. Löwdin. Nature des fonctions de la Mésomerie. (French) [Nature of mesomeric functions]. In Proc. Paris Symposium, volume LXXXII, pages 23–37. Centre National de la Recherche Scientifique, Paris, France, 1958.

Lowdin:1958:CFO

[Löw58b] Per-Olov Löwdin. Calcul des fonctions d’onde moléculaire. Technical report, Centre National de la Recherche Scientifique, Paris, France, 1958. 23 pp.

Lowdin:1958:KKB

[Löw58c] Per-Olov Löwdin. Den kovalenta kemiska bindningen i kvantmekanisk belysning. (Swedish) [The covalent chemical bond in a quantum-mechanical light]. Elementa: tidskrift for matematik, fysik och kemi, 40(??):1–16, 1958. CODEN EMNTAE. ISSN 0013-5933.

Lowdin:1958:EIV

[Löw58d] Per-Olov Löwdin. An elementary iteration-variation procedure for solving the Schrödinger equation. Preliminary Technical report 11, Kvantkemi, Upsala Universitet, Uppsala, Sweden, April 23, 1958. Published in [Löw63h].

Lowdin:1958:GHF

[Löw58e] Per-Olov Löwdin. Generalizations of the Hartree–Fock scheme. Annales Academiae regiae scientiarum upsaliensis, 2(??):127–135, 1958. CODEN ???? ISSN ????

Lowdin:1958:RSS

[Löw58f] Per-Olov Löwdin. Report from the Summer School in Quantum Chemistry held in Vålådalen, Sweden, July 26–August 30, 1958. Report ??, Kvantkemi, Upsala Universitet, Uppsala, Sweden, 1958.

Lowdin:1958:SAR

[Löw58g] Per-Olov Löwdin. Some aspects of the recent development of the theory of the electronic structure of atoms. In W. O. Milligan, editor, Proceedings of the Robert A. Welch Foundation Conferences on Chemical Research. II. Atomic Structure. December 1–3, 1958, Houston, Texas, pages 4–75. Robert A. Welch Foundation, Houston, TX, USA, 1958. CODEN PRAWAC. ISSN 0557-1588.

Lowdin:1958:SDP

[Löw58h] Per-Olov Löwdin. Spin degeneracy problem. Coll. Int. Centre Nat. Rech. Sci. Paris, 82(??):23–??, 1958.

Lowdin:1959:SPV

[Löw59a] P.-O. Löwdin. Scaling problem, virial theorem, and connected relations in quantum mechanics. Journal of Molecular Spectroscopy, 3(1):46–66, February 1959. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Lowdin:1959:CPM

[Löw59b] Per-Olov Löwdin. Correlation problem in many-electron quantum mechanics. I. Review of different approaches and discussion of some current ideas. Advances in Chemical Physics, 2:207–322, 1959. CODEN ADCPAA. ISSN 0065-2385 (print), 1934-4791 (electronic).

Lowdin:1959:IVA

[Löw59c] Per-Olov Löwdin. En iterations-variationsmetod for att Lösa egenvardesproblem. (Swedish) [An iteration-variation method for solving the eigenvalue problem]. In Nord-SAM, Karlskrona and Lund, Sweden, pages 199–209. ????, ????, 1959.

Lowdin:1960:ETT

[Löw60a] Per-Olov Löwdin. Expansion theorems for the total wave function and extended Hartree–Fock schemes. Reviews of Modern Physics, 32(2):328–334, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//rmp.aps.org/abstract/RMP/v32/i2/p328_1.

Lowdin:1960:QTE

[Löw60b] Per-Olov Löwdin. Quantum theory of electronic structure of molecules. Annual Review of Physical Chemistry, 11:107–132, 1960. CODEN ARPLAP. ISSN 0066-426X (print), 1545-1593 (electronic).

Lowdin:1960:SPT

[Löw60c] Per-Olov Löwdin. Studies in perturbation theory. V. Some aspects on the exact self-consistent field theory. Preprint 5, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, August 15, 1960. 33 pp.

Lowdin:1961:NST

[Löw61a] Per-Olov Löwdin. Note on the separability theorem for electron pairs. Journal of Chemical Physics, 35(1):78–81, July 1961. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

URL http://link.aip.org/link/doi/10.1063/1.1731935.

Lowdin:1961:PCC

[Löw61b] Per-Olov Löwdin. The principle of causality, the chemical bond and modern quantum chemistry. Annales Academiae regiae scientiarum upsaliensis, 5(??):63–78, 1961.

Lowdin:1962:QTN

[Löw62a] P.-O. Löwdin. Quantum theory of nonmetallic crystals. Band theory, valence bond, and tight-binding calculations. Journal of Applied Physics, 33(1):251–280, January 1962. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850.

Lowdin:1962:BTV

[Löw62b] Per-Olov Löwdin. Band theory, valence bond, and tight-binding calculations. Journal of Applied Physics, 33(1):251–280, January 1962. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (elec-

tronic), 1520-8850. URL http://link.aip.org/link/doi/10.
1063/1.1777106.

Lowdin:1962:EIV

[Löw62c] Per-Olov Löwdin. An elementary iteration-variation procedure for solving the Schrödinger equation. Unpublished manuscript., 1962.

Lowdin:1962:ECS

[Löw62d] Per-Olov Löwdin. Exchange, correlation, and spin effects in molecular and solid-state theory. Reviews of Modern Physics, 34(1):80– 87, January 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://rmp.aps.org/

abstract/RMP/v34/i1/p80_1.

Lowdin:1962:CSM

[Löw62e] Per-Olov Löwdin. Is cancer a sub-molecular disease caused by “proton exchange”? Unpublished manuscript., June 1, 1962.

Lowdin:1962:NCM

[Löw62f] Per-Olov Löwdin. The normal constants of motion in quantum mechanics treated by projection technique. Reviews of Modern Physics, 34(3):520–530, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Lowdin:1962:QGA

[Löw62g] Per-Olov Löwdin. Quantum genetics and the aperiodic solid. Technical report TN 85, Kvantkemi, Upsala Universitet, Uppsala, Sweden, November 11, 1962.

Lowdin:1962:QTM

[Löw62h] Per-Olov Löwdin. Quantum theory of molecules of importance in cancer research. Unpublished manuscript., June 1, 1962.

Lowdin:1962:SRP

[Löw62i] Per-Olov Löwdin. Some recent progress in the quantum theory of many-electron systems. Technical report TN 83, Kvantkemi, Upsala Universitet, Uppsala, Sweden, August 15, 1962.

Lowdin:1962:SPTa

[Löw62j] Per-Olov Löwdin. Studies in perturbation theory. IV. Solution of eigenvalue problem by projection operator formalism. Journal of Mathematical Physics, 3(5):969–982, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/
p969_s1.

Lowdin:1962:SPTb

[Löw62k] Per-Olov Löwdin. Studies in perturbation theory. V. Some aspects on the exact self-consistent field theory. Journal of Mathematical Physics, 3(6):1171–1184, November/December 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i6/p1171_s1.

Lowdin:1963:SDH

[Löw63a] P.-O. Löwdin. I. the symmetry dilemma in the Hartree–Fock scheme. II. symmetry properties of the NSO’s. III. Variation of the density matrix. Sanibel Notes., 1963.

Lowdin:1963:QTTa

[Löw63b] P.-O. Löwdin. Quantum theory of time-dependent phenomena treated by the evolution operator technique. Technical Report 100, Kvantkemi, Upsala Universitet, Uppsala, Sweden, June 1963. ???? pp. Published in [Löw67i].

Lowdin:1963:ISAa

[Löw63c] Per-Olov Löwdin, editor. International Symposium on Atomic and Molecular Quantum Mechanics: Proceedings in Honor of Professor Egil A. Hylleraas and arranged by the University of Florida at Sanibel Island, 14–19 January 1963, volume 35(3) of Reviews of Modern Physics. American Institute of Physics, Woodbury, NY, USA, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Lowdin:1963:ISAb

[Löw63d] Per-Olov Löwdin. International Symposium on Atomic and Molecular Quantum Mechanics: Proceedings in Honor of Professor Egil A. Hylleraas and arranged by the University of Florida at Sanibel Island, 14–19 January 1963: Introduction. Reviews of Modern Physics, 35(3):415, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Lowdin:1963:PTD

[Löw63e] Per-Olov Löwdin. Proton tunneling in DNA and its biological implications. Reviews of Modern Physics, 35(3):724–732, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Lowdin:1963:QG

[Löw63f] Per-Olov Löwdin. Quantum genetics. International Science and Technology, 17:64–68, 71–72, 74, 102, May 1963. CODEN ISCTAB. ISSN 0097-3777.

Lowdin:1963:RCI

[Löw63g] Per-Olov Löwdin. Remarks in connection with the International Conference on Atomic and Molecular Quantum Mechanics in Honour of Prof. E. A. Hylleraas on Sanibel Island, January 13–19, 1963. Technical report TN 95, Kvantkemi, Upsala Universitet, Uppsala, Sweden, January 13, 1963. 9 pp.

Lowdin:1963:SPTa

[Löw63h] Per-Olov Löwdin. Studies in perturbation theory. I. An elementary iteration-variation procedure for solving the Schrödinger equation by partitioning technique. Journal of Molecular Spectroscopy, 10(1):12–33, January 1963. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Lowdin:1963:WRO

[Löw63i] Per-Olov Löwdin. Wave and reaction operators in the quantum theory of many-particle systems. Reviews of Modern Physics, 35 (3):702–709, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Lowdin:1964:SPTd

[Löw64a] P.-O. Löwdin. Studies in perturbation theory. VII. Localized perturbation. Journal of Molecular Spectroscopy, 14(1):119–130, October 1964. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Lowdin:1964:SPTe

[Löw64b] P.-O. Löwdin. Studies in perturbation theory. VIII. Separation of Dirac equation and study of the spin-orbit coupling and Fermi contact terms. Journal of Molecular Spectroscopy, 14(1):131–144, October 1964. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Lowdin:1964:AQC

[Löw64c]  Per-Olov Löwdin, editor.  Advances in Quantum Chem-

istry, volume 1 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1964. CODEN AQCHA9. ISSN 0065-3276. xi + 385 pp. LCCN QD453 .A28 v.1.

URL http://www.sciencedirect.com/science/article/pii/
S006532760860369X.

Lowdin:1964:AMW

[Löw64d] Per-Olov Löwdin. Angular momentum wavefunctions constructed by projection operators. Reviews of Modern Physics, 36(4):966– 976, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Lowdin:1964:EPT

[Löw64e] Per-Olov Löwdin. Effect of proton tunneling in DNA on genetic information and problems of mutations, aging, and tumors. Biopolymers Symposia, 1(??):161–181, May 1964. ISSN 0523-6827.

Lowdin:1964:LAF

[Löw64f] Per-Olov Löwdin. Linear algebra and the fundaments of quantum theory. Technical report TN 125, Kvantkemi, Upsala Universitet, Uppsala, Sweden, July 1, 1964. 81 pp.

Lowdin:1964:MOE

[Löw64g] Per-Olov Löwdin. Molecular orbitals in the exact SCF theory. In Löwdin and Pullman [LP64], pages 37–55. LCCN QD461 .L872 1964.

Lowdin:1964:P

[Löw64h] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 1:vii–viii, 1964. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S006532760860372X.

Lowdin:1964:SAQ

[Löw64i] Per-Olov Löwdin. Some aspects of quantum biology. Biopolymers Symposia, 1(??):293–311, May 1964. ISSN 0523-6827.

Lowdin:1964:SAD

[Löw64j] Per-Olov Löwdin. Some aspects on DNA replication; incorporation errors and proton transfer. In Bernard Pullman, editor, Electronic Aspects of Biochemistry: proceedings of the International Symposium held at Ravello, Italy, September 16–18, 1963, page ?? Academic Press Inc., New York, USA, 1964. LCCN QH345.A1 I5 1963; QU4 I58 1963.

Lowdin:1964:SPTa

[Löw64k] Per-Olov Löwdin. Studies in perturbation theory. II. Generalization of the Brillouin–Wigner formalism. Journal of Molecular Spectroscopy, 13(3):326–331, July 1964. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Lowdin:1964:SPTb

[Löw64l] Per-Olov Löwdin. Studies in perturbation theory. III. Solution of the Schrödinger equation under a variation of a parameter. Journal of Molecular Spectroscopy, 13(3):331–337, July 1964. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Lowdin:1964:SPTc

[Löw64m] Per-Olov Löwdin. Studies in perturbation theory. VI. Contraction of secular equations. Journal of Molecular Spectroscopy, 14 (1):112–118, October 1964. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Lowdin:1965:CUL

[Löw65a] P.-O. Löwdin. The calculation of upper and lower bounds of energy eigenvalues in perturbation theory by means of partitioning technique. Technical Report 150, Kvantkemi, Upsala Universitet, Uppsala, Sweden, July 1965.

Lowdin:1965:AQC

[Löw65b]  Per-Olov Löwdin, editor.  Advances in Quantum Chem-

istry, volume 2 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1965. CODEN AQCHA9. ISSN 0065-3276. xi + 372 pp. LCCN QD453 .A28 v.2.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600660.

Lowdin:1965:AOD

[Löw65c] Per-Olov Löwdin. Arvsanlagen och deras Förändringar — ur kvantgenetisk synpunkt. (Swedish) [Inheritance and its changes — from a quantum genetic viewpoint]. Svensk naturvetenskap, ?? (??):??, 1965. CODEN SVNAB5. ISSN 0373-1359.

Lowdin:1965:IET

[Löw65d] Per-Olov Löwdin. Isotope effect in tunneling and its influence on mutation rates. Mutation Research, 2(3):218–221, June 1965. CO-DEN MUREAV. ISSN 0027-5107 (print), 1873-135X (electronic).

Lowdin:1965:P

[Löw65e] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 2:vii–viii, 1965. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600696.

Lowdin:1965:PISa

[Löw65f] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic and Molecular Orbital Quantum Theory, Dedicated to Professor Robert Sanderson Mulliken, Sanibel Island, Florida, 18–23 January 1965, volume 43 (10 part 2) of Journal of Chemical Physics. American Institute of Physics, Woodbury, NY, USA, November 15, 1965. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Lowdin:1965:PISb

[Löw65g] Per-Olov Löwdin. Proceedings of the International Symposium on Atomic and Molecular Orbital Quantum Theory, Dedicated to Professor Robert Sanderson Mulliken, Sanibel Island, Florida, 18–23 January 1965: Introduction. Journal of Chemical Physics, 43(10 (part 2)):S1, November 15, 1965. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.

aip.org/link/doi/10.1063/1.1701489.

Lowdin:1965:QGA

[Löw65h] Per-Olov Löwdin. Quantum genetics and the aperiodic solid. Some aspects of the biological problems of heredity, mutations, aging and tumors in view of the quantum theory of the DNA molecule. Advances in Quantum Chemistry, 2:213–360, 1965. CODEN

AQCHA9. ISSN 0065-3276. URL http://www.sciencedirect.
com/science/article/pii/S0065327608600763.

Lowdin:1965:QTT

[Löw65i] Per-Olov Löwdin. Quantum theory of time-dependent phenomena treated by the evolution operator technique. Technical report TN 100, Kvantkemi, Upsala Universitet, Uppsala, Sweden, June 15, 1965. 71 pp. Also issued as QTP TF number 77.

Lowdin:1965:SPTa

[Löw65j] Per-Olov Löwdin. Studies in perturbation theory. IX. Connection between various approaches in the recent development-evaluation of upper bounds to energy eigenvalues in Schrödinger’s perturbation theory. Journal of Mathematical Physics, 6(8):1341–1353, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i8/p1341_s1.

Lowdin:1965:SPTb

[Löw65k] Per-Olov Löwdin. Studies in perturbation theory. X. Lower bounds to energy eigenvalues in perturbation theory. Ground state. Physical Review, 139(2A):A357–A372, July 19, 1965. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Lowdin:1965:SPTc

[Löw65l] Per-Olov Löwdin. Studies in perturbation theory. XI. Lower bounds to energy eigenvalues, ground state, and excited states. Journal of Chemical Physics, 43(10 (part 2)):S175–S185, November 15, 1965. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Lowdin:1965:SPTd

[Löw65m] Per-Olov Löwdin. Studies in perturbation theory. XII. The calculation of upper and lower bounds of energy eigenvalues in perturbation theory by means of partitioning techniques. Report TN 150, Kvantkemi, Upsala Universitet, Uppsala, Sweden, July 1, 1965. 255–292 pp.

Lowdin:1966:PHFb

[Löw66a] P.-O. Löwdin. The projected Hartree–Fock method. Technical report TF 83, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, March 1966. 41 + 14 pp.

Lowdin:1966:SPT

[Löw66b] P.-O. Löwdin. Studies in perturbation theory. XII. The calculation of upper and lower bounds of energy eigenvalues in perturbation theory by means of partitioning techniques. In C. H. Wilcox, editor, Perturbation Theory, pages 255–294. John Wiley and Sons, Inc., New York, NY, USA, 1966.

Lowdin:1966:CUL

[Löw66c] Per-Olov Löwdin. The calculation of upper and lower bounds of energy eigenvalues in perturbation theory by means of partitioning techniques. In Perturbation Theory and its Applications in Quantum Mechanics (Proc. Adv. Sem. Math. Res. Center, U.S. Army, Theoret. Chem. Inst., Univ. of Wisconsin, Madison, Wis., 1965), pages 255–294. John Wiley and Sons, Inc., New York, NY, USA, 1966.

Lowdin:1966:CPJ

[Löw66d] Per-Olov Löwdin. Comments on Professor John C. Slater’s paper, “Cohesion in Monovalent Metals”. In Quantum theory of atoms, molecules, and the solid state: a tribute to John C. Slater [Löw66h], pages 15–?? LCCN QC174.1 .L827 1966 335.

Lowdin:1966:GAC

[Löw66e] Per-Olov Löwdin. Group algebra, convolution algebra, and applications to quantum mechanics. Technical report TN 176, Kvantkemi, Upsala Universitet, Uppsala, Sweden, July 15, 1966.

Lowdin:1966:PHFa

[Löw66f] Per-Olov Löwdin. The projected Hartree–Fock method. Technical report TN 170, Kvantkemi, Upsala Universitet, Uppsala, Sweden, March 1, 1966.

Lowdin:1966:PHFc

[Löw66g] Per-Olov Löwdin. The projected Hartree–Fock method. An extension of the independent-particles schemes. In Quantum theory of atoms, molecules, and the solid state: a tribute to John C. Slater [Löw66h], pages 601–?? LCCN QC174.1 .L827 1966 335.

Lowdin:1966:QTA

[Löw66h] Per-Olov Löwdin, editor. Quantum theory of atoms, molecules, and the solid state: a tribute to John C. Slater. Academic Press Inc., New York, USA, 1966. xvi + 641 pp. LCCN QC174.1 .L827 1966 335.

Lowdin:1966:RP

[Löw66i] Per-Olov Löwdin. Reports and publications. Report ??, Solid State and Molecular Theory Group, MIT, Cambridge, MA, July 1966. 97–101 pp.

Lowdin:1966:SPH

[Löw66j] Per-Olov Löwdin. Some properties of the hydrogen bond in biochemistry with particular reference to the stability of the genetic code. Technical report QB 33, Kvantkemi, Upsala Universitet, Uppsala, Sweden, August 31, 1966.

Lowdin:1967:AQC

[Löw67a]  Per-Olov Löwdin, editor.  Advances in Quantum Chem-

istry, volume 3 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1967. CODEN AQCHA9.

ISSN 0065-3276. xiv + 392 pp. LCCN QD453 .A28 v.3.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600805.

Lowdin:1967:EPLa

[Löw67b] Per-Olov Löwdin. Eigenvalue problem in a linearly dependent basis and the super secular equation. Preprint 108, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, June 1, 1967.

Lowdin:1967:EPLb

[Löw67c] Per-Olov Löwdin. Eigenvalue problem in a linearly dependent basis and the super secular equation. International Journal of Quantum Chemistry. Symposium, 1(S1):811–827, 1967. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1967:GAC

[Löw67d] Per-Olov Löwdin. Group algebra, convolution algebra, and applications to quantum mechanics. Reviews of Modern Physics, 39(2):259–287, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://falcon.

kvac.uu.se/swe/personal/PerOlovLowdinPub.html;   http:
//link.aps.org/doi/10.1103/RevModPhys.39.259;   http:/
/rmp.aps.org/abstract/RMP/v39/i2/p259_1.

Lowdin:1967:NQC

[Löw67e] Per-Olov Löwdin. Nature of quantum chemistry. International Journal of Quantum Chemistry, 1(1):7–12, January 1967. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1967:Pa

[Löw67f] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 3:vii–viii, 1967. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600830.

Lowdin:1967:PIS

[Löw67g] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, held in honor of John Clarke Slater at Sanibel Island, Florida, January 16–20, 1967, volume 1 of International journal of quantum chemistry. Symposium. Interscience Publishers, New York, NY, USA, 1967. LCCN QD 462 A1 I58 1967.

Lowdin:1967:Pb

[Löw67h] Per-Olov Löwdin. Program. International Journal of Quantum Chemistry, 1(1):1–6, January 1967. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1967:QTT

[Löw67i] Per-Olov Löwdin. Quantum theory of time-dependent phenomena treated by the evolution operator technique. Advances in Quantum Chemistry, 3:323–381, 1967. CODEN AQCHA9. ISSN

0065-3276. URL http://www.sciencedirect.com/science/
article/pii/S0065327608600921. Earlier edition available as

technical report [Löw63f].

Lowdin:1967:SCT

[Löw67j] Per-Olov Löwdin. Some comments on the treatment of symmetry properties in perturbation theory. Preprint 131, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, 1967. 21 pp.

Lowdin:1967:SPH

[Löw67k] Per-Olov Löwdin. Some properties of the hydrogen bonds in biochemistry with particular reference to the stability of the genetic code. Pontificaiae Academiae Scientiarum Scripta Varia, 31(??): 637–??, 1967. CODEN ???? ISSN ???? Semaine d’Etude sur les Forces Moleculaires.

Lowdin:1968:SACa

[Löw68a] P.-O. Löwdin. Some aspects of the correlation problem and possible extensions of the Independent Particle Model. Technical Report TF 128, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, 1968. Published in [Löw68b].

Lowdin:1968:SACb

[Löw68b] P.-O. Löwdin. Some aspects of the correlation problem and possible extensions of the Independent Particle Model. Advances in Chemical Physics, 14:283–340, 1968. CODEN ADCPAA. ISSN 0065-2385 (print), 1934-4791 (electronic). Proceedings of Frascati Summer School on the Correlation Problem, 1967; Eds. R. Lefebvre and C. Moser, Interscience, 1968. Also available as [Löw68a].

Lowdin:1968:AQC

[Löw68c]  Per-Olov Löwdin, editor.  Advances in Quantum Chem-

istry, volume 4 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1968. CODEN AQCHA9. ISSN 0065-3276. xiv + 334 pp. LCCN QD453 .A28 v.4.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608603858.

Lowdin:1968:MAB

[Löw68d] Per-Olov Löwdin. Molecular associations in biology — A brief summary. In Bernard Pullman, editor, Molecular Associations in Biology: Proceedings of an international symposium held in celebration of the 40th anniversary of the Institut de biologie physicochimique (Foundation Edmond de Rothschild) Paris, May 8–11, 1967, pages 539–549. Academic Press Inc., New York, USA, 1968. LCCN QH506 .M718 1967.

Lowdin:1968:NOP

[Löw68e] Per-Olov Löwdin. On the non-orthogonality problem. Technical Report TF 146, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, June 1, 1968. 21 pp. Published in [Löw70b].

Lowdin:1968:P

[Löw68f] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 4:vii–viii, 1968. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608603883.

Lowdin:1968:PIS

[Löw68g] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology, held at Sanibel Island, Florida, January 15–20, 1968, volume 2 of International journal of quantum chemistry. Symposium. Interscience Publishers, New York, NY, USA, 1968. LCCN QD453.A1 I58 S95, 1968.

Lowdin:1968:SAP

[Löw68h] Per-Olov Löwdin. Some aspects on the possible importance of the reading mechanism of DNA in carcinogenesis. In ????, editor, Proceedings of the Israel Academy of Sciences Symposium on Physico-Chemical Mechanism of Carcinogenesis, Jerusalem, page ?? ????, ????, 1968. LCCN ????

Lowdin:1968:SAS

[Löw68i] Per-Olov Löwdin. Some aspects on the structure and present situation of quantum chemistry as expressed at the Girona Workshop. In Ramón Carbó, editor, Proc. Girona Workshop on Quantum Chemistry, pages 605–?? Elsevier, Amsterdam, The Netherlands, 1968. ISBN ???? LCCN ????

Lowdin:1968:SCT

[Löw68j] Per-Olov Löwdin. Some comments on the treatment of symmetry properties in perturbation theory. International Journal of Quantum Chemistry. Symposium, 2(S2):137–150, 1968. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1968:SPT

[Löw68k] Per-Olov Löwdin. Studies in perturbation theory. XIII. Treatment of constants of motion in resolvent method, partitioning technique, and perturbation theory. International Journal of Quantum Chemistry, 2(6):867–931, November 1968. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1968:UIP

[Löw68l] Per-Olov Löwdin. The use of inner projections and rational approximations in the future of quantum chemistry. In Ramón Carbó, editor, Proc. Girona Workshop on Quantum Chemistry, pages 1–?? Elsevier, Amsterdam, The Netherlands, 1968. ISBN ???? LCCN ????

Lowdin:1969:PCU

[Löw69a] P.-O. Löwdin. Perturbation calculation of the upper and lower limits of energy eigenvalues by partitioning technique. Magyar Fizikai Folyóirat, 17(6):515–540, 1969. CODEN MGFFAC. ISSN 0025-0104.

Lowdin:1969:PHF

[Löw69b] P.-O. Löwdin. The projected Hartree–Fock method for the extension of the independent particle model. Magyar Fizikai Folyóirat, 17(6):495–514, 1969. CODEN MGFFAC. ISSN 0025-0104.

Lowdin:1969:BRB

[Löw69c] Per-Olov Löwdin. Book review: Selected scientific papers of Egil A. Hylleraas. Edited by: John Midtdal, Harald Wergeland, and Knut Thalberg. Published by: NTH Press, Trondheim, Norway, 1968. Price: $27 for both volumes. International Journal of Quantum Chemistry, 3(2):246, March 1969. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [MTW68].

Lowdin:1969:I

[Low69d] Per-Olov Lowdin. Introduction. International Journal of Quantum Chemistry, 3(S3a):iv–x, January 13–18, 1969. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory and Quantum Biology.

Lowdin:1969:PIS

[Löw69e] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology: held in honor of Henry Eyring at Sanibel Island, Florida, January 13–18, 1969, volume 3 of International journal of quantum chemistry. Symposium. Interscience Publishers, New York, NY, USA, 1969. LCCN QD453.A1I62 no.3.

Lowdin:1969:SAH

[Löw69f] Per-Olov Löwdin. Some aspects of the hydrogen bond in molecular biology. Annals of the New York Academy of Sciences, 158(1):86– 95, May 16, 1969. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Lowdin:1969:SCP

[Löw69g] Per-Olov Löwdin. Some comments on the periodic system of the elements. International Journal of Quantum Chemistry, 3(S3a): 331–334, January 13–18, 1969. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory and Quantum Biology.

Lowdin:1970:AQC

[Löw70a] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 5 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1970. CODEN AQCHA9. ISBN 0-12-034805-5 (hardcover). ISSN 0065-3276. xv + 303 pp. LCCN QD453 .A28

v.5. URL http://www.sciencedirect.com/science/article/
pii/S0065327608603317;   http://www.sciencedirect.com/
science/book/9780120348053.

Lowdin:1970:NP

[Löw70b] Per-Olov Löwdin. On the nonorthogonality problem. Advances in Quantum Chemistry, 5:185–199, 1970. CODEN AQCHA9. ISSN

0065-3276. URL http://www.sciencedirect.com/science/
article/pii/S0065327608603391. Contribution to Professor

Ivar Waller’s 70-Year Anniversary Volume. Earlier edition available as technical report [Löw68e].

Lowdin:1970:P

[Löw70c] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 5:xi–xii, 1970. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608603342.

Lowdin:1970:SAR

[Löw70d] Per-Olov Löwdin. Some aspects on the research process in the natural sciences. In Tore Dalenius, Georg Karlsson, and Sten Malmquist, editors, Scientists at Work: Festschrift in honour of Herman Wold, pages 112–135. Almqvist and Wiksell, Uppsala, Sweden, 1970. LCCN Q180.5. .S36 1970.

Lowdin:1970:SPI

[Löw70e] Per-Olov Löwdin. Some properties of inner projections. International Journal of Quantum Chemistry, 5(S4):231–237, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1971:QCM

[Löw71a] P.-O. Löwdin. Quantum chemistry and molecular physics of solids. Colloq. Int. Cont. Nat. Rech. Sci., 195(??):207–226, 1971. CO-DEN ???? ISSN ????

Lowdin:1971:SCT

[Löw71b] P.-O. Löwdin. Some comments on the time-dependent variation principle. QTP notes. Published in [?]., December 1971.

Lowdin:1971:SPI

[Löw71c] P.-O. Löwdin. Some properties of inner projections. International Journal of Quantum Chemistry. Symposium, 4:231–237, 1971. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1971:TECa

[Löw71d] P.-O. Löwdin. Treatment of exchange and correlation effects in crystals. An introduction. QTP notes., December 1971.

Lowdin:1971:I

[Löw71e] Per-Olov Löwdin. Introduction. International Journal of Quantum Chemistry, 5(S5):xv–xvi, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1971:PISa

[Löw71f] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology: Held at Sanibel Island, Florida, January 19–24, 1970, volume 4 of International journal of quantum chemistry. Symposium. Interscience Publishers, New York, NY, USA, 1971. LCCN W1 IN790P no.4 1970.

Lowdin:1971:PISb

[Löw71g] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular Solid-State Theory and Quantum Biology: held in honor of John H. Van Vleck at Sanibel Island, Florida, January 18–23, 1971, volume 5 of International journal of quantum chemistry. Symposium. Interscience Publishers, New York, NY, USA, 1971. LCCN QD453.A1I62 no.5.

Lowdin:1972:STN

[Löw72a] P.-O. Löwdin. A stability theorem for normal operators of importance in quantum mechanics. Technical Report TF 268, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, February 1, 1972. 9 pp.

Lowdin:1972:AQC

[Löw72b] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 6 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1972. CODEN AQCHA9. ISBN 0-12-034806-3 (hardcover). ISSN 0065-3276. xvi + 413 pp. LCCN QD453 .A28

v.6. URL http://www.sciencedirect.com/science/article/
pii/S0065327608605377;   http://www.sciencedirect.com/
science/book/9780120348060.

Lowdin:1972:P

[Löw72c] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 6:xi–xii, 1972. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608605407.

Lowdin:1972:PIS

[Löw72d] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology: Held at Sanibel Island, Florida, January 17–22, 1972, volume 6 of International journal of quantum chemistry. Symposium. Interscience Publishers, New York, NY, USA, 1972. LCCN QD453.A1I62 no.6.

Lowdin:1972:RRU

[Löw72e] Per-Olov Löwdin. Recent research in the Uppsala & Florida Quantum Theory Projects. In ????, editor, Selected Topics in Molecular Physics, pages 47–83. Verlag Chemie GmbH, ????, 1972. LCCN ????

Lowdin:1972:TECa

[Löw72f] Per-Olov Löwdin. Treatment of exchange and correlation effects in crystals. An introduction. Report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, January 1972. 14 pp.

Lowdin:1972:TECb

[Löw72g] Per-Olov Löwdin. Treatment of exchange and correlation effects in crystals: Introductory remarks. In Herman et al. [HDK72], pages 191–204. ISBN 0-306-30583-6, 1-4684-1979-X (print), 1-4684-1977-3 (e-book). LCCN QC176-176.9. URL http://link.

springer.com/chapter/10.1007/978-1-4684-1977-1_16/.

Lowdin:1973:AQC

[Löw73a] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 7 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1973. CODEN AQCHA9. ISBN 0-12-034807-1 (hardcover). ISSN 0065-3276. xvi + 457 pp. LCCN QD453 .A28

v.7. URL http://www.sciencedirect.com/science/article/
pii/S0065327608605547;   http://www.sciencedirect.com/
science/book/9780120348077.

Lowdin:1973:EMT

[Löw73b] Per-Olov Löwdin. Electronic mobility and transfer of energy and momentum as time-dependent processes. In M. (Maurice) Marois, editor, From theoretical physics to biology, proceedings, Versailles, June 21–26, 1971, pages 145–146. Karger, Basel, Switzerland; New York, NY, USA, 1973. ISBN 3-8055-1578-2. LCCN QH505 .I591 1971; QT34 I59 1971.

Lowdin:1973:O

[Löw73c] Per-Olov Löwdin. Obituary. International Journal of Quantum Chemistry, 7(1):i, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1973:P

[Löw73d] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 7:xi–xii, 1973. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608605572.

Lowdin:1973:PIS

[Löw73e] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology: Held at Sanibel Island, Florida, January 21–27, 1973, volume 7 of International journal of quantum chemistry. Symposium. Interscience Publishers, New York, NY, USA, 1973. LCCN QD453.A1I62 no.7.

Lowdin:1974:AQC

[Löw74a] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 8 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1974. CODEN AQCHA9. ISBN 0-12-034808-X (hardcover). ISSN 0065-3276. xvii + 291 pp. LCCN QD453 .A28

v.8. URL http://www.sciencedirect.com/science/article/
pii/S0065327608600532;   http://www.sciencedirect.com/
science/book/9780120348084.

Lowdin:1974:JMJ

[Löw74b] Per-Olov Löwdin. Josef-Maria Jauch, in memoriam. Eur. Phys. News, 12(7):??, 1974. CODEN ???? ISSN ????

Lowdin:1974:MOH

[Löw74c] Per-Olov Löwdin. Mäniskan och hennes psyke i den moderna kvantteorins världsbild. (Swedish) [Man and his psyche in modern quantum theory’s world view]. Forskning och Praktik, 6(??):121– 125, 1974. CODEN ???? ISSN ????

Lowdin:1974:P

[Löw74d] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 8:xi–xii (??), 1974. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600568.

Lowdin:1974:PISa

[Löw74e] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Statistics, held at Sanibel Island, Florida, January 20–26, 1974, volume 8 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1974. CODEN IJQBDZ. ISSN 0360-8832. LCCN QD453.A1I62 no.8.

Lowdin:1974:PISb

[Löw74f] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology, held at Sanibel Island, Florida, January 17–19, 1974, volume 1 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1974. CODEN IJQBDZ. ISSN 0360-8832. LCCN QH 345 I61p 1974.

Lowdin:1975:SAOa

[Löw75a] P.-O. Löwdin. Some aspects on objectivity and reality in modern science. Technical Report TU 462, Kvantkemi, Upsala Universitet, Uppsala, Sweden, 1975. 59 pp. Also published as TF 427.

Lowdin:1975:EHF

[Löw75b] Per-Olov Löwdin. 2.2. an extension of the Hartree–Fock method to include correlation effects. In Mehra [Meh75], chapter 14, pages 360–361. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Lowdin:1975:AQC

[Löw75c] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 9 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1975. CODEN AQCHA9. ISBN 0-12-034809-8 (hardcover). ISSN 0065-3276. x + 304 pp. LCCN QD453 .A28

v.9. URL http://www.sciencedirect.com/science/article/
pii/S0065327608600349;   http://www.sciencedirect.com/
science/book/9780120348091.

Lowdin:1975:P

[Löw75d] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 9:ix–x, 1975. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600374.

Lowdin:1975:PISa

[Löw75e] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics, held at Sanibel Island, Florida, January 19–25, 1975, volume 9 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1975. CODEN IJQBDZ. ISSN 0360-8832. LCCN QD462 .I61 1975.

Lowdin:1975:PISb

[Löw75f] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology, held at Sanibel Island, Florida, January 16–18, 1975, volume 2 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1975. CODEN IJQBDZ. ISSN 0360-8832. LCCN QH 345 I609 1975.

Lowdin:1975:SAOb

[Löw75g] Per-Olov Löwdin. Some aspects on objectivity and reality in modern science. Lecture prepared for Greek Symposium on “The Case of Objectivity”., 1975.

Lowdin:1976:FQC

[Löw76a] P.-O. Löwdin. Future of quantum chemistry — A symposium in honor of Per-Olov Löwdin: A collection of short research summaries of participants. QTP notes., September 1976.

Lowdin:1976:SAA

[Löw76b] P.-O. Löwdin. Some aspects on the American-Swedish exchange in quantum sciences, particularly the Uppsala-Florida exchange project. Technical Report TU 470, Kvantkemi, Upsala Universitet, Uppsala, Sweden, 1976. 21 pp. Partners in Progress, A Chapter in the American-Swedish Exchange of Knowledge.

Lowdin:1976:IAV

[Löw76c] Per-Olov Löwdin. Internationella aktiviteter vid Kvantkemiska Institutionen och forskargruppen vid Uppsala Universitet. (Swedish)

[International activities at the Quantum Chemistry Institute and research group at Uppsala University]. Universen, Uppsala Universitet informerar, 1/76:??, 1976. CODEN ???? ISSN ????

Lowdin:1976:QCU

[Löw76d] Per-Olov Löwdin. Quantum chemistry at Uppsala University. Technical Report TN 439, Kvantkemi, Upsala Universitet, Uppsala, Sweden, 1976. ?? pp. Published in Uppsala University’s 500th Anniversary volume in math.-physical sciences.

Lowdin:1977:QTT

[Löw77a] P.-O. Löwdin. Quantum theory as a trace algebra. Technical Report TF 515, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32611-8435, USA, 1977. 138 pp.

Lowdin:1977:STL

[Löw77b] P.-O. Löwdin. Set theory and linear algebra — some mathematical tools to be used in quantum theory. Technical Report TN 490, Kvantkemi, Upsala Universitet, Uppsala, Sweden, 1977.

Lowdin:1977:AQC

[Löw77c] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 10 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1977. CODEN AQCHA9. ISBN 0-12-034810-1 (hardcover). ISSN 0065-3276. xii + 338 pp. LCCN QD453 .A28

v.10. URL http://www.sciencedirect.com/science/article/
pii/S0065327608605730;   http://www.sciencedirect.com/
science/book/9780120348107.

Lowdin:1977:KCO

[Löw77d] Per-Olov Löwdin. Kvantkemi, cancer och åldrande. (Swedish) [Quantum chemistry, cancer, and aging]. Läkartidning, 74(40): 3419–3421, October 5, 1977. CODEN LAKAA3. ISSN 0023-7205 (print), 1652-7518 (electronic).

Lowdin:1977:P

[Löw77e] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 10:xi–xii, 1977. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608605766.

Lowdin:1977:SAA

[Löw77f] Per-Olov Löwdin. Some aspects on the American–Swedish exchange in quantum sciences, particularly the Uppsala–Florida exchange project. In A. Kastrup and N. Olsson, editors, Partners in Progress: A Chapter in the American–Swedish Exchange of Knowledge, pages 255–276. Swedish Council of America, North Central Publishing Company, St. Paul, MN, 1977. ISBN ???? LCCN ????

Lowdin:1977:SAI

[Löw77g] Per-Olov Löwdin. Some aspects on the importance of basic research for human welfare. In W. O. (Winfred Oliver) Milligan, editor, American chemistry: bicentennial. Proceedings of the twentieth conference held at the Shamrock Hilton Hotel, Houston, Texas, November 8–10, 1976, volume 20 of Proceedings of the Robert A. Welch Foundation conferences on chemical research, pages 264– 269. Robert A. Welch Foundation, Houston, TX, USA, 1977. LCCN QD18.U6 A52 1977.

Lowdin:1977:SCM

[Löw77h] Per-Olov Löwdin. Some current models of carcinogenesis. International Journal of Quantum Chemistry, 12(S4):185–196, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1978:AQC

[Löw78a] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 11 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1978. CODEN AQCHA9. ISBN 0-12-034811-X (hardcover). ISSN 0065-3276. xii + 487 pp. LCCN QD453 .A28

v.11. URL http://www.sciencedirect.com/science/article/
pii/S0065327608602282;   http://www.sciencedirect.com/
science/book/9780120348114.

Lowdin:1978:I

[Löw78b] Per-Olov Löwdin. Introduction. International Journal of Quantum Chemistry, 14(4):iii, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1978:LTP

[Löw78c] Per-Olov Löwdin. “Low Temperature Physics and Cosmic Microwave Background Radiation”: the 1978 Nobel prizes in physics.

In Enciclopedia della Scienca e della Technica (EST) Yearbook, page ?? Ed. Arnoldo Mondadow, Milano, Italy, 1978. ISBN ???? LCCN ????

Lowdin:1978:P

[Löw78d] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 11:xi–xii, 1978. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608602312.

Lowdin:1978:QTT

[Löw78e] Per-Olov Löwdin. Quantum theory as a trace algebra. International Journal of Quantum Chemistry. Symposium, 12:197–266, 1978. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). Not in symposium volumes 11–13. Only references in Science Citation Index are to this paper, and [Löw82d]. See addendum comment [Löw81b].

Lowdin:1979:P

[Löw79] Per-Olov Löwdin. Preface. International Journal of Quantum Chemistry, 16(1):1–2, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1980:PIC

[Löw80a] P.-O. Löwdin. Proceedings of the 3rd International Congress of Quantum Chemistry — Contributed Papers. 1. Held at Kyoto, Japan, 29 October 3 November 1979. International Journal of Quantum Chemistry, 18(1):1, ???? 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1980:AQC

[Löw80b] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 12 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1980. CODEN AQCHA9. ISBN 0-12-034812-8 (hardcover). ISSN 0065-3276. x + 325 pp. LCCN QD453 .A28

v.12. URL http://www.sciencedirect.com/science/article/
pii/S0065327608603081;   http://www.sciencedirect.com/
science/book/9780120348121.

Lowdin:1980:IRN

[Löw80c] Per-Olov Löwdin. Introductory remarks to Nobel Symposium on Many-Body Theory of Atomic Systems. Physica Scripta, 21(3–4): 229–234, ???? 1980. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Lowdin:1980:MSCa

[Löw80d] Per-Olov Löwdin. Molecular structure calculations. Advances in Quantum Chemistry, 12:263–316, 1980. CODEN AQCHA9. ISSN

0065-3276. URL http://www.sciencedirect.com/science/
article/pii/S0065327608603184.

Lowdin:1980:MSCb

[Löw80e] Per-Olov Löwdin. Molecular structure calculations — A brief review. In Rita G. Lerner and George L. Trigg, editors, Encyclopedia of Physics, page ?? Addison-Wesley, Reading, MA, USA, 1980. ISBN 0-201-04313-0. LCCN QC5 .E495 1980. Foreword by Walter Sullivan.

Lowdin:1980:OAI

[Löw80f] Per-Olov Löwdin. Opening address at the 3rd International Conference in Quantum Chemistry in Kyoto. In Kenichi Fukui and Bernard Pullman, editors, Horizons of quantum chemistry: proceedings of the third International Congress of Quantum Chemistry, held at Kyoto, Japan, October 29–November 3, 1979, pages xv–xviii. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1980. ISBN 90-277-1105-4. LCCN QD462.A1 I49 1979.

Lowdin:1980:Pa

[Löw80g] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 12:ix–x, 1980. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608603111.

Lowdin:1980:Pb

[Löw80h] Per-Olov Löwdin. Preface. International Journal of Quantum Chemistry, 18(1):1, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1980:SCP

[Löw80i] Per-Olov Löwdin. The size-consistency problem for large molecules. In Kenichi Fukui and Bernard Pullman, editors, Horizons of quantum chemistry: proceedings of the third International Congress of Quantum Chemistry, held at Kyoto, Japan, October 29–November 3, 1979, pages 193–195. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1980. ISBN 90-277-1105-4. LCCN QD462.A1 I49 1979.

Lowdin:1980:SAH

[Löw80j] Per-Olov Löwdin. Some aspects on the human mind in view of modern quantum chemistry. In ????, editor, Proc. Eur. Acad. of Arts, Sciences, and Humanities, pages 102–?? UNESCO, Paris, France, 1980. ISBN ???? LCCN ????

Lowdin:1981:PIC

[Löw81a] P.-O. Löwdin. Proceedings of the International Colloquium on the Nature of Binding in Molecular Associations, Atom Clusters, and Surface Phases Held a Sorrento, Italy 12–16 May 1980 — Preface. International Journal of Quantum Chemistry, 19(6):973, ???? 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1981:AQT

[Löw81b] Per-Olov Löwdin. Addendum to ‘Quantum theory as a trace algebra’. International Journal of Quantum Chemistry, 20(3):775–776, ???? 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1981:AQCa

[Löw81c] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 13 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1981. CODEN AQCHA9. ISBN 0-12-034813-6 (hardcover). ISSN 0065-3276. x + 251 pp. LCCN QD453 .A28

v.13. URL http://www.sciencedirect.com/science/article/
pii/S0065327608602877;   http://www.sciencedirect.com/
science/book/9780120348138.

Lowdin:1981:AQCb

[Löw81d] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 14 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1981. CODEN AQCHA9. ISBN 0-12-034814-4 (hardcover). ISSN 0065-3276. x + 248 pp. LCCN QD453 .A28

v.14. URL http://www.sciencedirect.com/science/article/
pii/S0065327608603214;   http://www.sciencedirect.com/
science/book/9780120348145.

Lowdin:1981:Pa

[Löw81e] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 13:ix–x, 1981. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608602907.

Lowdin:1981:Pb

[Löw81f] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 14:ix–x, 1981. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S006532760860324X.

Lowdin:1981:Pc

[Löw81g] Per-Olov Löwdin. Preface. International Journal of Quantum Chemistry, 19(6):973, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:AQCa

[Löw82a] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 15 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1982. CODEN AQCHA9. ISBN 0-12-034815-2 (hardcover). ISSN 0065-3276. x + 277 pp. LCCN QD453 .A28

v.15. URL http://www.sciencedirect.com/science/article/
pii/S0065327608601173;   http://www.sciencedirect.com/
science/book/9780120348152.

Lowdin:1982:AQCb

[Löw82b] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 16 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1982. CODEN AQCHA9. ISBN 0-12-034816-0 (hardcover). ISSN 0065-3276. x + 285 pp. LCCN QD453 .A28

v.16. URL http://www.sciencedirect.com/science/article/
pii/S0065327608603469;   http://www.sciencedirect.com/
science/book/9780120348169.

Lowdin:1982:OSH

[Löw82c] Per-Olov Löwdin. On operators, superoperators, Hamiltonians, and Liouvillians. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 22(S16):485–560, March 1–13, 1982. CODEN IJQSDI. ISSN 0161-3642. Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Lowdin:1982:AQT

[Löw82d] Per-Olov Löwdin. On the axioms of quantum theory formulated as a trace algebra. International Journal of Quantum Chemistry, 21(1):275–353, ???? 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:PTP

[Löw82e] Per-Olov Löwdin. Partitioning technique, perturbation theory, and rational approximations. International Journal of Quantum Chemistry, 21(1):69–92, ???? 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:Pa

[Löw82f] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 15:ix–x, 1982. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608601203.

Lowdin:1982:Pb

[Löw82g] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 16:ix–x, 1982. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608603494.

Lowdin:1983:AQM

[Löw83a] Per-Olov Löwdin. Advances in quantum molecular sciences — A tribute to Kenichi Fukui. Journal of Molecular Structure, 103: 3–24, August 1983. CODEN JMOSB4. ISSN 0022-2860 (print), 1872-8014 (electronic). Also as Theochem, vol. 12.

Lowdin:1983:SCE

[Löw83b] Per-Olov Löwdin. On the Sanibel coefficients in the expansion of spin-projected Slater determinants. International Journal of Quantum Chemistry, 24(6):729–745, ???? 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:SPP

[Löw83c] Per-Olov Löwdin. On the stability problem of a pair of adjoint operators. Journal of Mathematical Physics, 24(1):70–87, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v24/i1/p70_s1.

Lowdin:1983:TBM

[Löw83d] Per-Olov Löwdin. On the theoretical background of molecular biology. In Enrico Clementi and Ramaswamy H. Sarma, editors, Structure and dynamics, nucleic acids and proteins: proceedings of the International Symposium on Structure and Dynamics of Nucleic Acids and Proteins, held at the University of California, San Diego at La Jolla, September 5–9, 1982, under the auspices of the International Society of Quantum Biology, pages 13–53. Adenine Press, Guilderland, NY, USA, 1983. ISBN 0-940030-04-7. LCCN QP620 .I57 1982; QU58 I576 1983.

Lowdin:1983:SRD

[Löw83e] Per-Olov Löwdin. Some recent developments in theoretical molecular biology in relation to some general principles in physics and chemistry. In Tor Stubb, editor, Proceedings of the 60th Anniversary of the Swedish Academy of Technical Sciences, Helsinki, Finland, page ?? ????, ????, 1983. ISBN ???? LCCN ????

Lowdin:1984:SAR

[Löw84] Per-Olov Löwdin. Some aspects on the relation between the natural sciences and the human mind — an introduction to a panel discussion on the “origin of life and mind”. International Journal of Quantum Chemistry, 26(S11):31–43, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1985:YSS

[Löw85a] P.-O. Löwdin. 25 years of Sanibel Symposia — a brief historic and scientific survey. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 19:19–37, 1985. CODEN IJQSDI. ISSN 0161-3642.

Lowdin:1985:AQC

[Löw85b] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 17 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1985. CODEN AQCHA9. ISBN 0-12-034817-9 (hardcover). ISSN 0065-3276. x + 344 pp. LCCN QD453 .A28

v.17. URL http://www.sciencedirect.com/science/article/
pii/S006532760860297X;   http://www.sciencedirect.com/
science/book/9780120348176.

Lowdin:1985:ACL

[Löw85c] Per-Olov Löwdin. Approximate calculation of life-times of resonance states in the continuum from real stabilization graphs. International Journal of Quantum Chemistry, 27(4):495–500, ???? 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1985:CSA

[Löw85d] Per-Olov Löwdin. On the change of spectra associated with unbounded similarity transformations of a many-particle Hamiltonian and the occurrence of resonance states in the method of complex scaling. Part I: General theory. Technical Report TN 803, Kvantkemi, Upsala Universitet, Uppsala, Sweden, 1985. Published in [Löw88e].

Lowdin:1985:CBR

[Löw85e] Per-Olov Löwdin. On the connection between the resolvent methods and partitioning technique, rational approximations, and perturbation theory. Physica Scripta, 32(4):261–262, ???? 1985. CO-DEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Lowdin:1985:P

[Löw85f] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 17:ix–x, 1985. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608603007.

Lowdin:1985:SAH

[Löw85g] Per-Olov Löwdin. Some aspects on the Hamiltonian and Liouvillian formalism, the special propagator methods, and the equation of motion approach. Advances in Quantum Chemistry, 17:285–334, 1985. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608603056.

Lowdin:1985:SCE

[Löw85h] Per-Olov Löwdin. Some comments on the ethic and moral responsibilities of the research scientists. In ????, editor, Proc. UNESCO Conf., Bucharest, Romania, Dec. 1985, page ?? UNESCO, ????, 1985. ISBN ???? LCCN ????

Lowdin:1985:TFY

[Löw85i] Per-Olov Löwdin. Twenty-five years of Sanibel Symposia: A brief historic and scientific survey. International Journal of Quantum Chemistry, 28(S19):19–37, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1986:AQC

[Löw86a] Per-Olov Löwdin, editor. Advances in Quantum Chemistry, volume 18 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1986. CODEN AQCHA9. ISBN 0-12-034818-7 (hardcover). ISSN 0065-3276. viii + 350 pp. LCCN QD453 .A28

v.18. URL http://www.sciencedirect.com/science/article/
pii/S0065327608600441;   http://www.sciencedirect.com/
science/book/9780120348183.

Lowdin:1986:A

[Löw86b] Per-Olov Löwdin. Announcement. International Journal of Quantum Chemistry, 29(2):vii–viii, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1986:CSM

[Löw86c] Per-Olov Löwdin. Comments on some models of carcinogenesis. International Journal of Quantum Chemistry, 30(S13):311–317, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1986:GBT

[Löw86d] Per-Olov Löwdin. A generalization of Brillouin’s theorem and the stability conditions in the quantum-mechanical variation principle in the case of general trial wave functions. Proceedings of the Indian Academy of Sciences. Chemical sciences., 96(3–4):121–126, ???? 1986. CODEN PIAADM. ISSN 0253-4134. Basu Dedicatory Volume, Ed. D. Mukherjee.

Lowdin:1986:MDM

[Löw86e] Per-Olov Löwdin. The mathematical definition of a molecule and molecular structure. In Jean Maruani, editor, Molecules in physics, chemistry, and biology, (Paris, 1986), volume 2 (Physical aspects of molecular systems) of Topics in molecular organization and engineering, pages 3–60. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1986. ISBN 90-277-2596-9 (vol. 1); 90-277-2617-5 (set: vol. 1–3).

Lowdin:1986:NDQ

[Löw86f] Per-Olov Löwdin. New directions in quantum chemical calculations — particularly as to new materials. In Larry L. Hench and Donald R. Ulrich, editors, Science of Ceramic Chemical Processing: Proceedings of the Second International Conference on Ultrastructure Processing of Ceramics, Glasses, and Composites, held February 25–March 1, 1985 in Palm Coast, Florida, pages 577–?? John Wiley and Sons, Inc., New York, NY, USA, 1986. ISBN 0-471-82645-6. LCCN TP785 .U46 1985. US$54.00.

Lowdin:1986:SAQa

[Löw86g] Per-Olov Löwdin. On the state of the art of quantum chemistry. International Journal of Quantum Chemistry, 29(5):1651–1683, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Proceedings of the 5th International Congress of Quantum Chemistry (ICQC), August 18–24, 1985, Montréal, Canada, Part IV.

Lowdin:1986:P

[Löw86h] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 18:vii–viii, 1986. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600465.

Lowdin:1986:RAW

[Löw86i] Per-Olov Löwdin. Remarks about the wave-function representability of the first-order reduced density matrices. International Journal of Quantum Chemistry, 30(S20):769–771, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Lowdin:1986:SAQb

[Löw86j] Per-Olov Löwdin. Some aspects on quantum theory, consciousness of the mind and free will. In Sidney W. Fox, editor, Self-Organization: Principles in Science, Events in Society. Sonesta Beach Hotel, Key Biscayne, Florida, November 1–3, 1984, pages 1–3. Adenine Press, Guilderland, NY, USA, 1986. ISBN 0-940030-13-6. LCCN ????

Lowdin:1986:SCM

[Löw86k] Per-Olov Löwdin. Some comments on the method of complex scaling to find physical resonance states. International Journal of Quantum Chemistry, 30(S20):743–748, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Lowdin:1986:SCR

[Löw86l] Per-Olov Löwdin. Some comments on the reduction of the natural sciences and on the relation between mind and matter and between natural and artificial intelligence. In R. Daudel, editor, Proc. 5th Int. Conf. “Innovations and Society”, Dec. 1985, Bucharest, Romania, page ?? ????, Paris, France, 1986. ISBN ???? LCCN ????

Lowdin:1987:CCQ

[Löw87a] Per-Olov Löwdin. Citation classic — Quantum Theory of Many-Particle Systems. 1. Physical Interpretations by Means of Density-Matrices, Natural Spin-Orbitals, and Convergence Problems in the Method of Configurational Interaction, Phys. Rev. 97:1474– 1489, 1955. Current Contents: Physical, Chemical & Earth Sciences, 40(??):20, October 5, 1987. ISSN 0163-2574. URL http://

garfield.library.upenn.edu/classics1987/A1987K083700001.

pdf. Letter dated 30 June 1987.

Lowdin:1987:SAD

[Löw87b] Per-Olov Löwdin. Some aspects on the development of the theory of reduced density matrices and the representability problem. In Robert Erdahl and Vedene H. Smith, Jr., editors, Density Matrices and Density Functionals, pages 21–50. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1987. ISBN 90-277-2477-6. LCCN QC174.17.D44 D471 1987. Proceedings of a three day conference, held in August 1985 and hosted jointly by the Departments of Chemistry, and Mathematics and Statistics of Queen’s University.

Lowdin:1987:SAH

[Löw87c] Per-Olov Löwdin. Some aspects on the history of computational quantum chemistry — in view of the development of the supercomputers and large-scale parallel computers. In M. (Michel) Dupuis, editor, Supercomputer Simulations in Chemistry: proceedings of the Symposium on Supercomputer Simulations in Chemistry, held in Montreal, August 25–27, 1985 / sponsored by IBM-Kingston and IBM-Canada, volume 44 of Lecture Notes in Chemistry, pages 1–48. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-17178-9. LCCN QD462.A1 S98 1985.

Lowdin:1987:TJC

[Löw87d] Per-Olov Löwdin. A tribute to A. John Coleman — the “tame” mathematician. In Robert Erdahl and Vedene H. Smith, Jr., editors, Density Matrices and Density Functionals, pages 1–4. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1987. ISBN 90-277-2477-6. LCCN QC174.17.D44 D471 1987. Proceedings of a three day conference, held in August 1985 and hosted jointly by the Departments of Chemistry, and Mathematics and Statistics of Queen’s University.

Lowdin:1988:CCP

[Löw88a] P.-O. Löwdin. Call for contributions to paper symposium on the Pariser–Parr–Pople method. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 15:vi, 1988. CODEN IJQBDZ. ISSN 0360-8832.

Lowdin:1988:ANF

[Löw88b] Per-Olov Löwdin. Announcement of new format. Advances in Quantum Chemistry, 19:xiii, 1988. CODEN AQCHA9. ISSN

0065-3276. URL http://www.sciencedirect.com/science/
article/pii/S0065327608606127.

Lowdin:1988:IGQ

[Löw88c] Per-Olov Löwdin. The importance of general quantum statistics and the Liouville equation in physical chemistry — some recent trends in quantum chemistry. Zeitschrift für physikalische Chemie (Leipzig), 269(2):284–305, ???? 1988. CODEN ZPCLAH. ISSN 0323-4479.

Lowdin:1988:STI

[Löw88d] Per-Olov Löwdin. On some thermodynamical inequalities. International Journal of Quantum Chemistry, 34(S22):337–342, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Lowdin:1988:CSA

[Löw88e] Per-Olov Löwdin. On the change of spectra associated with unbounded similarity transformations of a many-particle Hamiltonian and the occurrence of resonance states in the method of complex scaling. Part I. General theory. Advances in Quantum Chemistry, 19:87–138, 1988. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608606140.

Lowdin:1988:P

[Löw88f] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 19:xi–xii, 1988. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608606115.

Lowdin:1988:PISb

[Löw88g] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at the University of Florida Whitney Marine Laboratory at Marineland, Florida, March 12–19, 1988, volume 15 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1988. CODEN IJQBDZ. ISBN 0-471-50693-1. ISSN 0360-8832. LCCN QH345.I625 1988.

Lowdin:1988:PISa

[Löw88h] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods: held at the University of Florida Whitney Marine Laboratory at Marineland, Florida, March 12–19, 1988, volume 22 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1988. CODEN IJQBDZ. ISBN 0-471-50692-3. ISSN 0360-8832. LCCN QD462 .I61 1988.

Lowdin:1988:RFE

[Löw88i] Per-Olov Löwdin. The role of free energy, entropy, and similar concepts in the future of quantum chemistry. In K. Fukui and T. Yamabe, editors, Proc. Kyoto Inauguration Institute for Fundamental Chemistry, page ?? ????, Kyoto, Japan, 1988. ISBN ???? LCCN ????

Lowdin:1988:SAS

[Löw88j] Per-Olov Löwdin. Some aspects on the structure and present situation of quantum chemistry as expressed as the Girona workshop. In Carbó [Car88], pages 605–622. ISBN 0-444-87494-1. LCCN QD450.S7 no.62; QD462.A1 I58 1988. NL FL375.

Lowdin:1988:RUP

[Löw88k] Per-Olov Löwdin. To reach for the unreachable — psi and gamma. Journal of Molecular Structure, 165(3–4):177–188, May 1988. CO-DEN JMOSB4. ISSN 0022-2860 (print), 1872-8014 (electronic). Also as Theochem, vol. 42, no. 3–4.

Lowdin:1988:UIP

[Löw88l] Per-Olov Löwdin. The use of inner projections and rational approximation in the future of quantum chemistry. In Carbó [Car88], pages 1–24. ISBN 0-444-87494-1. LCCN QD450.S7 no.62; QD462.A1 I58 1988. NL FL375.

Lowdin:1989:PIC

[Löw89a] P.-O. Löwdin. Proceedings of the 6th International Congress on Quantum Chemistry, Hebrew University, Jerusalem, Israel, 22–25 August 1988: Editor’s Preface. International Journal of Quantum Chemistry, 36(3):189–190, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1989:SCP

[Löw89b] P.-O. Löwdin. Some comments on the present situation of quantum chemistry in view of the discussions at the Dubrovnik workshop on the electronic structure of molecules. Pure and Applied Chemistry, 61(12):2185–2194, December 1989. CODEN PACHAS. ISSN 0033-4545 (print), 1365-3075 (electronic). Plenary and invited lectures presented at the International Symposium on the Electronic Structure and Properties of Molecules and Crystals held in Dubrovnik-Cavtat, Yugoslavia, 29 August – 3 September, 1988, ed. Z. B. Maksic.

Lowdin:1989:LWC

[Löw89c] Per-Olov Löwdin. On the long way from the Coulombic Hamiltonian to the electronic structure of molecules. Pure and Applied Chemistry, 61(12):2065–2074, December 1989. CODEN PACHAS. ISSN 0033-4545 (print), 1365-3075 (electronic). Plenary and invited lectures presented at the International Symposium on the Electronic Structure and Properties of Molecules and Crystals held in Dubrovnik-Cavtat, Yugoslavia, 29 August – 3 September, 1988, ed. Z. B. Maksic.

Lowdin:1989:PSQ

[Löw89d] Per-Olov Löwdin. On the present situation of quantum chemistry: Chairman’s introductory remarks at the 6th International Congress of Quantum Chemistry (ICQC). International Journal of Quantum Chemistry, 36(3):191–199, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Proc. 6th International Congress in Quantum Chemistry, in Jerusalem, Aug., 1988.

Lowdin:1989:P

[Löw89e] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 20:xi–xii, 1989. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608606243.

Lowdin:1989:PISb

[Löw89f] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at St. Augustine, Florida, April 1–8, 1989, volume 16 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1989. CODEN IJQBDZ. ISBN 0-471-52598-7. ISSN 0360-8832. LCCN QH345.I625 1989.

Lowdin:1989:RFE

[Löw89g] Per-Olov Löwdin. Role of free energy in future development of the quantum theory of microelectronics. Acta Polytechnica Scandinavica, Electrical Engineering Series, 64(El64):398–403, ???? 1989. CODEN APSEA5. ISSN 0001-6845. Proc. Helsinki Summer Institute on Micro-electronics, Ed. J. Sinkkonen.

Lowdin:1989:SRD

[Löw89h] Per-Olov Löwdin. Some recent developments in theoretical organic chemistry. In Imre G. Csizmadia, editor, WATOC Annual Report 1989, page ?? ????, Toronto, Ontario, Canada, 1989. ISBN ???? LCCN ????

Lowdin:1990:Eb

[Löw90a] P.-O. Löwdin. Editorial. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 24:xi–xii, 1990. CODEN IJQSDI. ISSN 0161-3642.

Lowdin:1990:CCM

[Löw90b] Per-Olov Löwdin. Concepts of convergence in mathematical chemistry. Journal of Mathematical Chemistry, 4(1):401–414, December 1990. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-

8897 (electronic). URL http://link.springer.com/article/

10.1007/BF01170023. Proc. Galveston Conference on Mathematical Chemistry. Accepted for publication.

Lowdin:1990:Ea

[Löw90c] Per-Olov Löwdin. Editorial. International Journal of Quantum Chemistry, 38(S17):v–vi, March 17–24, 1990. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1990:Ec

[Löw90d] Per-Olov Löwdin. Editorial. International Journal of Quantum Chemistry, 38(S17):V–VI, 1990. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1990:IJQ

[Löw90e] Per-Olov Löwdin. International Journal of Quantum Chemistry A journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology; editorial-program and policies. International Journal of Quantum Chemistry, 38 (S24):xi–xii, March 17–24, 1990. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Physics, and Computational Methods.

Lowdin:1990:I

[Löw90f] Per-Olov Löwdin. Introduction. International Journal of Quantum Chemistry, 37(4):317, ???? 1990. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1990:MCM

[Löw90g] Per-Olov Löwdin. Mathematics in chemistry and in mathematical chemistry. Journal of Mathematical Chemistry, 4(1):3–14, December 1990. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-

8897 (electronic). URL http://link.springer.com/article/

10.1007/BF01170000. Proc. Galveston Conference on Mathematical Chemistry. Accepted for publication.

Lowdin:1990:IQC

[Löw90h] Per-Olov Löwdin. On the importance of quantum chemistry in the future of chemistry. In Mu Shik John, editor, Proc. Second EurAsia Conference on Chemistry in Seoul, April 16–20, 1990, pages 3–?? KAIST, Seoul, Korea, 1990. ISBN ???? LCCN ????

Lowdin:1990:LWC

[Löw90i] Per-Olov Löwdin. On the long way from the Coulombic Hamiltonian to the intermolecular energy surfaces: Concluding remarks at the Tromsø EUCHEM Conference, June 20–22, 1989. International Journal of Quantum Chemistry, 38(5):741–758, ???? 1990. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1990:PISb

[Löw90j] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at the Ponce de Leon Conference Center, St. Augustine, Florida, March 17–24, 1990, volume 17 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1990. CODEN IJQBDZ. ISBN 0-471-54598-8. ISSN 0360-8832. LCCN QH345.I53 1990.

Lowdin:1990:PISa

[Löw90k] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Quantum Chemistry, Solid-State Physics, and Computational Methods: held at the Ponce de Leon Conference Center, St. Augustine, Florida, March 17–24, 1990, volume 24 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1990. CODEN IJQBDZ. ISBN 0-471-54599-6. ISSN 0360-8832. LCCN QD462 .I61 1990.

Lowdin:1991:E

[Löw91a] P. O. Löwdin. Editorial. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 25:ix, 1991. CODEN IJQSDI. ISSN 0161-3642.

Lowdin:1991:IHM

[Löw91b] P.-O. Löwdin. Issue in honor of Michael Kasha — introduction. International Journal of Quantum Chemistry, 39(3):231, ???? 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1991:CRS

[Löw91c] Per-Olov Löwdin. Concluding remarks at the Second WATOC World Congress in Toronto, July 8–14, 1991. Journal of Molecular Structure, 233:353–355, September 1991. CODEN JMOSB4. ISSN 0022-2860 (print), 1872-8014 (electronic). Also as Theochem, vol. 79.

Lowdin:1991:IJQa

[Löw91d] Per-Olov Löwdin. International Journal of Quantum Chemistry — A journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. editorial — program and policies. International Journal of Quantum Chemistry, 39(1): 1–2, 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1991:IJQb

[Löw91e] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. editorial program and policies. International Journal of Quantum Chemistry, 40 (1):1–2, ???? 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1991:IJQc

[Löw91f] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology: Editorial program and policies. International Journal of Quantum Chemistry, 40 (S18):v–vi, ???? 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceeding of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1991:IJQd

[Löw91g] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. editorial program and policies. International Journal of Quantum Chemistry, 40 (S25):ix–x, ???? 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceeding of the International Symposium on Quantum Chemistry Solid-State Physics, and Computational Methods.

Lowdin:1991:MSC

[Löw91h] Per-Olov Löwdin. Molecular structure calculations — a review. In Rita G. Lerner and George L. Trigg, editors, Encyclopedia of Physics, pages 753–761. VCH Publishers, New York, NY, USA, second edition, 1991. ISBN 0-89573-752-3. LCCN QC5 .E545 1991.

Lowdin:1991:OOA

[Löw91i] Per-Olov Löwdin. Om osäkerhetsrelationerna i akustiken och deras betydelse vid definitionen av begreppet ton. (Swedish) [On uncertainty relations in acoustics and their meaning by the definition of the concept of tone]. Invited paper at Colloquium on “Quid est Sonus - Quid est Tonus” at the Swedish Royal Academy of Music, Stockholm., February 1991.

Lowdin:1991:NMD

[Löw91j] Per-Olov Löwdin. On nuclear motion and the definition of molecular structure. Journal of Molecular Structure, 230:13–15, May 1991. CODEN JMOSB4. ISSN 0022-2860 (print), 1872-8014 (electronic). Also as Theochem, vol. 76.

Lowdin:1991:HDV

[Löw91k] Per-Olov Löwdin. On the historical development of the valence bond method and the non-orthogonality problem. Journal of Molecular Structure, 229:1–14, May 6, 1991. CODEN JMOSB4. ISSN 0022-2860 (print), 1872-8014 (electronic). Also as Theochem, vol. 75.

Lowdin:1991:ITF

[Löw91l] Per-Olov Löwdin. On the importance of theory in the future development of chemistry. Journal of Molecular Structure, 230:1–3, May 24, 1991. CODEN JMOSB4. ISSN 0022-2860 (print), 1872-8014 (electronic). Also as Theochem, vol. 76. Proceedings of the 2nd World Condensed of Theoretical Organic Chemistry (WA-TOC), July 9–14, 1990, University of Toronto, Toronto, Ontario, Canada.

Lowdin:1991:OTS

[Löw91m] Per-Olov Löwdin. On the origin of the tone system and the theory of harmony. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 135(4): 533–556, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.org/

stable/986815.

Lowdin:1991:PTE

[Löw91n] Per-Olov Löwdin. On the pairing theorem and its extension. Israel Journal of Chemistry, 31(4):297–302, 1991. CODEN ISJCAT. ISSN 0021-2148. Ruben Pauncz dedicatory volume.

Lowdin:1991:P

[Löw91o] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 22:ix–x, 1991. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608603603.

Lowdin:1991:PISb

[Löw91p] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at the Ponce de Leon Conference Center, St. Augustine, Florida, March 9–16, 1991, volume 18 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1991. CODEN IJQBDZ. ISBN 0-471-57348-5. ISSN 0360-8832. LCCN QH301.I52 1991.

Lowdin:1991:PISa

[Löw91q] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Quantum Chemistry, Solid-State Physics, and Computational Methods: held at the Ponce de Leon Conference Center, St. Augustine, Florida, March 9–16, 1991, volume 25 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1991. CODEN IJQBDZ. ISBN 0-471-57349-3. ISSN 0360-8832. LCCN QD462 .I61 no.25 1991.

Lowdin:1991:SCP

[Löw91r] Per-Olov Löwdin. Some current problems in quantum chemistry as described by their key words and the International Journal of Quantum Chemistry. International Journal of Quantum Chemistry, 39(1):3–10, January 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1991:TOH

[Löw91s] Per-Olov Löwdin. Toner och harmoni: Något om tonsystemets uppkomst och harmonilärans grunder. (Swedish) [Tone and harmony: Something about the tone systems development and the foundations of harmony]. Proc. of the Royal Academy of Arts and Sciences in Uppsala, 28(??):33–??, 1991. CODEN ???? ISSN ????

Lowdin:1992:E

[Löw92a] P.-O. Löwdin. Editorial. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 19:v, 1992. CODEN IJQBDZ. ISSN 0360-8832.

Lowdin:1992:EPP

[Löw92b] P.-O. Löwdin. Editorial — program and policies. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 19:v–vi, 1992. CODEN IJQBDZ. ISSN 0360-8832.

Lowdin:1992:IJQb

[Löw92c] P.-O. Löwdin. International journal of quantum chemistry — A journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology — introduction. International Journal of Quantum Chemistry, 41(1):5, ???? 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:SIH

[Löw92d] P.-O. Löwdin. Special issue in honor of Enrico Clementi. 1. introduction. International Journal of Quantum Chemistry, 42(4): 525–526, ???? 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:IJQa

[Löw92e] Per-Olov Löwdin. International journal of quantum chemistry — A journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology — editorial program and policies. International Journal of Quantum Chemistry, 41 (1):1–2, ???? 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:IJQe

[Löw92f] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. International Journal of Quantum Chemistry, 44(S19):v–vi, March 14–21, 1992. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on the Application of Fundamental Theory to Problems of Biology and Pharmacology.

Lowdin:1992:IJQc

[Löw92g] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. editorial program and policies. International Journal of Quantum Chemistry, 42(1): 1–2, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:IJQd

[Löw92h] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. editorial program and policies. International Journal of Quantum Chemistry, 43(1): 3–4, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:Ia

[Löw92i] Per-Olov Löwdin. Introduction. International Journal of Quantum Chemistry, 41(1):5, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:Ib

[Löw92j] Per-Olov Löwdin. Introduction. International Journal of Quantum Chemistry, 42(4):525–526, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:BAI

[Löw92k] Per-Olov Löwdin. On Boolean algebra and its importance for the computer sciences. International Journal of Quantum Chemistry, 42(4):719–726, May 20, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:LAL

[Löw92l] Per-Olov Löwdin. On linear algebra, the least square method, and the search for linear relations by regression analysis in quantum chemistry and other sciences. Advances in Quantum Chemistry, 23:83–126, 1992. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600295.

Lowdin:1992:RSC

[Löw92m] Per-Olov Löwdin. On rotations as special cases of unitary transformations with some applications to the theory of spin. International Journal of Quantum Chemistry, 41(1):213–242, January 5, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:Pa

[Löw92n] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 23:ix–x, 1992. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600271.

Lowdin:1992:Pb

[Löw92o] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 24:ix–x, 1992. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600994.

Lowdin:1992:PISa

[Löw92p] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods: held at the Ponce de Leon Conference Center, St. Augustine, Florida, March 14–21, 1992, volume 26 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1992. CODEN IJQBDZ. ISBN 0-471-59463-6. ISSN 0360-8832. LCCN QD462.A1 I52 no.26.

Lowdin:1992:PISb

[Löw92q] Per-Olov Löwdin, editor. Proceedings of the International Symposium on the Application of Fundamental Theory to Problems of Biology and Pharmacology: held at the Ponce de Leon Conference Center, St. Augustine, Florida, March 14–21, 1992, volume 19 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1992. CODEN IJQBDZ. ISBN 0-471-59454-7. ISSN 0360-8832. LCCN QH345.I54 1992.

Lowdin:1992:SAO

[Löw92r] Per-Olov Löwdin. Some aspects of objectivity and reality in modern science. Foundations of Physics, 22(1):41–105, January 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). In honor of Sir Karl Popper.

Lowdin:1993:Ea

[Löw93a] P.-O. Löwdin. Editorial. International Journal of Quantum Chemistry, 45(1):1–??, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1993:Eb

[Löw93b] P.-O. Löwdin. Editorial. International Journal of Quantum Chemistry, 47(1):1–??, July 5, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1993:IJQa

[Löw93c] Per-Olov Löwdin. International Journal of Quantum Chemistry — A Journal Devoted to Quantum Theory and Computations in Chemistry, Condensed Matter Physics, and Biology: Editorial — program and policies. International Journal of Quantum Chemistry, 45(1):1–2, ???? 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1993:IJQb

[Löw93d] Per-Olov Löwdin. International Journal of Quantum Chemistry — A Journal Devoted to Quantum Theory and Computations in Chemistry, Condensed Matter Physics, and Biology: Editorial — program and policies. International Journal of Quantum Chemistry, 47(1):1–2, July 5, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1993:OIT

[Löw93e] Per-Olov Löwdin. On operator inequalities as tools in quantum theory. Theoretica Chimica Acta, 85(1–3):137–153, March 1993. CODEN TCHAAM. ISSN 0040-5744. Volume dedicated to Alberte Pullman.

Lowdin:1993:SPG

[Löw93f] Per-Olov Löwdin. On some properties of general linear operators. International Journal of Quantum Chemistry, 46(3):505–510, ????

1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Submitted to the Proc. Nobel Symposium in Uppsala Dec. 6–7, 1991.

Lowdin:1993:PISb

[Löw93g] Per-Olov Löwdin, editor. Proceedings of the International Symposium on the Application of Fundamental Theory to Problems of Biology and Pharmacology: held at St. Augustine, Florida, March 13–20, 1993, volume 20 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1993. CODEN IJQBDZ. ISBN 0-471-02504-6. ISSN 0360-8832. LCCN QH345.I54 1993.

Lowdin:1993:SRR

[Löw93h] Per-Olov O. Löwdin. Some remarks on the resemblance theorems associated with various orthonormalization procedures. International Journal of Quantum Chemistry, 48(4):225–232, November 10, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1994:IJQb

[Löw94a] P.-O. Löwdin. International journal of quantum chemistry — A journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. International Journal of Quantum Chemistry, 49(1):1–2, ???? 1994. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1994:A

[Löw94b] Per-Olov Löwdin. Announcement. International Journal of Quantum Chemistry, 52(5):1247, 1994. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1994:IJQa

[Löw94c] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. editorial program and policies. International Journal of Quantum Chemistry, 51 (1):1–2, ???? 1994. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1994:P

[Löw94d] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 25:ix–x, 1994. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608600167.

Lowdin:1994:PISa

[Löw94e] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods: held at the Marriott at Sawgrass Resort, Ponte Vedra Beach, Florida, February 12–19, 1994, volume 28 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1994. CODEN IJQBDZ. ISBN 0-471-12125-8. ISSN 0360-8832. LCCN QD453.A1 I58 S95, no.28.

Lowdin:1994:PISb

[Löw94f] Per-Olov Löwdin, editor. Proceedings of the International Symposium on the Application of Fundamental Theory to Problems of Biology and Pharmacology: held at Ponte Vedra Beach, Florida, February 12–19, 1994, volume 21 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1994. CODEN IJQBDZ. ISBN 0-471-12016-2. ISSN 0360-8832. LCCN QH345 .I625 1994.

Lowdin:1994:SCP

[Löw94g] Per-Olov Löwdin. Some current problems in theoretical chemical physics to be solved. International Journal of Quantum Chemistry, 51(6):473–485, ???? 1994. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Proc. Girona Congress in Theoretical Chemical Physics 1993.

Lowdin:1994:SRP

[Löw94h] Per-Olov Löwdin. Some remarks on propagators and resolvents. Theoretica Chimica Acta, 89(5–6):277–285, December 1994. CO-DEN TCHAAM. ISSN 0040-5744. Special issue in honor of Jan Linderberg, 27 October 1934 –.

Lowdin:1995:IJQb

[Löw95a] P.-O. Löwdin. International journal of quantum chemistry — A journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. International Journal of Quantum Chemistry, 56(1):1–2, October 5, 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1995:HDE

[Löw95b] Per-Olov Löwdin. The historical development of the electron correlation problem. International Journal of Quantum Chemistry, 55(2):77–102, ???? 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Proceedings of the meeting on “Electron Correlation in Atoms and Molecules: New Methods and Applications”, Smolenice Castle, Bratislava, Slovakia, 14–18 June 1994.

Lowdin:1995:IJQc

[Löw95c] Per-Olov Löwdin. International journal of quantum chemistry — A journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. International Journal of Quantum Chemistry, 54(1):1–2, ???? 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1995:IJQe

[Löw95d] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. International Journal of Quantum Chemistry, 56(1):1–2, 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1995:IJQa

[Löw95e] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. editorial program and plicies. International Journal of Quantum Chemistry, 53(1): 1–2, ???? 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1995:IJQd

[Löw95f] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. editorial program and policies. International Journal of Quantum Chemistry, 55 (1):1–2, ???? 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1995:FLC

[Löw95g] Per-Olov Löwdin. On fundamentals, logic, and the connection between the natural sciences. International Journal of Quantum Chemistry, 53(1):97–103, ???? 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Submitted to the Proc. Nobel Symposium in Uppsala Dec. 8, 1991.

Lowdin:1995:P

[Löw95h] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 26:ix–x, 1995. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608601094.

Lowdin:1995:PISa

[Löw95i] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods: held at Ponce de Leon Resort, St. Augustine, Florida, February 25–March 4, 1995, volume 29 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1995. CODEN IJQBDZ. ISSN 0360-8832. LCCN QD453.A1 I62 no.29.

Lowdin:1995:PISb

[Löw95j] Per-Olov Löwdin, editor. Proceedings of the International Symposium on the Application of Fundamental Theory to Problems of Biology and Pharmacology: held at Ponce de Leon Resort, St. Augustine, Florida, February 25–March 4, 1995, volume 22 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1995. CODEN IJQBDZ. ISSN 0360-8832. LCCN QH345.I54 1995.

Lowdin:1995:SRP

[Löw95k] Per-Olov Löwdin. Some recent progress in the quantum theory of molecular spectra. ????, ??(??):??, 1995. Contribution to the Nagakura Symposium in Hayama, Japan, Oct. 22–26, 1995. To be published in Japan.

Lowdin:1996:MJL

[Löw96a] Per-Olov Löwdin. In memory of Jean-Louis Calais. International Journal of Quantum Chemistry, 57(1):1–2, ???? 1996. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1996:IJQa

[Löw96b] Per-Olov Löwdin. International journal of quantum chemistry — A journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. International Journal of Quantum Chemistry, 58(1):1–2, ???? 1996. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1996:IJQb

[Löw96c] Per-Olov Löwdin. International Journal of Quantum Chemistry: a journal devoted to quantum theory and computations in chemistry, condensed matter physics, and biology. editorial program and policies. International Journal of Quantum Chemistry, 59 (1):1–2, ???? 1996. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1996:P

[Löw96d] Per-Olov Löwdin. Preface. Advances in Quantum Chemistry, 27:xi–xii, 1996. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S006532760860249X.

Lowdin:1996:PISa

[Löw96e] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods: held at Ponce de Leon Resort, St. Augustine, Florida, February 24–March 2, 1996, volume 60(7) of International journal of quantum chemistry. John Wiley and Sons, Inc., New York, NY, USA, December 15, 1996. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). LCCN ???? Quantum Chemistry Symposium No. 30.

Lowdin:1996:PISb

[Löw96f] Per-Olov Löwdin, editor. Proceedings of the International Symposium on the Application of Fundamental Theory to Problems of Biology and Pharmacology: held at Ponce de Leon Resort, St. Augustine, Florida, February 24–March 2, 1996, volume 60(8) of International journal of quantum chemistry. John Wiley and Sons, Inc., New York, NY, USA, December 20, 1996. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). LCCN ???? Quantum Biology Symposium No. 23.

Lowdin:1996:SCP

[Löw96g] Per-Olov Löwdin. Some comments on the properties of unitary transformations on linear spaces having an indefinite metric and the connection with the theory of spin. International Journal of Quantum Chemistry, 58(6):549–568, ???? 1996.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-bin/

abstract?ID=60513. Paper based on a Lecture given at the Symposium on “The Chemical Bond”, in Copenhagen, Denmark, in honor of Jens Peder Dahl’s 60th Birthday, July 28–29, 1994.

Lowdin:1996:SPL

[Löw96h]  Per-Olov Löwdin.  Some properties of linear functionals

and adjoint operators. Advances in Quantum Chemistry, 27:371–397, 1996. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608602555.

Lowdin:1997:RPV

[Löw97a] P.-O. Löwdin. On the representability problem in variational quantum theory. Abstracts of Papers of the American Chemical Society, 213(??):82–COMP, April 1997. CODEN ACSRAL. ISSN 0065-7727.

Lowdin:1997:PISa

[Löw97b] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods: held at Ponce de Leon Resort, St. Augustine, Florida, March 1–March 7, 1997, volume 65(5) of International journal of quantum chemistry. John Wiley and Sons, Inc., New York, NY, USA, 1997. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). LCCN QD462.A1. Quantum Chemistry Symposium No. 31.

Lowdin:1997:PISb

[Löw97c] Per-Olov Löwdin, editor. Proceedings of the International Symposium on the Application of Fundamental Theory to Problems of Biology and Pharmacology: held at Ponce de Leon Resort, St. Augustine, Florida, March 1–7, 1997, volume 65(6) of International journal of quantum chemistry. John Wiley and Sons, Inc., New York, NY, USA, 1997. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). LCCN QH345.I54 1995. Quantum Biology Symposium No. 24.

Lowdin:1997:SAD

[Löw97d] Per-Olov Löwdin. Some aspects on the development of the natural sciences and their importance for modern society and for our global environment. International Journal of Quantum Chemistry, 64(2):157–169, August 5, 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract?ID=42677;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?

ID=42677&PLACEBO=IE.pdf. Special Issue: Second Triennial Congress of the International Society for Theoretical Chemical Physics. Issue Edited by Seán P. McGlynn, Kresimir Rupnik, Janos Ladik.

Lowdin:1998:LAQ

[Löw98a] Per-Olov Löwdin. Linear Algebra for Quantum Theory. John Wiley and Sons, Inc., New York, NY, USA, 1998. ISBN 0-471-19958-3 (cloth). xi + 458 pp. LCCN QC174.17.M35 L69 1998.

Lowdin:1998:PISb

[Löw98b] Per-Olov Löwdin, editor. Proceedings of the International Symposium on the Application of Fundamental Theory to Problems of Biology and Pharmacology: held at Ponce de Leon Resort, St. Augustine, Florida, February 21–27, 1998, volume 70(6) of International journal of quantum chemistry. John Wiley and Sons, Inc., New York, NY, USA, 1998. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). LCCN QH345.I54 1995. Quantum Biology Symposium No. 25.

Lowdin:1998:SAB

[Löw98c] Per-Olov Löwdin. Some aspects on the Bloch–Lindgren equation and a comparison with the partitioning technique. Advances in Quantum Chemistry, 30:415–432, 1998. CO-DEN AQCHA9. ISBN 0-12-034830-6. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608605201.

Lowdin:1998:SCF

[Löw98d] Per-Olov Löwdin. Some comments on the foundations of physics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1998. ISBN 981-02-2913-5. xvi + 121 pp. LCCN QC173.58 .L69 1998.

Lowdin:1999:CBS

[Löw99a] Per-Olov Löwdin. Connection between semi-empirical and ab initio methods in the quantum theory of molecular electronic spectra. International Journal of Quantum Chemistry, 72(4):379–391, ???? 1999. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X

(electronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract?ID=40006316;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=40006316&PLACEBO=IE.pdf.

Special Issue: In Honor of Lionel Goodman. Issue Edited by Michael Kasha.

Lowdin:1999:PISb

[Löw99b] Per-Olov Löwdin, editor. Proceedings of the International Symposium on the Application of Fundamental Theory to Problems of Biology and Pharmacology: held at Ponce de Leon Resort, St. Augustine, Florida, February 27–March 5, 1999, volume 75(6) of International journal of quantum chemistry. John Wiley and Sons, Inc., New York, NY, USA, 1999. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic). LCCN QH345.I54 1995. Quantum Biology Symposium No. 26.

Lowdin:19xx:STLb

[Löwxxa] P.-O. Löwdin. Set theory and linear algebra — some mathematical tools to be used — binary product spaces, and their operators, Part II. Technical Report TN 491, Kvantkemi, Upsala Universitet, Uppsala, Sweden, 19xx. ?? pp.

Lowdin:19xx:STLa

[Löwxxb] P.-O. Löwdin. Set theory and linear algebra — some mathematical tools to be used in quantum theory, 19xx. Unpublished book manuscript.

Lowdin:19xx:SAD

[Löwxxc] Per-Olov Löwdin. Some aspects on the development of the natural sciences and their importance for modern society. ????, ??(??): ??, 19xx. Paper prepared for publication by the National Science Council of ROC, Taiwan.

Lowdin:19xx:SPL

[Löwxxd] Per-Olov Löwdin. Some properties of linear functionals and adjoint operators. ????, ??(??):??, 19xx. Ready for publication.

Lowdin:2001:QTC

[Low01] P. O. Lowdin. Quantum theory of cohesive properties of solids. Advances in Physics, 50(6):597–756, September 2001. CODEN ADPHAH. ISSN 0001-8732 (print), 1460-6976 (electronic).

Lowdin:2007:CPM

[Löw07a] Per-Olov Löwdin. Correlation Problem in Many-Electron Quantum Mechanics I. Review of Different Approaches and Discussion of Some Current Ideas, pages 207–322. John Wiley and Sons, Inc., New York, NY, USA, 2007. ISBN 0-470-14348-7.

Lowdin:2007:SABb

[Löw07b] Per-Olov Löwdin. Some Aspects of the Biological Problems of Heredity, Mutations, Aging, and Tumours in View of the Quantum Theory of the DNA Molecule, pages 177–179. John Wiley and Sons, Inc., New York, NY, USA, 2007. ISBN 0-470-14354-1.

Lowdin:2007:SAC

[Löw07c] Per-Olov Löwdin. Some Aspects on the Correlation Problem and Possible Extensions of the Independent-Particle Model, pages 283– 340. John Wiley and Sons, Inc., New York, NY, USA, 2007. ISBN 0-470-14359-2.

Lowdin:2007:SRDa

[Löw07d] Per-Olov Löwdin. Some Recent Developments in the Quantum Theory of Many-Electron Systems and the Correlation Problem, pages 2–4. John Wiley and Sons, Inc., New York, NY, USA, 2007. ISBN 0-470-14354-1.

Lowdin:1964:MOC

[LP64] Per-Olov Löwdin and Bernard Pullman, editors. Molecular Orbitals in Chemistry, Physics, and Biology: A Tribute to R. S. Mulliken. Academic Press Inc., New York, USA, 1964. xiii + 578 pp. LCCN QD461 .L872 1964.

Lowdin:1983:NHQ

[LP83] Per-Olov Löwdin and Bernard Pullman, editors. New horizons of quantum chemistry: proceedings of the Fourth International Congress of Quantum Chemistry, held at Uppsala, Sweden, June 14–19, 1982. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1983. ISBN 90-277-1526-2. LCCN QD462.A1 I57 1982.

Lowdin:1960:CIO

[LPdH60] P.-O. Löwdin, R. Pauncz, and J. de Heer. On the calculation of the inverse of the overlap matrix in cyclic systems. Journal of Mathematical Physics, 1(6):461–467, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v1/i6/p461_s1.

Lykos:1963:DHF

[LPLH63] P. Lykos, G. W. Pratt, P.-O. Löwdin, and Gilda M. Harris. Discussion on “The Hartree–Fock Approximation”. Reviews of Modern Physics, 35(3):496–501, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Lowdin:1959:CUM

[LR59] Per-Olov Löwdin and Lajos Rèdei. Combined use of the methods of superposition of configurations and correlation factor on the ground states of the helium-like atoms. Physical Review, 114 (3):752–757, May 1, 1959. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v114/i3/p752_1.

Lowdin:1951:NNC

[LS51] Per-Olov Löwdin and Alf Sjölander. A note on the numerical calculation of asymptotic phases with a numerical study of Hulthén’s variational principle. Arkiv för fysik / utgivet av K. Svenska Vetenskapsakademien, 3(11):155–166, 1951. CODEN AFYSA7. ISSN 0365-2440.

Lowdin:1954:NEE

[LS54] Per-Olov Löwdin and H. Sponer. Les niveaux d’energie electronique dans l’ethylene. Le Journal de physique et le radium. Series 8, 15(??):607–611, 1954.

Lowdin:1955:LEN

[LS55] Per-Olov Löwdin and H. Shull. Letter to the editor: Natural spin-orbitals for helium. Journal of Chemical Physics, 23(8): 1565, August 1955. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://jcp.aip.org/resource/1/
jcpsa6/v23/i8/p1565_s1.

Lowdin:1956:NOQ

[LS56] Per-Olov Löwdin and Harrison Shull. Natural orbitals in the quantum theory of two-electron systems. Physical Review, 101 (6):1730–1739, March 15, 1956. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v101/i6/p1730_1.

Lowdin:1976:PISb

[LS76] Per-Olov Löwdin and John R. Sabin, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology, held at Sanibel Island, Florida, January 12–14, 1976, volume 3 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1976. CODEN IJQBDZ. ISBN ???? ISSN 0360-8832. LCCN QH 345 I609 1976.

Lowdin:1977:PISb

[LS77] Per-Olov Löwdin and John R. Sabin, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology, held at Sanibel Island, Florida, January 9– 12, 1977, volume 4 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1977. CODEN IJQBDZ. ISSN 0360-8832. LCCN QH 345 I609 1977; QH345.I53 1977.

Lowdin:1978:PISa

[LS78] Per-Olov Löwdin and John R. Sabin, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at Flagler Beach, Florida, March 3–11, 1978, volume 5 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1978. CODEN IJQBDZ. ISBN 0-471-05635-9 (paperback). ISSN 0360-8832. LCCN QH345.I53 1978.

Lowdin:1979:PISb

[LS79] Per-Olov Löwdin and John R. Sabin, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology, held in honor of Alberte and Bernard Pullman at Flagler Beach, Florida, March 4–10, 1979, volume 6 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1979. CODEN IJQBDZ. ISSN 0360-8832. LCCN QH 345 I609 1979; QH345.I53 1979. US$30.00 (paperback).

Lowdin:1980:PISa

[LS80] Per-Olov Löwdin and John R. Sabin, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology, held at Palm Coast, Florida, March 5–8, 1980, volume 7 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1980. CODEN IJQBDZ. ISBN 0-471-09190-1. ISSN 0360-8832. LCCN QH345.I53 1980.

Lowdin:1981:PISb

[LS81a] Per-Olov Löwdin and John R. Sabin, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology, held at Palm Coast, Florida, March 5–7, 1981, volume 8 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1981. CODEN IJQBDZ. ISBN 0-471-86671-7. ISSN 0360-8832. LCCN QH 345 I609 1981.

Lowdin:1981:PISc

[LS81b] Per-Olov Löwdin and John R. Sabin. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology Held in Palm Coast, Florida, 5–7 March 1981 — Introduction. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 8:vii, 1981. CODEN IJQBDZ. ISSN 0360-8832.

Lowdin:1982:PISc

[LS82a] P.-O. Löwdin and J. R. Sabin. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology Held at Palm Coast, Florida, March 4–6, 1982 — Introduction. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 9:vii, 1982. CO-DEN IJQBDZ. ISSN 0360-8832.

Lowdin:1982:PISb

[LS82b] Per-Olov Löwdin and John R. Sabin, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at Palm Coast, Florida, March 4–6, 1982, volume 9 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1982. CODEN IJQBDZ. ISBN 0-471-89123-1. ISSN 0360-8832. LCCN QH345.I53 1982.

Lowdin:1983:ECa

[LSZ83a] Per-Olov Löwdin, John R. Sabin, and Michael Zerner. Editorial comments. International Journal of Quantum Chemistry, 23(1): vii, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:ECb

[LSZ83b] Per-Olov Löwdin, John R. Sabin, and Michael Zerner. Editorial comments. International Journal of Quantum Chemistry, 23(2): vii, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:ECc

[LSZ83c] Per-Olov Löwdin, John R. Sabin, and Michael Zerner. Editorial comments. International Journal of Quantum Chemistry, 23(3): v, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:ECd

[LSZ83d] Per-Olov Löwdin, John R. Sabin, and Michael Zerner. Editorial comments. International Journal of Quantum Chemistry, 23(4): xi, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1984:PISa

[LSZ84a] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry, held at Flagler Beach, Florida, March 1–15, 1984, volume 18 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1984. CODEN IJQBDZ. ISBN 0-471-81329-X. ISSN 0360-8832. LCCN QD462 .I61 1984.

Lowdin:1984:PISb

[LSZ84b] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at Palm Coast, Florida, March 12–15, 1984, volume 11 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1984. CODEN IJQBDZ. ISBN 0-471-81331-1. ISSN 0360-8832. LCCN QH345.I53 1984.

Lowdin:1984:PISd

[LSZ84c] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology Held at Palm Coast, Florida, March 12–15, 1984 — Introduction. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 11:vii–viii, 1984. CODEN IJQBDZ. ISSN 0360-8832.

Lowdin:1988:AQC

[LSZ88] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Advances in Quantum Chemistry, volume 19 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1988. CODEN AQCHA9. ISBN 0-12-034819-5 (hardcover). ISSN 0065-3276. xii + 365 pp. LCCN QD453 .A28

v.19. URL http://www.sciencedirect.com/science/article/
pii/S0065327608606085;   http://www.sciencedirect.com/
science/book/9780120348190.

Lowdin:1989:AQC

[LSZ89] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Advances in Quantum Chemistry, volume 20 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1989. CODEN AQCHA9. ISBN 0-12-034820-9 (hardcover). ISSN 0065-3276. xiii + 453 pp. LCCN QD453 .A28

v.20. URL http://www.sciencedirect.com/science/article/
pii/S0065327608606218;   http://www.sciencedirect.com/
science/book/9780120348206.

Lowdin:1991:AQCa

[LSZ91] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Advances in Quantum Chemistry, volume 22 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1991. CODEN AQCHA9. ISBN 0-12-034822-5 (hardcover). ISSN 0065-3276. x + 385 pp. LCCN QD453 .A28

v.22. URL http://www.sciencedirect.com/science/article/
pii/S0065327608603573;   http://www.sciencedirect.com/
science/book/9780120348220.

Lowdin:1992:AQCb

[LSZ92a] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Advances in Quantum Chemistry, volume 23 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1992. CODEN AQCHA9. ISBN 0-12-034823-3 (hardcover). ISSN 0065-3276. x + 363 pp. LCCN QD453 .A28

v.23. URL http://www.sciencedirect.com/science/article/
pii/S0065327608600246;   http://www.sciencedirect.com/
science/book/9780120348237.

Lowdin:1992:AQCc

[LSZ92b] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Advances in Quantum Chemistry, volume 24 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1992. CODEN AQCHA9. ISBN 0-12-034824-1 (hardcover). ISSN 0065-3276. x + 298 pp. LCCN QD453 .A28

v.24. URL http://www.sciencedirect.com/science/article/
pii/S0065327608600969;   http://www.sciencedirect.com/
science/book/9780120348244.

Lowdin:1994:AQC

[LSZ94] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Advances in Quantum Chemistry, volume 25 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1994. CODEN AQCHA9. ISBN 0-12-034825-X (hardcover). ISSN 0065-3276. x + 320 pp. LCCN QD453 .A28

v.25. URL http://www.sciencedirect.com/science/article/
pii/S0065327608600131;   http://www.sciencedirect.com/
science/book/9780120348251.

Lowdin:1995:AQC

[LSZ95] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Advances in Quantum Chemistry, volume 26 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1995. CODEN AQCHA9. ISBN 0-12-034826-8 (hardcover). ISSN 0065-3276. x + 367 pp. LCCN QD453 .A28

v.26. URL http://www.sciencedirect.com/science/article/
pii/S0065327608601069;   http://www.sciencedirect.com/
science/book/9780120348268.

Lowdin:1996:AQC

[LSZ96] Per-Olov Löwdin, John R. Sabin, and Michael C. Zerner, editors. Advances in Quantum Chemistry, volume 27 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1996. CODEN AQCHA9. ISBN 0-12-034827-6 (hardcover). ISSN 0065-3276. xii + 407 pp. LCCN QD453

.A28 v.27. US$126.00. URL http://www.apcatalog.com/cgi-
bin/AP?ISBN=0120348276&LOCATION=US&FORM=FORM2;   http:
//www.hbuk.co.uk/ap/;   http://www.sciencedirect.com/
science/book/9780120348275.

Lowdin:1997:AQCb

[LSZ^{+}97a] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, Erkki Brändas, Laszlo Kövér, Jun Kawai, and Hirohiko Adachi, editors. Advances in Quantum Chemistry: Electronic structure of clusters, volume 29 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1997. CODEN AQCHA9. ISBN 0-12-034829-2 (hardcover). ISSN 0065-3276. xvi + 471 pp. LCCN QD453 .A28 v.29. US$110.00. URL http://www.

apcatalog.com/cgi-bin/AP?ISBN=0120348292&LOCATION=US&
FORM=FORM2;   http://www.hbuk.co.uk/ap/;   http://www.
sciencedirect.com/science/book/9780120348299. This vol-

ume includes the papers presented at the first International Workshop on DV-X Methods. DVA-X ’96 took place in the ATOMKI, Debrecen, Hungary, on 2–3 September 1996.

Lowdin:1997:AQCa

[LSZ^{+}97b] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, Jacek Karwowski, and Mati Karelson, editors. Advances in Quantum Chemistry: Recent advances in computational chemistry, volume 28 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1997. CODEN AQCHA9. ISBN 0-12-034828-4 (hardcover). ISSN 0065-3276. xix + 370 pp. LCCN QD453

.A28 v.28. US$139.95. URL http://www.apcatalog.com/cgi-
bin/AP?ISBN=0120348284&LOCATION=US&FORM=FORM2;   http:
//www.hbuk.co.uk/ap/;   http://www.sciencedirect.com/
science/article/pii/S0065327608601975;   http://www.sciencedirect.
com/science/book/9780120348282. Papers written to honor

Geerd H. F. Diercksen on his 60th birthday. Most of the papers were presented at the symposium Recent Advances in Computational Chemistry (Munich, 20–30 March 1996).

Lowdin:1998:AQCa

[LSZ^{+}98a] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, Erkki Brändas, Dag Hanstorp, and Hans Persson, editors. Advances in quantum chemistry: Modern trends in atomic physics, volume 30. Academic Press Inc., New York, USA, 1998. CODEN AQCHA9. ISBN 0-12-034830-6. ISSN 0065-3276. xiii + 506 pp. LCCN QD453 .A28. US$125.00. URL http://www.apcatalog.

com/cgi-bin/AP?ISBN=0120348306&LOCATION=US&FORM=FORM2;
 http://www.hbuk.co.uk/ap.

Lowdin:1998:AQCd

[LSZ^{+}98b] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, Erkki Brändas, and Jorge M. Seminario, editors. Advances in Quantum Chemistry: Density functional theory, volume 33 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1998. CODEN AQCHA9. ISBN 0-12-034832-2 (hardcover). ISSN 0065-3276. xviii + 398 pp. LCCN QD453 .A28

v.33. US$120.00. URL http://www.academicpress.com/pecs/
chem/serials.html;   http://www.apcatalog.com/cgi-bin/
AP?ISBN=0120348322&LOCATION=US&FORM=FORM2;   http://www.
hbuk.co.uk/ap/;   http://www.sciencedirect.com/science/
book/9780120348329.

Lowdin:1998:AQCb

[LSZ^{+}98c] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, Erkki Brändas, S. Wilson, J. Maruani, Y. G. Smeyers, P. J. Grout, and R. McWeeny, editors. Advances in Quantum Chemistry: Quantum systems in chemistry and physics, part I, volume 31 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1998. CODEN AQCHA9. ISBN 0-12-034831-4 (hardcover). ISSN 0065-3276. xix + 345 pp. LCCN QD453 .A28

v.31. US$99.95. URL http://www.academicpress.com/pecs/
chem/serials.html;   http://www.apcatalog.com/cgi-bin/
AP?ISBN=0120348314&LOCATION=US&FORM=FORM2;   http://www.
hbuk.co.uk/ap/;   http://www.sciencedirect.com/science/
article/pii/S0065327608601744;   http://www.sciencedirect.
com/science/article/pii/S0065327608601756;   http://www.
sciencedirect.com/science/book/9780120348312. Copyright

year 1999.

Lowdin:1998:AQCc

[LSZ^{+}98d] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, Erkki Brändas, S. Wilson, J. Maruani, Y. G. Smeyers, P. J. Grout, and R. McWeeny, editors. Advances in Quantum Chemistry: Quantum systems in chemistry and physics, part II, volume 32 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1998. CODEN AQCHA9. ISBN 0-12-034833-0 (hardcover). ISSN 0065-3276. xix + 338 pp. LCCN QD453 .A28

v.32. US$99.95. URL http://www.academicpress.com/pecs/
chem/serials.html;   http://www.apcatalog.com/cgi-bin/
AP?ISBN=0120348330&LOCATION=US&FORM=FORM2;   http://www.
hbuk.co.uk/ap/;   http://www.sciencedirect.com/science/
article/pii/S0065327608604001;   http://www.sciencedirect.

com/science/article/pii/S0065327608604220;   http://www.
sciencedirect.com/science/book/9780120348336.

Lowdin:1999:AQCc

[LSZ^{+}99a] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, Erkki Brändas, Alessandro Lami, and Vincenzo Barone, editors. Advances in Quantum Chemistry: From electronic structure to timedependent processes: a volume in honor of Giuseppe Del Re, volume 36 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1999. CODEN AQCHA9. ISBN 0-12-034836-5 (hardcover). ISSN 0065-3276. xxii + 394 pp. LCCN QD453 .A28

v.36. US$145.00. URL http://www.academicpress.com/pecs/
chem/serials.html;   http://www.apcatalog.com/cgi-bin/
AP?ISBN=0120348365&LOCATION=US&FORM=FORM2;   http://www.
hbuk.co.uk/ap/;   http://www.sciencedirect.com/science/
article/pii/S0065327608604712;   http://www.sciencedirect.
com/science/book/9780120348367. Copyright year 2000.

Lowdin:1999:AQCb

[LSZ^{+}99b] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, Erkki Brändas, J. V. Ortiz, and Henry Kurtz, editors. Advances in Quantum Chemistry: Propagating insight: A tribute to Yngve Öhrn, volume 35 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1999. CODEN AQCHA9. ISBN 0-12-034835-7 (hardcover). ISSN 0065-3276. x + 380 pp. LCCN QD453 .A28 v.35. US$145.00. URL http://

www.academicpress.com/pecs/chem/serials.html;   http:/
/www.apcatalog.com/cgi-bin/AP?ISBN=0120348357&LOCATION=
US&FORM=FORM2;   http://www.hbuk.co.uk/ap/;   http://
www.sciencedirect.com/science/article/pii/S0065327608604487;
 http://www.sciencedirect.com/science/book/9780120348350.

Lowdin:2000:AQCa

[LSZ^{+}00] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, Erkki Brändas, Hirohiko Adachi, Masayuki Uda, and Rika Sekine, editors. Advances in Quantum Chemistry: DV-Xα for atomic spectroscopy and materials science, volume 37 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 2000. CODEN AQCHA9. ISBN 0-12-034837-3 (hardcover). ISSN 0065-3276. xvi + 393 pp. LCCN QD453 .A28

v.37. US$145.00. URL http://www.academicpress.com/pecs/
chem/serials.html;   http://www.apcatalog.com/cgi-bin/
AP?ISBN=0120348373&LOCATION=US&FORM=FORM2;   http://www.

hbuk.co.uk/ap/;   http://www.sciencedirect.com/science/

book/9780120348374. Copyright year 2001.

Lowdin:1999:AQCa

[LSZB99] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, and Erkki Brändas, editors. Advances in Quantum Chemistry, volume 34 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1999. CODEN AQCHA9. ISBN 0-12-034834-9 (hardcover). ISSN 0065-3276. x + 388 pp. LCCN QD453 .A28

v.34. US$125.00. URL http://www.academicpress.com/pecs/
chem/serials.html;   http://www.apcatalog.com/cgi-bin/
AP?ISBN=0120348349&LOCATION=US&FORM=FORM2;   http://www.
hbuk.co.uk/ap/;   http://www.sciencedirect.com/science/
article/pii/S0065327608605262;   http://www.sciencedirect.
com/science/book/9780120348343.

Lowdin:2000:AQCb

[LSZB00] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, and Erkki Brändas, editors. Advances in Quantum Chemistry: From electronic structure to time-dependent processes, volume 38 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 2000. CODEN AQCHA9. ISBN 0-12-034838-1 (hardcover). ISSN 0065-3276. ix + 287 pp. LCCN QD453 .A28

v.38. US$145.00. URL http://www.academicpress.com/pecs/
chem/serials.html;   http://www.apcatalog.com/cgi-bin/
AP?ISBN=0120348381&LOCATION=US&FORM=FORM2;   http://www.
hbuk.co.uk/ap/;   http://www.sciencedirect.com/science/
book/9780120348381.

Lowdin:1989:PISa

[LSZÖ89] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, and N. Yngve Öhrn, editors. Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Molecular Dynamics: held at St. Augustine, Florida, April 1–8, 1989, volume 23 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1989. CODEN IJQBDZ. ISBN 0-471-52597-9. ISSN 0360-8832. LCCN QD462 .I61 1989.

Lowdin:1990:AQC

[LSZT90] Per-Olov Löwdin, John R. Sabin, Michael C. Zerner, and Samuel B. Trickey, editors. Advances in Quantum Chemistry: Density functional theory of many-fermion systems, volume 21 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 1990. CODEN AQCHA9. ISBN 0-12-034821-7 (hardcover). ISSN 0065-3276. xii + 405 pp. LCCN QD453 .A28

v.21. URL http://www.sciencedirect.com/science/article/
pii/S0065327608605857;   http://www.sciencedirect.com/
science/book/9780120348213.

Lundqvist:1976:SDC

[Lun76] Stig Lundqvist. Static and dynamic correlations in many-electron systems. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 201–214. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_13/.

Bidon-Chanal:2004:SSI

[Luq04] Axel Bidon-Chanal F. Javier Luque. Solute–solvent interactions from QM SCRF methods. In Brändas and Kryachko [BK04], pages 475–495. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-

book). URL http://link.springer.com/book/10.1007/978-
94-017-0448-9.

Lowdin:1983:PISc

[LZ83a] P.-O. Löwdin and M. Zerner. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology — Held at Palm Coast, Florida, March 14–16, 1983 — Introduction. International Journal of Quantum Chemistry. Quantum Biology Symposium: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology., 10:vii, 1983. CODEN IJQBDZ. ISSN 0360-8832.

Lowdin:1983:I

[LZ83b] Per-Olov Löwdin and Michael Zerner. Introduction. International Journal of Quantum Chemistry, 24(S10):vii, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:PISb

[LZ83c] Per-Olov Löwdin and Michael C. Zerner, editors. Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology: held at Palm Coast, Florida, March 14–16, 1983, volume 10 of International journal of quantum chemistry. Quantum biology symposium. John Wiley and Sons, Inc., New York, NY, USA, 1983. CODEN IJQBDZ. ISBN 0-471-88168-6. ISSN 0360-8832. LCCN QH345.I53 1983.

Lowdin:1991:CCP

[LZ91] Per-Olov Löwdin and Michael Zerner. Call for contributions to paper symposium on semi-empirical methods in quantum chemistry and solid-state theory. International Journal of Quantum Chemistry, 40(2):287, 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1992:IPS

[LZ92] Per-Olov Löwdin and Michael C. Zerner. Introduction to the paper symposium on semiempirical methods in quantum chemistry and solid-state theory. International Journal of Quantum Chemistry, 44(4):421–422, October 15, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1994:IPS

[LZ94] Per-Olov Löwdin and Michael C. Zerner. Introduction to a paper symposium on “the role of quantum chemistry as a foundation for the principles of chemistry”. International Journal of Quantum Chemistry, 49(3):129–130, January 20, 1994. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1984:PISc

[LZS84] P.-O. Löwdin, M. C. Zerner, and J. R. Sabin. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry, Held at Flagler Beach, Florida, March 1–15, 1984 — Introduction. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 18:iii–iv, 1984. CODEN IJQSDI. ISSN 0161-3642.

Maksic:1981:COB

[Mak81] Z. B. Maksic. Construction of optimal basis-sets in approximate theories of the electronic-structure of molecules — weighted Löwdin orthogonalization. Zeitschrift für Naturforschung Section A — a Journal of Physical Sciences, 36(4):373–377, ???? 1981. ISSN 0932-0784 (print), 1865-7109 (electronic).

Malli:2004:RQC

[Mal04] Gulzari L. Malli. Relativistic quantum chemistry of heavy and superheavy elements: Fully relativistic coupled-cluster calculations for molecules of heavy and transactinide superheavy elements. In Brändas and Kryachko [BK04], pages 323–363. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http://link.springer.

com/book/10.1007/978-94-017-0448-9.

Manne:1976:POL

[Man76] Rolf Manne. Per-Olov Löwdin in the scientific literature 1965– 1974. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 25–31. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_3/.

Mayer:2002:HFM

[May02a] I. Mayer. Hermitian Fock matrix in the chemical Hamiltonian approach: Fulfilling Löwdin’s prediction. International Journal of Quantum Chemistry, 90(1):89–91, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:2002:LMS

[May02b] I. Mayer. On Löwdin’s method of symmetric orthogonalization. International Journal of Quantum Chemistry, 90(1):63–65, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:2010:LPT

[May10] I. Mayer. Löwdin’s pairing theorem and some of its applications. Molecular Physics, 108(21–23):3273–3278, ???? 2010. CODEN MOPHAM. ISSN 0026-8976 (print), 1362-3028 (electronic).

Mamedov:2009:CTCb

[MC09] B. A. Mamedov and E. Copuroglu. Calculation of two-center nuclear attraction integral over Slater type orbital in molecular coordinate system using Löwdin alpha-radial function and Guseinov one-center charge density expansion formulae. Match — Communications in Mathematical and in Computer Chemistry, 61(2): 553–560, ???? 2009. ISSN 0340-6253.

Mamedov:2010:CTC

[MÇ10a] B. A. Mamedov and E. Çopuroğlu. Calculation of three-center nuclear attraction integral over Slater type orbitals in molecular coordinate system using Löwdin α-radial function and Guseinov’s two-center charge density expansion formulae. Journal of Mathematical Chemistry, 47(1):345–353, January 2010.

CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (elec-

tronic). URL http://link.springer.com/article/10.1007/

s10910-009-9577-6. See erratum [MÇ10b].

Mamedov:2010:ECT

[MÇ10b] B. A. Mamedov and E. Çopuroğlu. Erratum: Calculation of three-center nuclear attraction integral over Slater type orbitals in molecular coordinate system using Löwdin α-radial function and Guseinov’s two-center charge density expansion formulae. Journal of Mathematical Chemistry, 47(1):354, January 2010. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10910-009-9601-x;   http://link.springer.com/content/
pdf/10.1007/s10910-009-9601-x.pdf. See [MÇ10a].

Mamedov:2011:UBC

[MÇ11a] B. A. Mamedov and E. Çopuroğlu. Use of binomial coefficients in fast and accurate calculation of Löwdin-alpha radial functions. Journal of Mathematical Chemistry, 49(1):201–207, January 2011. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10910-010-9735-x.

Mamedov:2011:UGO

[MC11b] B. A. Mamedov and E. Copuroglu. Use of Guseinov’s one-center expansion formulae and Löwdin alpha radial function in calculation of two-center overlap integrals over Slater type orbitals with noninteger principal quantum numbers. Acta Physica Polonica A, 119(3):332–335, March 2011. CODEN ATPLB6. ISSN 0587-4246 (print), 1898-794x (electronic).

Mamedov:2012:CTC

[MÇ12] Bahtiyar A. Mamedov and Ebru Çopuroğlu. Calculation of twocenter nuclear attraction integrals of Slater type orbitals with noninteger principal quantum numbers using Guseinov’s onecenter expansion formulas and Löwdin-α radial function. Applied Mathematics and Computation, 218(15):7893–7897, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300312001208.

McIntosh:1976:QGF

[McI76] Harold V. McIntosh. Quantization and a Green’s function for systems of linear ordinary differential equations. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 277–294. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_17/.

McLaughlin:1981:BRB

[McL81] Don R. McLaughlin. Book review: Quantum Chemistry — A scientific melting pot, edited by Per-Olov Löwdin, Jean-Louis Calais and Osvaldo Goscinski. Journal of Chemical Education, 58(4): A153–??, April 1981. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Meissner:2003:LPT

[Mei03] H. Meissner. On Löwdin’s projection technique and the energycorrected approaches. Journal of Chemical Physics, 119(8):4126– 4133, August 2003. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Mezey:1997:QSM

[Mez97] Paul G. Mezey. Quantum similarity measures and Löwdin’s transform for approximate density matrices and macromolecular forces. International Journal of Quantum Chemistry, 63(1):39–48, May 15, 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X

(electronic). URL http://www3.interscience.wiley.com/
cgi-bin/abstract?ID=42579;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=42579&PLACEBO=IE.pdf.

Mayer:2004:TNP

[MH04] István Mayer and Andrea Hamza. Treating nonadditivity as a perturbation: a quasi-particle formalism. In Brändas and Kryachko [BK04], pages 187–198. ISBN 90-481-6687-X (print), 94-017-0448-

1 (e-book). URL http://link.springer.com/book/10.1007/
978-94-017-0448-9.

Micha:1976:MBT

[Mic76] David A. Micha. Many-body theory of molecular collisions. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 367–380. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_24/.

Mamedov:2008:CTC

[MK08] B. A. Mamedov and H. Koç. Calculation of two-center overlap integral in molecular coordinate system over Slater type orbital using Löwdin α-radial and Guseinov rotation-angular functions. Journal of Mathematical Chemistry, 44(2):365–372, August 2008. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (elec-

tronic). URL http://link.springer.com/article/10.1007/

s10910-007-9313-z. See corrigendum [MK09].

Mamedov:2009:CTCa

[MK09] B. A. Mamedov and H. Koç. Calculation of two-center overlap integral in molecular coordinate system over Slater type orbital using Löwdin α-radial and Guseinov rotation-angular functions [vol 44, pg 365, 2008]. Journal of Mathematical Chemistry, 45(3):867, March 2009. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-

8897 (electronic). URL http://link.springer.com/article/
10.1007/s10910-008-9449-5;   http://link.springer.com/
content/pdf/10.1007/s10910-008-9449-5.pdf. See [MK08].

Macintyre:1968:EED

[ML68] Walter M. Macintyre and Per-Olov Löwdin. Electronic energy of the DNA replication plane. International Journal of Quantum Chemistry. Symposium, 2(S2):207–217, 1968. CODEN IJQSAF. ISSN 0020-7608 (print), 1097-461X (electronic).

Micha:1983:IP

[ML83a] David A. Micha and Per-Olov Löwdin. Introduction, Part I. International Journal of Quantum Chemistry, 24(S17):1, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Micha:1983:PIS

[ML83b] David A. Micha and Per-Olov Löwdin. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry — Held at Flagler Beach, Florida, March 3–12, 1983 — Introduction. Part I. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 17:1, 1983. CODEN IJQSDI. ISSN 0161-3642.

Mayer:1993:SCG

[ML93] István Mayer and Per-Olov Löwdin. Some comments on the general Hartree–Fock method. Chemical Physics Letters, 202(1–2):1– 6, January 15, 1993. CODEN CHPLBC. ISSN 0009-2614 (print), 1873-4448 (electronic).

Morikawa:1978:VAL

[Mor78] T. Morikawa. Variation approach to Löwdin canonical orthonormalization. Physics Letters A, 68(3–4):323–325, ???? 1978. CO-DEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Mayhall:2010:OBQ

[MRH10] Nicholas J. Mayhall, Krishnan Raghavachari, and Hrant P. Hratchian. ONIOM-based QM:QM electronic embedding method using Löwdin atomic charges: energies and analytic gradients. Journal of Chemical Physics, 132(11):114107, March 21, 2010. CODEN JCPSA6. ISSN 1089-7690.

Midtdal:1968:SSPa

[MTW68] John Midtdal, Knut Thalberg, and Harald Wergeland, editors. Selected Scientific Papers of Egil A. Hylleraas, volume 1. NTH-Trykk, Trondheim, Norway, 1968. viii + 445 pp. LCCN Q143.H98 A25 1968.

Masovic:1981:BSC

[MV81] D. R. Masovic and F. Vukajlovic. Band-structure calculations of cubic semiconductors on the basis of Löwdin perturbation technique. Computer Physics Communications, 24(2):181–184, ???? 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Mahalakshmi:2004:LLI

[MY04] S. Mahalakshmi and Danny L. Yeager. Low-lying ionization potentials of B_{3}N and photodetachment energies of B_{3}N^{−} using the multiconfigurational spin tensor electron propagator method. In Brändas and Kryachko [BK04], pages 547–561. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http://link.springer.

com/book/10.1007/978-94-017-0448-9.

Nagy:2012:SEE

[NA12] I. Nagy and I. Aldazabal. Series expansions for an exact twoelectron wave function in terms of Löwdin’s renormalized natural orbitals. Physical Review A (Atomic, Molecular, and Optical Physics), 85(3):034501:1–034501:4, March 5, 2012. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Naidu:2004:LCO

[NS04] A. Ramesh Naidu and Vipin Srivastava. Löwdin’s canonical orthogonalization: Getting round the restriction of linear independence. International Journal of Quantum Chemistry, 99(6):882– 888, ???? 2004. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nagy:2012:MOR

[NSS12] Péter R. Nagy, Péter R. Surján, and Ágnes Szabados. Mayer’s orthogonalization: relation to the Gram–Schmidt and Löwdin’s symmetrical scheme. Theoretical Chemistry Accounts, 131(2): 1109, February 2012. CODEN TCACFW. ISSN 1432-881x (print), 1432-2234 (electronic).

Nye:1993:CPT

[Nye93] Mary Jo Nye. From Chemical Philosophy to Theoretical Chemistry: Dynamics of Matter and Dynamics of Disciplines, 1800– 1950. University of California Press, Berkeley, CA, USA, 1993. ISBN 0-520-08210-9. xvii + 328 pp. LCCN QD452 .N94 1993.

URL http://www.loc.gov/catdir/bios/ucal051/92043114.
html;   http://www.loc.gov/catdir/description/ucal041/
92043114.html.

Ohno:1976:POL

[Ohn76] Kimio Ohno. Per-Olov Löwdin — conqueror of scientific, educational and rocky mountains. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 1–12. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.springer.

com/chapter/10.1007/978-1-4757-1659-7_1/.

Ohrn:1983:IPI

[ÖL83] Yngve Öhrn and Per-Olov Löwdin. Introduction, Part II. International Journal of Quantum Chemistry, 24(S17):199, March 12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Ohrn:1983:PIS

[Öö83] Yngve Öhrn and Per-Olov öwdin. Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry

— Held at Flagler Beach, Florida, March 3–12, 1983 — Introduction. Part II. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 17:199, 1983. CODEN IJQSDI. ISSN 0161-3642.

Ohrn:2002:SID

[OS02] N. Y. Ohrn and J. R. Sabin. Special issue: Dedicated to Per-Olov Löwdin — Part I of II — introduction. International Journal of Quantum Chemistry, 90(1):12–13, October 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461x (electronic).

Ohrn:2001:POL

[ÖSB01] Yngve Öhrn, John R. Sabin, and Erkki Brändas. Per-Olov Löwdin 1916–2000. International Journal of Quantum Chemistry, 81(1):i–ii, ???? 2001. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract/75502160/
START;   http://www3.interscience.wiley.com/cgi-bin/fulltext/
75502160/FILE?TPL=ftx_start;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=75502160&PLACEBO=IE.pdf.

Ohrn:1987:PIS

[ÖSZL87] Yngve Öhrn, John R. Sabin, Michael C. Zerner, and Per-Olov Löwdin, editors. Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods: held at Marineland, Florida, March 12–21, 1987, volume 21 of International journal of quantum chemistry. Quantum chemistry symposium. John Wiley and Sons, Inc., New York, NY, USA, 1987. CODEN IJQBDZ. ISSN 0360-8832. LCCN QD462 .I61 1987.

Panin:1982:MRCa

[Pan82a] A. I. Panin. Method of recurrent construction of Löwdin spinadapted wave functions. I. Addition and subtraction operators. International Journal of Quantum Chemistry, 22(3):557–573, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panin:1982:MRCb

[Pan82b] A. I. Panin. Method of recurrent construction of Löwdin spinadapted wave functions. II. Local representation of creation– annihilation operators. International Journal of Quantum Chemistry, 22(6):1177–1188, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panin:1983:MRCc

[Pan83] A. I. Panin. Method of recurrent construction of Löwdin spinadapted wave functions. III. Löwdin basis and its permutation symmetry: Evaluation of overlap integrals. International Journal of Quantum Chemistry, 24(3):279–306, September 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panin:1984:MRC

[Pan84] A. I. Panin. Method of recurrent construction of Löwdin spinadapted wave functions. IV. Matrix element evaluation problem. International Journal of Quantum Chemistry, 25(5):767–793, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Parr:1966:BRB

[Par66] Robert G. Parr. Book review: Advances in quantum chemistry, by Per-Olov Löwdin. American Scientist, 54(1):114A–114A, March 1966. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.jstor.org/stable/27836334.

Park:2003:HQC

[Par03] Buhm Soon Park. The ‘hyperbola of quantum chemistry’: the changing practice and identity of a scientific discipline in the early years of electronic digital computers, 1945–65. Annals of Science, 60(3):219–247, 2003. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.ingentaconnect.com/
content/tandf/tasc/2003/00000060/00000003/art00001.

Park:2009:BAM

[Par09] Buhm Soon Park. Between accuracy and manageability: Computational imperatives in quantum chemistry. Historical Studies in the Natural Sciences, 39(1):32–62, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Patterson:1962:DMR

[Pat62] J. D. Patterson. Density matrix representations. American Journal of Physics, 30(12):894–897, December 1962. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

ajp.aapt.org/resource/1/ajpias/v30/i12/p894_s1.

Pauncz:1976:IPP

[Pau76] Ruben Pauncz. Investigations into the properties of projected spin functions. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 357–365. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_23/.

Pauncz:1962:SAMa

[PdHL62a] R. Pauncz, J. de Heer, and P.-O. Löwdin. Studies on the alternant molecular orbital method. I. General energy expression for an alternant system with closed-shell structure. Journal of Chemical Physics, 36(9):2247–2256, May 1, 1962. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://

link.aip.org/link/doi/10.1063/1.1732872.

Pauncz:1962:SAMb

[PdHL62b] R. Pauncz, J. de Heer, and P.-O. Löwdin. Studies on the alternant molecular orbital method. II. Application to cyclic systems. Journal of Chemical Physics, 36(9):2257–2265, May 1, 1962. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

URL http://link.aip.org/link/doi/10.1063/1.1732873.

PeresdeAndrade:2012:PCP

[Per12] Paulo Cesar Peres de Andrade. Probability current in protein electron transfer: Löwdin population analysis. International Journal of Quantum Chemistry, 112(20):3325–3332, October 15, 2012. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461x (electronic).

Pullman:1983:NHQ

[PL83] Bernard Pullman and Per-Olov Löwdin. New horizons of quantum chemistry: proceedings of the Fourth International Congress of Quantum Chemistry, held at Uppsala, Sweden, June 14–19, 1982. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1983. ISBN 90-277-1526-2. ix + 457 pp. LCCN QD462.A1 I57 1982. Held under the auspices of the International Academy of Quantum Molecular Science.

PonsViver:2019:PIL

[Pon19] Miquel Pons Viver. The practical implementation of Löwdin’s method for spin projection. International Journal of Quantum Chemistry, 119(4):e25770:1–e25770:??, February 15, 2019. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Yos20].

Pullman:1976:PGQ

[PP76] Alberte Pullman and Bernard Pullman. The phosphate group in quantum biochemistry. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 517–537. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.springer.

com/chapter/10.1007/978-1-4757-1659-7_38/.

Quack:2004:TTR

[Qua04] Martin Quack. Time and time reversal symmetry in quantum chemical kinetics. In Brändas and Kryachko [BK04], pages 423–474. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book).

URL http://link.springer.com/book/10.1007/978-94-017-
0448-9.

Redei:1976:LPR

[Réd76] L. B. Rédei. Laguerre polynomials, reminiscences from Uppsala. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 305–306. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online).

LCCN QC174.125 .Q371. URL http://link.springer.com/
accesspage/chapter/10.1007/978-1-4757-1659-7_19?coverImageUrl=
/static/0.7725/sites/link/images/abstract_cover_placeholder.
png.

Reid:1976:LBE

[Rei76] Charles E. Reid. Lower bounds to energy eigenvalues. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 315–347. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125

.Q371. URL http://link.springer.com/chapter/10.1007/
978-1-4757-1659-7_21/.

Reiher:2004:BRBa

[Rei04a] Markus Reiher. Book review: Fundamental World of Quantum Chemistry. A Tribute to the Memory of Per-Olov Löwdin. 2 Bände. Fundamental World of Quantum Chemistry. A Tribute to the Memory of Per-Olov Löwdin. 2 Bände. Herausgegeben von Erkki J. Brändas und Eugene S. Kryachko. Angewandte Chemie, 116(25):3281–3282, June 21, 2004. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/ange.200385109.

Reiher:2004:BRBb

[Rei04b] Markus Reiher. Book review: Fundamental World of Quantum Chemistry. A Tribute to the Memory of Per-Olov Löwdin. Vols. I and II. Edited by Erkki J. Brändas and Eugene S. Kryachko. Angewandte Chemie, International Edition, 43(25):3219–3220, June 21, 2004. CODEN ACIEF5. ISSN 1433-7851 (print), 1521-

3773 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/anie.200385109/abstract.

Rein:1976:EMA

[RG76] Robert Rein and Ramon Garduno. Energetics and mechanism of 2-aminopurine induced mutations. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 549–560. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.

springer.com/chapter/10.1007/978-1-4757-1659-7_40/.

Roos:2000:PQT

[Roo00] Björn Roos. Perspective on “Quantum theory of many-particle systems I, II, and II”, Löwdin P-O (1955) Phys. Rev. 97:1474– 1520. Theoretical Chemistry Accounts, 103(3):228–230, ???? 2000. CODEN TCACFW. ISSN 1432-881X (print), 1432-2234 (electronic). This paper gives a historical review of the influence of the famous three 1955 papers.

Ratner:1976:PMI

[RS76a] Mark A. Ratner and J. R. Sabin. On proton mobilities in individual hydrogen bonds. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 577–588. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/content/
pdf/10.1007/978-1-4757-1659-7_43.pdf.

Ruedenberg:1976:MSC

[RS76b] Klaus Ruedenberg and Kenneth R. Sundberg. MCSCF studies of chemical reactions: Natural reaction orbitals and localized reaction orbitals. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 505–515. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_37/.

Rokob:2008:NSP

[RSS08] Andras T. Rokob, Agnes Szabados, and Peter R. Surjan. A note on the symmetry properties of Löwdin’s orthogonalization schemes. Collection of Czechoslovak Chemical Communications, 73(6–7): 937–944, ???? 2008. CODEN CCCCAK. ISSN 0010-0765 (print), 1212-6950 (electronic).

Sabin:2002:AQC

[SB02] John S. Sabin and Erkki J. Brändas, editors. Advances in Quantum Chemistry: A tribute to the life and work of Per-Olov Löwdin, volume 41 of Advances in Quantum Chemistry. Academic Press Inc., New York, USA, 2002. ISBN 0-12-034841-1 (hardcover). ISSN 0065-3276. xvi + 270 pp. LCCN QD453

.A28 v.41. URL http://www.sciencedirect.com/science/

book/9780120348411. Founding editor Per-Olov Löwdin (1916– 2000).

Sabin:2017:AQCa

[SB17] John R. Sabin and Erkki Brändas, editors. Advances in Quantum Chemistry: Löwdin volume, volume 74. Elsevier, Amsterdam, The Netherlands, 2017. CODEN AQCHA9. ISBN 0-12-809988-7 (hardcover), 0-12-810400-7. ISSN 0065-3276. xiii + 392 pp. LCCN

QD453 .A28 v.74 2017. URL http://www.sciencedirect.com/
science/bookseries/00653276/74.

Scerri:2006:CAK

[Sce06] Eric R. Scerri. Commentary on Allen & Knight’s response to the Löwdin Challenge. Foundations of Chemistry, 8(3):285–292, October 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/
10.1007/s10698-006-9009-7.

Scerri:2008:CPP

[Sce08] Eric R. Scerri, editor. Collected papers on philosophy of chemistry. Imperial College Press, London, UK, 2008. ISBN 1-84816-137-9. x + 235 pp. LCCN QD6 .S34 2008. URL http://www.loc.gov/

catdir/toc/fy0903/2008278137.html.

Scerri:2009:LRA

[Sce09a] Eric R. Scerri. Löwdin’s remarks on the aufbau principle and a philosopher’s view of ab initio quantum chemistry. In Selected papers on the periodic table [Sce09b], pages 91–110. ISBN 1-84816-425-4. LCCN QD467 .S347 2009. URL http://www.

worldscientific.com/worldscibooks/10.1142/p655.

Scerri:2009:SPP

[Sce09b] Eric R. Scerri, editor. Selected papers on the periodic table. Imperial College Press, London, UK, 2009. ISBN 1-84816-425-4. ix + 143 pp. LCCN QD467 .S347 2009.

Scharfenberg:1977:NAS

[Sch77] Peter Scharfenberg. A new algorithm for the symmetric (Löwdin) orthonormalization. International Journal of Quantum Chemistry, 12(5):827–839, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schwarz:2004:TPE

[Sch04] W. H. Eugen Schwarz. Towards a physical explanation of the periodic table (PT) of chemical elements. In Brändas and Kryachko [BK04], pages 645–669. ISBN 90-481-6687-X (print), 94-017-0448-

1 (e-book). URL http://link.springer.com/book/10.1007/
978-94-017-0448-9.

Scofield:1973:NLO

[Sco73] D. F. Scofield. A note on Löwdin orthogonalization and the square root of a positive self-adjoint matrix. International Journal of Quantum Chemistry, 7(3):561–568, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seliger:1965:BRB

[Sel65] Howard H. Seliger. Book review: Molecular orbitals in chemistry, physics, and biology, edited by Per-Olov Löwdin and Bernard Pullman. Quarterly review of biology, 40(4):418–419, December 1965. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718

(electronic). URL http://www.jstor.org/stable/2819433. See

[LP64].

Simoes:2014:RPL

[SG14] Ana Simões and Kostas Gavroglu. Reviews: P.-O. Löwdin and the International Journal of Quantum Chemistry: a kaleidoscopic agenda for quantum chemistry. International Journal of Quantum Chemistry, 114(2):116–127, January 15, 2014. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schwalbe:2018:FLO

[SHL^{+}18] Sebastian Schwalbe, Torsten Hahn, Simon Liebing, Kai Trepte, and Jens Kortus. Fermi–Löwdin orbital self-interaction corrected density functional theory: Ionization potentials and enthalpies of formation. Journal of Computational Chemistry, 39(29):2463– 2471, November 5, 2018. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shull:1955:LER

[SL55] Harrison Shull and Per-Olov Löwdin. Letter to the editor: Role of the continuum in superposition of configurations. Journal of Chemical Physics, 23(7):1362, July 1955. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://jcp.aip.

org/resource/1/jcpsa6/v23/i7/p1362_s1.

Shull:1956:CSH

[SL56] Harrison Shull and Per-Olov Löwdin. Correlation splitting in helium-like ions. Journal of Chemical Physics, 25(5):1035–1040, November 1956. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://jcp.aip.org/resource/1/
jcpsa6/v25/i5/p1035_s1.

Shull:1958:LEV

[SL58a] Harrison Shull and Per-Olov Löwdin. Letter to the editor: Variation theorem for excited states. Physical Review, Series 2, 110 (6):1466–1467, June 15, 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v110/i6/p1466_1.

Shull:1958:SCN

[SL58b] Harrison Shull and Per-Olov Löwdin. Superposition of configurations and natural spin orbitals. application to the He problem. Preliminary Technical report 13, Kvantkemi, Upsala Universitet, Uppsala, Sweden, June 15, 1958. Published in [SL59].

Shull:1959:SCN

[SL59] Harrison Shull and Per-Olov Löwdin. Superposition of configurations and natural spin orbitals. Applications to the He problem. Journal of Chemical Physics, 30(3):617–626, March 1959. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Slater:1976:PSM

[Sla76] J. C. Slater. Power series methods for cellular calculations on atoms, molecules and solids. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 215– 247. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN QC174.125 .Q371. URL http://link.springer.

com/chapter/10.1007/978-1-4757-1659-7_14/.

Szczepanik:2013:SAL

[SM13] D. Szczepanik and J. Mrozek. On several alternatives for Löwdin orthogonalization. Computational & Theoretical Chemistry, 1008 (??):15–19, March 15, 2013. CODEN CTCOA5. ISSN 2210-271X.

Sidi:2014:AEB

[SPH14] A. Sidi, D. Pinchon, and P. E. Hoggan. Asymptotic expansions of Barnett–Coulson–Löwdin functions of high order. Advances in Quantum Chemistry, 68:43–76, 2014. CO-DEN AQCHA9. ISBN 0-12-800536-X. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
B9780128005361000034.

Szabados:2004:ASP

[SS04] Péter R. Surján and Ágnes Szabados. Appendix to “Studies in Perturbation Theory”: The problem of partitioning. In Brändas and Kryachko [BK04], pages 129–185. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http://link.springer.

com/book/10.1007/978-94-017-0448-9.

Srivastava:2017:CLM

[SS17] Vipin Srivastava and Suchitra Sampath. Cognition of learning and memory: What have Löwdin’s orthogonalizations got to do with that? Advances in Quantum Chemistry, 74:299–319, 2017. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327616300430.

Sovic:2005:NME

[ST05] N. Sovic and J. D. Talman. Numerical methods for the evaluation of the Löwdin alpha-function. Collection of Czechoslovak Chemical Communications, 70(8):1035–1054, ???? 2005. CODEN CCCCAK. ISSN 0010-0765 (print), 1212-6950 (electronic).

Szabados:2014:LBF

[ST14] Ágnes Szabados and Zsuzsanna Tóth. Löwdin’s bracketing function revisited. Journal of Mathematical Chemistry, 52(8):2210– 2221, September 2014. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (electronic). URL http://link.springer.

com/article/10.1007/s10910-014-0379-0.

Stepanov:1977:KSG

[Ste77] N. F. Stepanov. On the Kashiwagi–Sasaki generalization of the Löwdin orthogonalization and the inverse vibrational problem. International Journal of Quantum Chemistry, 12(2):317–326, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sutcliffe:2002:SOP

[Sut02] Brian Sutcliffe. Some observations on P.-O. Löwdin’s definition of a molecule. International Journal of Quantum Chemistry, 90 (1):66–79, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Suzuki:1984:ESC

[Suz84a] Noboru Suzuki. Erratum: “Simple calculation of Löwdin’s alphafunction”, [J. Math. Phys. 25, 1133 (1984)]. Journal of Mathematical Physics, 25(10):3135, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v25/i10/p3135_s2. See

[Suz84b].

Suzuki:1984:SCL

[Suz84b] Noboru Suzuki. Simple calculation of Löwdin’s alpha-function. Journal of Mathematical Physics, 25(4):1133–1138, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p1133_s1. See erratum [Suz84a].

Suzuki:1985:SCL

[Suz85] Noboru Suzuki. Simple calculation of Löwdin’s alpha function. II. Easier procedure for evaluating b_{Kk}(LM|l), and vanishing of h_{n,2n−i}(LM|l) for special values of i and n. Journal of Mathematical Physics, 26(12):3193–3199, December 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suzuki:1987:CCE

[Suz87] Noboru Suzuki. Comment on “Compact expression for Löwdin’s alpha function”, [J. Math. Phys. 26 (1985), no. 5, 940–942; MR 86j:33014] by A. A. Antone. Journal of Mathematical Physics, 28(4):769–773, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [Ant85].

Suzuki:1990:SCL

[Suz90] Noboru Suzuki. Simple calculation of Löwdin’s alpha function. III. Procedure for calculating h_{n,2n−i}(LM|l)/(2n−i)! successively. Journal of Mathematical Physics, 31(9):2314–2326, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suzuki:1992:SCL

[Suz92] Noboru Suzuki. Simple calculation of Löwdin’s alpha function. IV. Procedure for evaluating the coefficients in expansion of a Slatertype atomic orbital. Journal of Mathematical Physics, 33(12): 4288–4299, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smirnov:1992:CGC

[SW92] Y. F. Smirnov and R. F. Wehrhahn. The Clebsch–Gordan coefficients for the 2-parameter quantum algebra SU_{pq}(2) in the Löwdin–Shapiro approach. Journal of Physics A (Mathematical and General), 25(21):5563–5576, November 7, 1992. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Slater:19xx:B

[SYLxx] J. C. Slater, H. Yoshizumi, and P.-O. Löwdin. Bibliographies. Uppsala and Florida use only., 19xx.

Tchougreeff:1999:GFL

[Tch99] A. L. Tchougreeff. Group functions, Löwdin partition, and hybrid QC/MM methods for large molecular systems. Physical Chemistry Chemical Physics: PCCP, 1(6):1051–1060, March 15, 1999. CO-DEN PPCPFQ. ISSN 1463-9076 (print), 1463-9084 (electronic).

Taurian:1981:SRP

[TL81] Oscar E. Taurian and Per-Olov Löwdin. Some remarks on the projector associated with the intersection of 2 linear manifolds. Acta physica Academiae Scientiarum Hungaricae, 51(1–2):5–12, ???? 1981. CODEN APAHAQ. ISSN 0001-6705.

Taurian:1982:SRP

[TL82] Oscar E. Taurian and Per-Olov Löwdin. Some remarks on the projector associated with the intersection of two linear manifolds. Acta physica Academiae Scientiarum Hungaricae, 51(1–2):5–12, 1982. CODEN APAHAQ. ISSN 0001-6705. R. Gaspar dedicatory volume.

Trepte:2019:AAG

[TSH^{+}19] Kai Trepte, Sebastian Schwalbe, Torsten Hahn, Jens Kortus, Der-You Kao, Yoh Yamamoto, Tunna Baruah, Rajendra R. Zope, Kushantha P. K. Withanage, Juan E. Peralta, and Koblar A. Jackson. Analytic atomic gradients in the Fermi–Löwdin orbital self-interaction correction. Journal of Computational Chemistry, 40(6):820–825, March 5, 2019. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tsipis:2004:ECC

[TTK04] C. A. Tsipis, A. C. Tsipis, and C. E. Kefalidis. Exploring the catalytic cycle of the hydrosilylation of alkenes catalyzed by hydridobridged diplatinum complexes using electronic structure calculation methods. In Brändas and Kryachko [BK04], pages 619–643. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-book). URL http:

//link.springer.com/book/10.1007/978-94-017-0448-9.

Tulub:2004:CMS

[Tul04] A. V. Tulub. The cavity model with a surface formed by two intersecting spheres. an analytical treatment. In Brändas and Kryachko [BK04], pages 497–520. ISBN 90-481-6687-X (print), 94-017-0448-

1 (e-book). URL http://link.springer.com/book/10.1007/
978-94-017-0448-9.

UziKaldor:2004:SHE

[Uzi04] Ephraim Eliav Uzi Kaldor, Arie Landau. Study of heavy elements by relativistic Fock space and intermediate Hamiltonian coupled cluster methods. In Brändas and Kryachko [BK04], pages 365–406. ISBN 90-481-6687-X (print), 94-017-0448-1 (e-

book). URL http://link.springer.com/book/10.1007/978-
94-017-0448-9.

VanLeuven:1976:DAP

[Van76] P. Van Leuven. Deformed atoms and the Projected Hartree– Fock method. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 499–504. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_36/.

Varga:1975:LPO

[VB75] J. A. Varga and C. A. L. Becker. Löwdin projection operator and construction of strong crystalline field D_{3} eigenfunctions. Texas Journal of Science, 26(3–4):577–584, ???? 1975. CODEN TJS-CAU. ISSN 0040-4403.

Vikictopic:1995:LAC

[VBHG95] D. Vikictopic, E. D. Becker, M. Hodoscek, and A. Graovac. Löwdin atomic charges for modeling long-range deuterium-isotope shifts in C-13 NMR-spectra of binuclear aromatic compounds. Croatica Chemica Acta, 68(1):193–203, April 1995. CODEN CCA-CAA. ISSN 0011-1643 (print), 1334-417x (electronic).

Weinhold:1976:PTD

[Wei76] Frank Weinhold. Partitioning technique for determinantal equations. In Quantum science: methods and structure: a tribute to Per-Olov Löwdin [CGLÖ76b], pages 307–314. ISBN 0-306-30968-8, 1-4757-1661-3 (print), 1-4757-1659-1 (online). LCCN

QC174.125 .Q371. URL http://link.springer.com/chapter/
10.1007/978-1-4757-1659-7_20/.

Walk:2012:ALO

[WJ12] Philipp Walk and Peter Jung. Approximation of Löwdin orthogonalization to a spectrally efficient orthogonal overlapping PPM design for UWB impulse radio. Signal Processing, 92(3):649–666, March 2012. CODEN SPRODR. ISSN 0165-1684 (print), 1872-7557 (electronic).

Wilson:1981:SLP

[WS81] R. G. Wilson and C. S. Sharma. Studies in Löwdin’s partitioning technique in quantum theory. I. Invertibility of the operator K. International Journal of Quantum Chemistry, 20(4):927–932, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wilson:1982:SLP

[WS82] R. G. Wilson and C. S. Sharma. Studies in Löwdin’s partitioning technique in quantum theory. II. Operators T and Ω. International Journal of Quantum Chemistry, 22(1):183–190, July 1982. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wehrhahn:1992:CGC

[WS92] R. F. Wehrhahn and Yu. F. Smirnov. The Clebsch–Gordan coefficients for the two parameter quantum algebras SU_{p,q}(2) in the Löwdin–shapiro approach. Report DESY-92-024, Universität Hamburg, Hamburg, Germany, February 1992. 25 pp.

Williams:1976:LEL

[WW76] A. R. Williams and D. Weaire. Letter to the Editor: The Löwdin technique as an exact folding procedure in band-structure calculations. Journal of Physics C: Solid State Physics, 9(2):L47–L50, 1976. CODEN JPSOAW. ISSN 0022-3719. URL http://stacks.

iop.org/0022-3719/9/i=2/a=005.

Yim:1981:DFR

[YE81] Yung-Chang Yim and Henry Eyring. Derivation of a formula for the resonance integral for a nonorthogonal basis set. Proceedings of the National Academy of Sciences of the United States of America, 78(5):2649–2641, ???? 1981. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

www.pnas.org/content/78/5/2649.full.pdf+html.

Yoshizawa:2020:MDP

[Yos20] Terutaka Yoshizawa. A mathematical discussion of Pons Viver’s implementation of Löwdin’s spin projection operator. International Journal of Quantum Chemistry, 120(12):e26215:1– e26215:??, June 15, 2020. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Pon19].

Zerner:1994:IPS

[ZL94] M. C. Zerner and P.-O. Löwdin. Introduction to a paper symposium on “The Role of Quantum Chemistry as a Foundation for the Principles of Chemistry”. International Journal of Quantum Chemistry, 49(3):129–??, January 20, 1994. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).